### v3

# The physical basis of teleological evolutions

Abstract
It is easily seen that the mere existence of an optimal quantum computational speedup logically implies the mutually exclusive $or$ of well-defined causal loops. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and used this knowledge to produce the solution with correspondingly fewer computation steps. Involving only the measurements of commuting observables, quantum algorithms are submitted to the same (classical) logic and should therefore host the quantum superposition of the causal loops in question. However, being causal in character, their ordinary quantum description cannot describe causal loops. We argue that it must therefore be incomplete and show that it is completed by time-symmetrizing it. This leaves the ordinary quantum description mathematically unaltered but replaces its single causal process by the quantum superposition of causal loops in question. In their completed quantum description, quantum algorithms respond to the pre-scientific notion of teleological evolution, that is, an evolution toward a goal (the solution) with an attractor in the goal it will produce in the future (the solution again). Once applied to the evolutions of the living, this notion was dismissed with the advent of modern science for the alleged absence of a physical basis. We show that, under a quantum cosmological interpretation of the Anthropic Principle, the same quantum superposition of causal loops underlying the teleological character of quantum algorithms becomes the missing physical basis of the teleological character of natural evolutions.
1 Foreword
The present work is about causal loops, of course a puzzling notion: in our natural way of seeing the world, we firmly believe that cause always precedes effect in time. In a causal loop, the effect, in turn, changes backward in time its previous cause. This may seem paradoxical; it is in fact the source of well-known paradoxes such as the grandfather one.
So, let us say right away that the causal loops we will be dealing with are quantum and of a special time-symmetric kind that should be imposed by quantum mechanics. Therefore, as long as the latter is consistent, they should be exempt from the paradoxes that plague the usual kind. They would be the physical reason for the quantum computational speedup, the fact that solving a problem in the quantum world can require fewer logical operations (computation steps) than in the classical world222 The notion that computation is possible in the quantum world, with the related notion of quantum bit, was first introduced by Finkelstein $[1]$ in 1969. In 1982, Feynman $\left[2\right]$ observed that the classical simulation of a quantum process can require an exponentially higher amount of physical resources. In 1985, Deutsch $\left[3\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup. . More in general, they would always be present in quantum correlations.
The present work is divided into two parts.
The first is a review and further clarification of the retrocausal approach to the quantum computational speedup $\left[4-8\right]$. It concerns a (time-symmetric) quantum causal loops description of quantum algorithms that provides a unified quantitative explanation of their quantum speedups. According to it, quantum algorithms are quantum superpositions of causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future. Then, she could use this knowledge to produce the solution with fewer computation steps. In equivalent terms, the computational complexity of the problem to be solved, or the dimension of the Hilbert space in which the search for the solution takes place, correspondingly reduces.
The number of computation steps needed to solve the problem in an optimal quantum way is the number needed to solve it in an optimal logical way benefiting from the advanced knowledge in question. We call this the advanced knowledge rule.
The first part of the work ends with the observation that, under their time-symmetric quantum causal loops description, quantum algorithms turn out to be teleological evolutions, namely evolutions toward a goal (the solution of the problem) with an attractor in the goal they will produce in the future (the solution again) – a causal loop in fact.
As is well known, the evolutions of the living were often thought to be teleological in character. The teleological notion was dismissed with the advent of modern science for the alleged absence of a physical basis. That was, of course, under the notion of causal evolution that is dominant in science also today and that the present work challenges.
The second part of the work is about the possibility that the time-symmetric quantum causal loops at the physical basis of the teleological character of quantum algorithms are also the physical basis of the teleological character of the natural evolutions for which the notion of teleological evolution was conceived. We will see that, under a quantum cosmological interpretation of the Anthropic Principle, this possibility turns out to be plausible.
2 Rationale
As the present work implies a fundamentally different way of seeing the role of causality in quantum correlation, we replace the usual introduction by a more extended rationale that describes this way of seeing in sufficient detail. Quantum correlation is of course the common feature of quantum algorithms (quantum correlation between the problem-setting and the corresponding solution), quantum nonlocality (quantum correlation between the outcomes of measuring two maximally entangled observables in a state of spatial separation of the respective subsystems), and a quantum cosmological interpretation of the Anthropic Principle (quantum correlation between the value of the fundamental constants since the beginning of time and the “value” of life today, in particular “sentient life” or “no life”).
The rationale is divided into two parts.
The first provides the concept of the time-symmetric quantum causal loops, showing that they stand at the physical basis of the quantum computational speedup and that, consequently, quantum algorithms respond to the notion of teleological evolution. This naturally raises the question whether the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithms could also be the missing physical basis of a teleological character of the natural evolutions for which the teleological notion was conceived.
The second part of the rationale proposes an affirmative answer to this question. It is, of course, a speculation. It will imply seeing quantum features – the quantum speedup – in evolutions that can be the size of the Earth and last billions of years. We will show how this becomes possible under a quantum cosmological interpretation of the Anthropic Principle.
2.1 The time-symmetric quantum causal loops – the concept
We go now to the first part of the work. We begin by showing, by the way very simply, that the sheer existence of the optimal quantum speedup delivered by Grover $\left[9\right]$ quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops.
We start with the simplest instance of the quantum algorithm. The problem to be solved is as follows. Bob, the problem-setter, hides a ball in a chest of four drawers; the number of the drawer in which to hide the ball – the problem-setting – is selected at random. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open from one to three drawers. If the ball is only in the third drawer opened, she has solved the problem; if not even in that, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well. In the quantum case, with the optimal quantum search algorithm of Grover, she always locates the ball by opening just one drawer, in a quantum superposition of the four possible drawer numbers. Always solving the problem by opening just one drawer would obviously be impossible in the classical case.
The fact that Alice always solves the problem of locating the ball by opening just one drawer logically implies that it is as if she always knew in advance, before beginning her problem-solving action, that the ball is in one of the three pairs of drawers formed by pairing the drawer with the ball with another drawer.
Indeed, in any of these three mutually exclusive cases, she opens either drawer; if the ball is in it, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well.
The question is then: from where does Alice always get (in a mutually exclusive way) this advanced knowledge of one of the possible halves of the information that specifies the solution? Since, to her, the only possible source of it is the solution she will produce and measure in the future, the only possible answer is that it comes to her from her future measurement of the solution. This is, of course, a causal loop that allows her to produce the solution by opening just one drawer. There must be, in a mutually exclusive way, one causal loop for each of the three possible ways of knowing in advance one of the possible halves of the information that specifies the solution.
We consider now $N$ drawers. We should use to Long’s $\left[10\right]$ version of Grover algorithm that always provides the solution with certainty. Grover algorithm (also in Long’s version) yields a quadratic quantum speedup, that is, always solves the problem by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers against the $\operatorname*{O}\left(N\right)$ of the best classical case. Also now, this quantum speedup is logically justified by the fact that Alice knows in advance, in a mutually exclusive way, one of the possible halves of the information about the number of the drawer with the ball. This means that she knows in advance (in a mutually exclusive way) that the ball is in one of the possible tuples of $\sqrt{N}$ drawers. She then solves the problem, in the best possible logical way, by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers, which is indeed what Grover algorithm does.
We can therefore formulate the following advanced knowledge rule:
The number of computation steps (drawer openings) needed to solve Grover’s problem in an optimal quantum way is that needed to solve it in an optimal logical way given the advanced knowledge of one of the possible halves of the information about the solution.
Of course, involving only the measurements of commuting observables (the problem-setting and the corresponding solution), Grover algorithm, as any other quantum algorithm, is submitted to the same (classical) logic. Its quantum description should therefore describe the quantum superposition of the causal loops in question (logical mutual exclusivity of possibilities becomes their quantum superposition in the quantum world).
However, the above immediately raises a problem. The current quantum description of quantum algorithms, as that of any quantum process, is causal in character. In other words, causality always goes in the same time direction. Obviously, it is a description that cannot describe causal loops; we ascribe this to its incompleteness and show that it is completed by time-symmetrizing it.
First, we should emend a trivial incompleteness of the usual quantum description of quantum algorithms. It is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It is obviously incomplete because it lacks the initial measurement. Then, we extend it to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition (or, indifferently, mixture) of all possible problem-settings. In the case of Grover algorithm, where the solution is an invertible function of the problem-setting, we obtain a unitary evolution (encoding the process of solving the problem) between an initial and a final one-to-one correlated measurement outcomes (encoding respectively the problem-setting selected by the initial measurement and the corresponding solution read by the final measurement). Of course, this unitary evolution never changes the problem-setting selected by the initial measurement.
In view of going beyond Grover algorithm, we note that the one-to-one correlation between the two measurement outcomes is no restriction: if the solution is not an invertible function of the problem-setting, to have this correlation it suffices that, at the end, the problem-solver measures both the solution and the problem setting. The only constraint is that the quantum algorithm determines the solution with certainty.
We call the above quantum description of the quantum algorithm its ordinary quantum description, ordinary because it is still a causal description that cannot describe causal loops. It is the “less incomplete” quantum description that we should complete by the time-symmetrizing it.
In order that it generates the causal loops logically implied by the sheer existence of the quantum speedup, the time-symmetrization must be as follows.
Instead of ascribing the selection of the pair of one-to-one correlated measurement outcomes all to the initial measurement, as customary, we require that it time-symmetrically shares between the initial and final measurements, of course in a quantum superposition of all the possible ways. In each time-symmetrization instance, the initial measurement selects one of the possible halves of the information that specifies the sorted out pair of correlated measurement outcomes (it selects a quantum superposition of problem-settings instead of the customary single one). Its measurement outcome unitarily propagates forward in time until becoming the state immediately before the final measurement, a quantum superposition of tensor products each a problem-setting multiplying the corresponding solution. The final measurement completes the selection, of course by selecting a single pair problem-setting and corresponding solution. Its measurement outcome, obviously the same of the ordinary quantum description, time-symmetrically propagates backward in time, by the inverse of the time-forward unitary transformation, until replacing the previous initial measurement outcome.
This backward in time propagation, which inherits both selections, is an instance of the time-symmetrization of the ordinary quantum description of the quantum algorithm. Note that, mathematically, i.e. disregarding the time-direction of causality, it is its ordinary quantum description back again.
Also note that, together, the forward and backward in time propagations describe a causal loop that changes its own past (we call it time-symmetric since it is generated by a time-symmetrization procedure). As seen by the problem-solver (see Subsection 4.2.), from whom the outcome of the initial measurement should always be concealed (would give her information about the solution), it is one of the causal loops logically implied by the sheer existence of the optimal quantum speedup.
Eventually, we should take the quantum superposition of all the causal loops.
As we can see, this time-symmetrization leaves the unitary evolution between the two measurement outcomes mathematically unaltered, but changes the structure of causality along it. As seen by the problem-solver, the single causal process of the ordinary quantum description is replaced by the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup. As already said, in each causal loop it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future. Then, she can use this knowledge to produce the solution with fewer computation steps. When she exploits it entirely, the quantum speedup is optimal.
In equivalent terms, in each causal loop, the computational complexity of the problem to be solved reduces – this will become tangible in the formalizations of Section 4. The number of computation steps needed to solve the reduced complexity problem in an optimal quantum way is that needed to solve it in an optimal logical way. In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place correspondingly reduces.
In view of what will follow, it is important to note that these results have been obtained independently of the distinctive character of the unitary evolution between the two correlated measurement outcomes (i.e. the actual quantum algorithm). It sufficed to know that it is a unitary evolution that one-to-one correlates two initially unrelated observables (the problem-setting observable and the solution observable) never changing the eigenvalue of the observable selected by the initial measurement.
In the present work, we will always know that there is such a unitary evolution. It is Grover algorithm in the present case, but it could be any other quantum algorithm that provides the solution with certainty – see also $\left[7\right]$. Under the assumption that between two one-to-one correlated measurement outcomes there can always be a unitary evolution of the kind in question, the advanced knowledge rule holds for any oracle problem.
Eventually, we note that the time-symmetrization in question, and the consequent time-symmetric quantum causal loops, are imposed by logic and quantum mechanics. Summarizing, logic says that the sheer existence of an optimal quantum speedup logically implies the mutually exclusive $or$ of well defined causal loops. Of course, these causal loop must appear in the quantum description of the quantum algorithm that delivers the quantum speedup, which is submitted to the same (classical) logic. In turn, quantum mechanics says that the only way of obtaining this is by time-symmetrizing the ordinary quantum description of the quantum algorithm. This, in turn, generates the description of the quantum superposition of causal loops in question. Therefore, as long as logic and quantum mechanics are consistent, these time-symmetric quantum causal loops should be exempt from the paradoxes that plague the usual kind.
Interestingly, the time-symmetrization we are dealing with turns out to be mandatory also to satisfy a symmetry implicit in any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement.
In fact, in this kind of process, we can postpone the initial measurement at will – all results must remain unaltered. When the two measurements become simultaneous, the time-symmetrization in question must already be there to become an atemporal symmetrization mandatory to represent the perfect symmetry between the simultaneous measurements of two maximally entangled observables.
This further motivation for time-symmetrizing the unitary evolutions we are dealing with, besides further justifying the previous motivation, makes things much more general. Any such unitary evolution must be time-symmetrized, what leaves it mathematically unaltered but changes its ordinary, causal, quantum description into that a quantum superposition of time-symmetric causal loops.
If the unitary evolution is a quantum algorithm that determines the solution with certainty, in each causal loop it is as if the problem solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future. However, this just creates the room for its occurring with quantum speedup; the quantum algorithm does not need any more to be optimal like in the case of Grover algorithm. How much of this room is exploited depends on the problem-solver, i.e. on the quantum algorithm.
If the unitary evolution in question is not a quantum algorithm, but anyhow creates one-to-one quantum correlation between two initially unrelated observables (without changing the eigenvalue of the observable selected by the initial measurement), it can always occur with quantum speedup. In this more general case, the problem-solver advanced knowledge might loose any meaning but, of course, can be replaced by the reduction of the dimension of the Hilbert space in which the search for one-to-one correlation takes place. This reduction, in turn, creates the room for a quantum speedup.
There is also the case that the unitary evolution does not build up any quantum correlation. It is the case of quantum nonlocality, namely the measurement in the same basis of two maximally entangled observables whose respective subsystems are in a state of spatial separation. Here maximal entanglement between the two observables is already there before the space-separation process. After the time-symmetrization, this process becomes a quantum superposition of time-symmetric quantum causal loops locally going from one to the other measurement outcome via the time the two subsystems were not yet spatially separate. Here, always locally, these causal loops tell either observable the state to assume in view of its future measurement – see $\left[8\right]$. These causal loops play the role of the hidden variables envisaged by Einstein and the others $\left[11\right]$ to locally (i.e. also physically) describe quantum nonlocality.
Note that, as a consequence, the incompleteness we are currently ascribing to the quantum description of quantum algorithms is the same that Einstein and the others ascribed to the quantum description of quantum nonlocality. In both cases, what is incomplete is the quantum description of a quantum correlation. With this, we end the first part of the work.
2.2 Teleological character of natural evolutions – a possible physical basis
The second part of the work examines the possibility that the time-symmetric quantum causal loops underlying the teleological character of quantum algorithms are also the missing physical basis of the teleological character of the natural evolutions for which the teleological notion was conceived.
This, of course, immediately raises the problem of explaining the existence of quantum features – indeed the quantum speedup – in evolutions that can be the size of the Earth and have a duration of billions of years. The obvious impossibility of scaling up the quantum features of the microscopic world to such levels suggests the possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe. Naturally, at the same time we should adopt quantum cosmology $\left[12,13\right]$, namely the assumption that the evolution of the universe, a perfectly isolated system since by definition there is nothing outside it, is quantum and unitary in character.
An immediate consequence is having to also adopt a quantum cosmological interpretation of the Anthropic Principle $\left[14-18\right]$, the “fine-tuned universe” version of it. This version of the principle asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental physical constants since the beginning of time and the “value” of life today, in particular “sentient life” or “no life”. Let us assume, for simplicity, that this correlation today is one-to-one. The significance of this principle lies in the fact that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
Under a quantum cosmological interpretation of it, the correlation addressed by the Anthropic Principle becomes quantum. We should think of a value of the fundamental constants observable and a value of life observable. At the beginning, there should be a quantum superposition (or mixture) of all the possible universes, each with its value of the fundamental constants. An initial measurement of the value of the fundamental constants observable would select one value at random. The subsequent unitary evolution of the universe would produce, at the present time, the corresponding value of life, which would then be measured. In Subsection 5.1., after the formalization of the notion of the time-symmetric quantum causal loops, we will provide a formalized toy model of the building up of the correlation in question on the part of the universe’s evolution.
By the way, since the value of the fundamental constants never changes along the universe’s evolution, in the universe we are in we can do without the initial measurement (when there were no observers). Today, there would be a quantum superposition of tensor products, each a possible value of the fundamental constants multiplying the corresponding value of life. The measurement of either observable today would select the eigenvalues of both observables. All measurement outcomes are as if the value of the fundamental constants was selected at the beginning of time.
Summing up, we are dealing with a unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue selected by the initial measurement (here the value of the fundamental constants). We have seen that this unitary evolution can always occur with quantum speedup. Being a quantum search (for one-to-one correlation), it should occur with quantum speedup even if the search (by the universe’s evolution under the fundamental physical laws and constants) is random $\left[19\right]$. It should consequently be teleological in character.
In the universe we are in, the search for one-to-one correlation between the value of the fundamental constants and the value of life becomes search for sentient life. It is thus the universe’s evolution toward sentient life that occurs with quantum speedup and is consequently teleological in character.
In principle, the evolution of life on Earth can be obtained from that of the universe toward sentient life by tracing out all the rest. Evolution times would be identical and the amount of the resources employed by the evolution of life on Earth could not exceed those employed by its image in the evolution of the universe that comprises them. In the end, the evolution of life on Earth, even if decohered, would occur with quantum speedup and would thus be teleological in character. This should also hold for the evolutions of species, etc..
Of course, all the above is puzzling.
A physical basis to the dismissed notion of teleological evolution would be a paradigm shift for the natural sciences. As is well known, empirically, the teleological notion better fits with the functional structure and the evolutions of the living. If it has been discarded, it is only because of the assumed absence of a physical basis. For the time being, the fact that quantum algorithms are certainly teleological evolutions, and that, under a quantum cosmological interpretation of the Anthropic Principle, natural evolutions would be teleological in character too, may shake the current belief that such a physical basis cannot exist.
That quantum algorithms are teleological evolutions might even be acceptable from the psychological standpoint: it would be the strangeness of the microscopic quantum world and its short lived evolutions soon killed by decoherence. Things become quite different in the case of the evolutions of the living. The fact that what happens around and within us would be driven not only by the past but also by a future even billions of years ahead is more than puzzling. It might go along with ancient beliefs, but of a kind deliberately ignored by science. Once science was accused of heresy for denying a physical basis to beliefs of the kind in question. We could be facing the reverse problem now: proposing a physical basis to the notion of teleological evolution could well be seen as a scientific heresy.
3 Positioning the work
We try and position the present work. Of course, we are not alone in challenging the predominant idea that causality only goes forward in time. Besides the retrocausal approach to the quantum computational speedup $\left[4-8\right]$, we should mention time-symmetric quantum mechanics $\left[20-24\right]$, in particular its study of the possibility that a future choice changes a past quantum measurement outcome $\left[23\right]$, and ongoing attempts to study quantum processes without the need to preassume a definite causal structure $\left[25-28\right]$.
The present approach is naturally inspired by time-symmetric quantum mechanics and has in common with the attempts in question the fact of being open to the violation of the ordinary notion of causality, not in common the way of looking at the quantum description of quantum algorithms. For example, in $\left[27\right]$, quantum algorithms are considered examples of processes with a definite causal structure because, in their current quantum description, causality always goes forward in time. So far so good but, consequently, they are excluded (by the approach in question) from the phenomena that can accommodate causality violations. Of course, for us, this means putting too much faith in their current quantum description. As we have seen, with their quantum computational speedups, quantum algorithms should be a most clear example of violation of one-way causality – what obviously implies changing their current quantum description.
Some concepts of the work have precedents.
One is the local explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[29\right]$. Consider the measurements in the same basis of two maximally entangled observables in a state of spatial separation of the respective subsystems. Costa de Beauregard assumed that the outcome of the first of the two measurements locally propagates backward in time, by the inverse of the unitary transformation that spatially separated the two subsystems, until these were not yet spatially separate. Here, still locally, it would tell the other observable the state to assume in view of its future measurement. This retrocausal, local, explanation of quantum nonlocality is similar to ours. The difference is that sending backward in time an entire measurement outcome can originate paradoxes and, of course, is not imposed by quantum mechanics.
Interestingly, at the time, Costa de Beauregard was an assistant of De Broglie, who forbade him to publish that work for years until the publication, by Wheeler and Feynman, of their emitter-absorber theory that also resorted to retrocausality – see $\left[30\right]$ for an account of this. This tells us a lot on the physicist’s position toward retrocausality, even at the top levels. By the way, despite the importance of the author’s name, Costa de Beauregard’s explanation of quantum nonlocality never took off.
Other precedents are Wheeler’s Participatory Anthropic Principle $\left[31,32\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[12\right]$. These two principles share with the present work the notion of mutual causality between the observation today and the value of the fundamental physical laws and constants since the beginning of time mediated by a quantum evolution of the universe in between. Dating from before Deutsch’s 1985 discovery, the two principles in question make no reference to the quantum speedup of the universe’s evolution and, a-fortiori, its consequent teleological character.
4 The time-symmetric quantum causal loops – formalization
We formalize the notion of the time-symmetric quantum causal loops on the four drawer instance of Grover $\left[9\right]$quantum search algorithm. As anticipated in the Rationale, this immediately generalizes first to any number of drawers (under Long’s $\left[10\right]$ version of Grover algorithm that always provides the solution with certainty), then to any unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement. The formalization consists of the following steps:
(1) Extend the usual quantum description of the quantum algorithm, limited to the process of solving the problem, to the upstream process of setting the problem. We call this extended quantum description the ordinary quantum description of the quantum algorithm, ordinary because it is still a causal description that cannot describe causal loops.
(2) This point is propaedeutic to point (4). It shows how to relativize with respect to the problem-solver the ordinary quantum description of the quantum algorithm, which is with respect to the customary external observer. Of course, the problem-solver must be concealed from the problem-setting selected by the initial measurement, which would give her information about the solution of the problem.
(3) Complete the ordinary quantum description of the quantum algorithm by time-symmetrizing it. Mathematically, the quantum description of the process between the two correlated measurement outcomes remains unaltered. However, the description of the behavior of causality along it changes: the single causal process of the ordinary quantum description is replaced by a quantum superposition of time-symmetric causal loops.
(4) Relativize the quantum description of the causal loops generated at point (3) with respect to the problem-solver. This produces the quantum description of the quantum superposition of causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
4.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
The usual quantum description of quantum algorithms is limited to the process of solving the problem, where the quantum speedup resides, and lacks the initial measurement. Of course, it is trivially incomplete: the canonical quantum description of a quantum process must consist of initial measurement, unitary evolution, and final measurement. We extend the usual quantum description to the upstream process of setting the problem. An initial measurement of the problem-setting, performed in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings, selects one of them at random. This is followed by the unitary evolution that computes the corresponding solution and the final measurement of the solution.
Let us introduce the notation. We make reference to the simplest four drawer instance of Grover algorithm:
We number the four draws in binary notation: $00,01,10,11$.
We will need two quantum registers of two quantum bits each:
A register $B$ under the control of the problem-setter Bob contains the problem-setting (the number of the drawer in which the ball is hidden). The content of register $B$ is the problem-setting observable $\hat{B}$ of eigenstates and corresponding eigenvalues:
$|00\rangle_{B}$ and $00$; $|01\rangle_{B}$ and $01$; $|10\rangle_{B}$ and $10$; $|11\rangle_{B}$ and $11$.
A register $A$ under the control of the problem-solver Alice contains the number of the drawer opened by her. The content of register $A$ is the observable $\hat{A}$ of eigenstates and corresponding eigenvalues:
$|00\rangle_{A}$ and $00$; $|01\rangle_{A}$ and $01$; $|10\rangle_{A}$ and $10$; $|11\rangle_{A}$ and $11$;
we call it content of $A$ observable. Measuring $\hat{A}$ at the end of the quantum algorithm yields the number of the drawer with the ball selected by the initial measurement, namely the solution of the problem. Note that $\hat{B}$ and $\hat{A}$ commute.
The quantum description of the extended quantum algorithm, namely its ordinary quantum description is:
time $\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, meas. of}\;\hat{B}\end{array}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\text{, meas. of}\;\hat{A}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&&\\ \Downarrow&&\\ |01\rangle_{B}\left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table I)
Here and in the following we disregard normalization.
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow333 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operantion that changes nothing for us. .
It is now Alice’s turn. The unitary part of her problem-solving action, $\hat{\mathbf{U}}_{1,2}$, produces the solution of the problem, of course, never changing the problem-setting. Note that, of $\hat{\mathbf{U}}_{1,2}$, it is sufficient to know that it does this, whereas we will never need to know how it does it. Of course, we must also know that there exists a unitary transformation that does this, Grover algorithm in the present case.
At the end of Alice’s problem-solving action, at time $t_{2}$, register $A$ contains the solution of the problem, i.e. the number of the drawer with the ball – bottom-right corner of the diagram.
Eventually, by measuring $\hat{A}$ (the content of $A$), which is already in one of its eigenstates, Alice acquires the solution with probability $1$ – the quantum state in the bottom-right corner of the diagram naturally remains unaltered throughout this measurement. Correspondingly, there is a unitary transformation between the initial and final measurement outcomes; the process between them is reversible despite comprising the final measurement.
4.2 Relativizing the ordinary quantum description of the quantum algorithm with respect to the problem-solver
The ordinary quantum description of the quantum algorithm of Table I is with respect to the customary external observer, namely an observer external to the quantum process. It is not a description for the problem-solver (Alice). In fact, it would tell her, immediately after the initial measurement of $\hat{B}$, the number of the drawer with the ball selected by Bob, which is also the solution of the problem – see the state of register $B$ in the bottom-left corner of Table I diagram. Of course, in the quantum description of the quantum algorithm with respect to Alice, she must be concealed from the outcome of the initial measurement444 More in general, one should assume that the problem-solver has a complete knowledge of the problem, thus also of the relation between problem-setting and solution. Knowledge of the problem-setting selected by the initial measurement would tell her also the solution. .
Since Alice’s problem-solving action never changes the number of the drawer with the ball selected by Bob and the observables $\hat{B}$ and $\hat{A}$ commute, describing this concealment is simple. It suffices to postpone after the end of Alice’s problem-solving action the projection of the quantum state associated with the initial measurement of $\hat{B}$ (or, indifferently, to postpone the very measurement of $\hat{B}$). The result is the following description of the quantum algorithm with respect to Alice:
time $\begin{array}[]{ccc}\begin{array}[]{c}\text{time}\text{ }t_{1}\text{, meas. of}\;\hat{B}\end{array}&t_{1}\rightarrow t_{2}&\text{\text{time }\text{}}t_{2}\text{, meas. of}\;\hat{A}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&\begin{array}[]{c}|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}+\\ |10\rangle_{B}|10\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}\\ &&\Downarrow\\ &&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table II)
To Alice, the initial measurement of $\hat{B}$ leaves the initial state of register $B$ unaltered. At the beginning of her problem-solving action she is completely ignorant of the number of the drawer with the ball selected by Bob – see the state of register $B$ on the left of $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$.
The unitary transformation $\hat{\mathbf{U}}_{1,2}$ is then performed for a quantum superposition of the four possible numbers of the drawer with the ball. The state at the end of it – on the right of the horizontal arrows – is a quantum superposition of tensor products, each a possible number of the drawer with the ball in register $B$ tensor product the corresponding solution (that same number) in register $A$. Since $\hat{B}$ and $\hat{A}$ commute, the final measurement of $\hat{A}$ must project this superposition on the tensor product of the number of the drawer with the ball initially selected by Bob and the corresponding solution – bottom-right corner of Table II diagram.
Note that this relativization relies on relational quantum mechanics $\left[33,34\right]$, where the quantum state is relative to the observer. In the following, we will continue to call the quantum description with respect to the external observer the ordinary quantum description of the quantum algorithm. Instead we will always specify that the other quantum description is that with respect to the problem-solver.
4.3 Time-symmetrizing the ordinary quantum description of the quantum algorithm
As we have already said, in order to obtain the quantum superposition of the causal loops whose mutually exclusive $or$ is implied by the sheer existence of an optimal quantum speedup, we should first time-symmetrize the ordinary quantum description of the quantum algorithm, then relativize with respect to the problem-solver the causal loops obtained.
To start with, we should require that the selection of the information that specifies the sorted out pair of correlated measurement outcomes among all the possible pairs time-symmetrically shares between the initial and final measurements. The half information selected by the initial measurement should propagate forward in time, by $\hat{\mathbf{U}}_{1,2}$ until becoming the state immediately before the final measurement. Time-symmetrically, the complementary selection performed by the final measurement should propagate backward in time, by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$, until replacing the previous outcome of the initial measurement.
We consider the time-symmetrization instance where the initial measurement of $\hat{B}$ reduces to the measurement of $\hat{B}_{l}$, the left bit of the two-bit number contained in register $B$; say that it randomly selects left bit 00. The final measurement of $\hat{A}$ should correspondingly reduce to that of $\hat{A_{r}}$ the right bit of the two-bit number contained in register $A$; say that it randomly selects right bit $1$. The randomly selected number of the drawer with the ball is thus $01$.
In this particular time-symmetrization instance, the ordinary quantum description of the quantum algorithm with respect to the external observer of Table I becomes:
time $\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, }\text{}\text{meas. of}\;\hat{B}_{l}\end{array}&t_{1\rightleftarrows}t_{2}&\text{time }\text{}t_{2}\text{, meas. of}\;\hat{A_{r}}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&&\\ \Downarrow&&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}\\ &&\Downarrow\\ |01\rangle_{B}\left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)&\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table III)
The initial measurement of $\hat{B}_{l}$, selecting the 00 of $01$, projects the initial quantum superposition of all the basis vectors of register $B$ on the superposition of those beginning with 00 – vertical arrow on the left of the diagram. Under $\hat{\mathbf{U}}_{1,2}$, the latter superposition causally propagates forward in time into the superposition of the two tensor products on the right of the right looking horizontal arrows. Then the final measurement of $\hat{A_{r}}$, selecting the $1$ of $01$, projects the latter superposition on the term ending with $1$ under the right vertical arrow. Note that this outcome of the final measurement, which inherits both selections, coincides with that of the ordinary quantum description – see the bottom-right corner of Table I diagram. Time-symmetrically, this final measurement outcome causally propagates backward in time by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until becoming the definitive outcome of the initial measurement. This backward in time propagation (which inherits both selections) is an instance of the time-symmetrized description of the quantum algorithm – bottom line of the diagram. Note that, since the symbol “$\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$” is mathematically equivalent to “$\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$”, this instance, up to the time-direction of causality, is the unitary process between the two one-to-one correlated measurement outcomes back again.
Also note that causality performs a causal loop that changes its own past; the outcome of its initial measurement of $\hat{B}_{l}$ is changed into the outcome of the initial measurement of $\hat{B}$ of the ordinary quantum description; compare the state of register $B$ under the left vertical arrow with the state in the bottom-left corner of the diagram.
We should eventually take the quantum superposition of all the time-symmetrization instances. We can see that, no matter how we evenly share the selection between initial and final measurements, the final measurement outcome, which inherits both selections, is always that of the final measurement of $\hat{A}$ of the ordinary quantum description of the quantum algorithm. Correspondingly, its subsequent backward in time propagation by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ (bottom line of the diagram), mathematically (disregarding the direction of causality), is always the ordinary quantum description of the unitary process between the two correlated measurement outcomes, namely the bottom line of Table I diagram.
Summing up, mathematically, time-symmetrization leaves the unitary process between the two measurement outcomes unaltered but changes the structure of causality along it. The causal process of the ordinary quantum description is replaced by a quantum superposition of causal loops.
4.4 Relativizing the causal loops with respect to the problem-solver
The causal loops generated by time-symmetrizing the ordinary quantum description of the quantum algorithm, which is relative to the external observer, do not yet map on those logically implied by the sheer existence of the quantum speedup. This is because the latter causal loops, concerning the problem-solver advanced knowledge of one of the halves of the information about the solution, must of course be relative to the problem-solver.
Therefore, we should relativize with respect to the problem-solver the causal loop of Table III by postponing after her problem-solving action the projection of the quantum state due to the initial measurement of $\hat{B}_{l}$ (whose outcome should be concealed from Alice). This yields the causal loop in the following table:
time $\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, meas. of}\;\hat{B}_{l}\end{array}&t_{1\rightleftarrows}t_{2}&\text{time }t_{2}\text{,meas. of}\;\hat{A_{r}}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&\begin{array}[]{c}|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}+\\ |10\rangle_{B}|10\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}\\ &&\Downarrow\\ \begin{array}[]{c}\left(|01\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table IV)
The projection of the initial state of register $B$ due to the measurement of $\hat{B}_{l}$ – top-left corner of the diagram – is postponed after Alice’s problem-solving action, namely outside the present table that is limited to that action. With this postponement, the initial state of register $B$ goes unaltered through the initial measurement of $\hat{B}_{l}$ becoming, by $\hat{\mathbf{U}}_{1,2}$, the quantum superposition of tensor products above the vertical arrow. Alice’s final measurement of $\hat{A_{r}}$, which in present assumptions selects the value $1$ of the right bit of the number contained in register $A$, projects the state above the vertical arrow on the state below it. Then (in causal order), this final measurement outcome propagates backward in time by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until becoming the definitive outcome of the initial measurement. This backward in time propagation is an instance of the time-symmetrized description of the quantum algorithm with respect to the problem-solver.
Note that we would have obtained the identical diagram by time-symmetrizing the description of the quantum algorithm relativized to the problem-solver of Table II.
Also note that, for the mathematical equivalence between $\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$ and $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$, the bottom line of the diagram of Table IV can also be read in the usual left to right way:
time $\begin{array}[]{ccc}\text{time }t_{1}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\\ &&\\ \begin{array}[]{c}(|01\rangle_{B}+|11\rangle_{B})\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table V)
Note that the entire causal loop is the zigzag diagram of table IV. It is a causal loop that changes its own past. The initial state in the top-left part of the diagram, where the state of register $B$ tells that Alice is completely ignorant of the number of the drawer with the ball, changes into the initial state in the bottom-left part of the diagram, where the state of register $B$ tells that Alice knows in advance, before beginning her problem-solving action, that the ball is in drawer either $01$ or $11$. Since what happens along the time-symmetric causal loop obeys classical logic, the number of drawer openings needed to find the solution in the optimal quantum case is that needed in the optimal logical way, namely just one.
In other words, the computational complexity of the problem to be solved by her reduces. The problem of locating the ball in one of $4$ drawers – as for the state of register $B$ in the upper-left corner of the diagram of Table IV – becomes that of locating it in one of $2=\sqrt{4}$ drawers – as for the state of register $B$ in the bottom-left corner of Table IV diagram (or the left of Table V diagram). In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place reduces from $4$ basis vectors to $2=\sqrt{4}$.
In equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place quadratically reduces.
The generalization of the above is straightforward. Note that, in all our considerations, we never needed to know the unitary transformation between the two measurement outcomes $\hat{\mathbf{U}}_{1,2}$. It sufficed to know that it unitarily changes its input state, where the two observables ($\hat{B}$ and $\hat{A}$ ) are unrelated, into its output state, where they are one-to-one correlated, without changing the eigenvalue of the observable selected by the initial measurement.
First, the causal loops explanation of the quantum speedup, and the fact that the latter is quadratic in kind, remain there if we increase the number of quantum bits in the two registers of an even number and in an equal way. In fact, all our considerations remain identical. This means, in particular, Grover algorithm with number of drawers any even power of $2$.
The same, of course, holds, for any unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement. This will apply to the teleological evolutions addressed in the next section.
5 The physical basis of the notion of teleological evolution
We have seen, rather simply, that the complete quantum description of quantum algorithms is a quantum superposition of time-symmetric causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. In the optimal quantum case, the number of computation steps is that needed to solve the problem in the optimal logical way benefiting of the advanced knowledge in question. In equivalent terms, the computational complexity of the problem to be solved or the dimension of the Hilbert space in which the search for the solution takes place correspondingly reduces. As already observed, in quantum search algorithms, also a random quantum search for the solution is endowed with (a suboptimal) quantum speedup.
Therefore, we can say that quantum algorithms are evolutions toward a goal (the solution of the problem) with an attractor in the goal they will produce in the future. Of course, this is necessarily a causal loop and responds to the notion of teleological evolution. As is well known, before the scientific revolution, the evolutions of the living were often considered to be teleological in character. The teleological notion was then dismissed with the advent of modern science for the alleged absence of a physical basis. That was, of course, under the causal view of evolutions that is dominant in science also today.
Now, the fact that quantum algorithms turn out to be teleological evolutions should already cast a doubt on this way of seeing. It also, naturally, raises the question whether the time-symmetric quantum causal loops that stand at the basis of their teleological character could be the missing physical basis of the teleological character of natural evolutions.
As already seen in the Rationale, under a quantum cosmological interpretation of the Anthropic Principle, the search for the value of life by the evolution of the universe under the fundamental physical laws and constants has the same form a quantum search algorithm. This search can therefore occur with quantum speedup and does even if it is random. Consequently, the evolution of a universe toward the value of life turns out to be teleological in character. In the universe we are in, what turns out to be teleological in character is the evolution of the universe toward sentient life.
Eventually, the evolution of life on earth, in principle obtainable from that of the universe by tracing out all the rest, even if decohered, would inherit the quantum speedup and the consequent teleological character of the evolution of the universe toward life.
In the end, the notion that the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithm are also the physical basis of the teleological character of natural evolutions would seem to be physically plausible, this, of course, within a quantum cosmological view.
This section is a discussion of the many impacts that the present teleological view of natural evolutions might have on different areas of science and knowledge; of course, not forgetting the natural philosophy that gave rise to the notion of teleological evolution.
5.1 Decohered evolutions of life endowed with quantum speedup: how would they look?
All natural evolutions are necessarily described in a probabilistic way. We can think of the stochastic description of a Darwinian evolution: an alternation of random changes in the species DNA pool each followed by natural selection of the DNA changes more suited to the environment. Of course, in a classical stochastic process, transition probabilities can depend on the past of the process but not on the future of it.
Instead, a decohered process endowed with quantum speedup must have an attractor in its future. This means that those DNA mutations that in the future will survive natural selection must make, in the present, the transition toward them more probable. By the way, since the consequent teleological character of the evolution would be hidden in transition probabilities, the process would be difficult to distinguish from a classical stochastic process. Let us call time-symmetric also the kind of stochastic evolution that we have just outlined.
We mention what could be an emblematic example of time-symmetric stochastic evolution. It is known that, for many millions of years, the ancestors of the flying birds were feathered dinosaurs who did not fly. Very light, strong but flexible, cave, and air-proof, feathers are of course ideal for flight. So, why anticipating something designed for flight millions of years in advance?
The usual explanation is of course a causal one: at first, feathers appeared at random and survived natural selection as good thermal insulators, later they proved useful for speeding escape from predators, perhaps for gliding from trees to the ground, and only eventually they turned out to be what was needed to conquer the air.
According to the present teleological view of the evolutions of the living, even during the flightless period of the feathered dinosaurs, those random DNA mutations that would have led in the future to the conquest of the air – obviously an extremely successful outcome from the point of view of survival – would have more likely occurred.
The above is puzzling, of course. We should keep in our mind that we are under the quantum cosmological assumption that might make macroscopic evolutions, the partial trace over all the rest of a quantum evolution of the universe, as strange as the microscopic quantum world.
5.2 Teleological evolution of the universe toward life – a toy model
A toy model of a teleological evolution of the universe toward life can be trivially mapped on the simplest instance of Grover algorithm.
We should imagine a value of the fundamental physical constants observable $\hat{B}$ and a value of life observable $\hat{A}$. Say that the eigenvalue $01$ of $\hat{B}$ is the value of the fundamental physical constants compatible with the development of sentient life and the eigenvalues $00,10,11$ the values of those incompatible. We assume that, correspondingly, the eigenvalue $01$ of the measurement of $\hat{A}$ at the present time would mean existence of sentient life and the eigenvalues $00,10,11$ absence of life.
Initially, $\hat{B}$ should be in a quantum superposition (or, indifferently, mixture) of all its eigenstates (i.e of all the possible universes each with its fundamental physical constants). We can assume that also $\hat{A}$, the value of life observable, is initially in a quantum superposition of all its eigenstates.
The universe’s unitary evolution under the fundamental physical laws and constants would have finally, at the present time, built up a quantum superposition of tensor products, each an eigenstate of $\hat{B}$ tensor product the corresponding eigenstate of $\hat{A}$ – see table II. In our universe, at the present time one with observers, the measurement of $\hat{A}$ at the present time would select the eigenvalue $01$ of $\hat{A}$, namely “existence of sentient life” and the corresponding eigenvalue $01$ of the value of the fundamental physical constants observable $\hat{B}$. Since the universe’s evolution does not change the value of the fundamental physical constants, all is as if the eigenvalue of $\hat{B}$ was selected by a measurement of $\hat{B}$ at the beginning of time, when there could be no observers.
Of course, this is exactly the quantum search algorithm described in the previous section, the version with respect to the problem-solver of Table II. As shown by its time-symmetrization (Table IV), it can undergo with a quantum speedup, and does even if the quantum search for the value of sentient life is random.
5.3 How to model a teleological natural evolution
We address the problem of how to model a decohered evolution endowed with quantum speedup – say that of the living on Earth after the beginning of the liquid water stage of it. The canonical way would be to obtain this evolution from that of the universe toward life by tracing out all the rest. Of course, in the present case, this holds only in principle. However, it would be interesting to apply the canonical way to a toy model of the situation we are dealing with.
One should identify the purely quantum evolution of a system (the universe) endowed with quantum speedup, where the achievement of the goal is mostly localized in a subsystem thereof (the Earth). Then, one should obtain a decohered evolution of the subsystem from that of the system by tracing out all the rest. Note that the interaction of the subsystem with the rest of the system, in the present case, would do the opposite than decoherence. It would be what allows the evolution of the subsystem toward the goal to occur with quantum speedup as part of the purely quantum evolution of the system.
Let us note that, since the liquid water stage of it, the evolution of life on Earth should have been almost completely independent from the evolution of the rest of the universe. If it were completely independent, it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup. Therefore, an obvious, approximate, way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact it is decohered and model it in a quantum way. There are encouraging precedents to this. For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[35\right]$.
As already said, a further possibility should be to model a decohered evolution endowed with quantum speedup as a stochastic process with transition probabilities time-symmetrically depending on both the past and the future of the process, something just outlined for the time being.
5.4 Impact on the natural sciences
The current position of the natural sciences toward teleology can be considered in some measure conflicting. On the one hand, the teleological view would seem to allow a deeper understanding of the functional structure and the evolutions of living beings. On the other, as this vision is considered without physical basis and therefore misleading, there is outright hostility to it.
This tension is captured by the following sentence by the evolutionary biologist Haldane $\left[36\right]$:
Teleology is like a mistress to a biologist: he cannot live without her, but he is unwilling to be seen with her in public.
It would seem that, today, no evolutionary biologist supports the teleological view, some explicitly assert that it is an unworkable view because it implies retrocausality.
No biologist, of course, does not mean no scientist. A paradigmatic example is the American philosopher Nagel $\left[37\right]$, who argues that the principles that explain the emergence of life should be teleological, rather than materialistic or mechanistic in character.
We would put together Nagel’s view in support of the teleological character of the evolutions of the living and Haldane’s assertion that a biologist cannot live without teleology to affirm that the notion of teleological natural evolution is more “scientific” than the mainstream one of causal natural evolution. This in the sense that it better answers empirical evidence – of course the first canon of science. The only reason biologists would be against teleology in spite of empirical evidence would come from another science. We mean the notion – current in today’s physics – that a teleological evolution is physically impossible.
Of course, the sheer notion that quantum computation is teleological in character may already put a strain on the position of today’s biology: actually, although in the quantum world, teleological evolutions are physically possible. The strain should increase with the notion that, under a quantum cosmological interpretation of the Anthropic Principle, the evolution of the universe toward life should be teleological in character. It should further increase with the consequent notion that also the evolution of life on Earth, obtainable in principle from that of the universe by tracing out all the rest, even if decohered, should inherit the teleological character of the universe’s evolution.
Summing up, given the above, the impact of the present teleological view of natural evolutions on today’s evolutionary biology would seem to be almost a head on collision.
Almost because, on the one side, in recent years a group of evolutionary biologists $\left[38\right]$ expressed the need to rethink Darwin’s theory of evolution. That was after the discovery that some species’ evolution steps occurred far too quickly to be solely ascribed to random changes in the species’ DNA pool. Of course, the present teleological view of evolutions, with the associated quantum speedup, would offer a way out.
On the other side, there are scientists who have began to apply essentially quantum mechanical features like quantum entanglement and nonlocality to the macroscopic game of species, showing that this can explain some empirically ascertained features of it otherwise difficult to explain $\left[35\right]$.
5.5 Impact on philosophy
Of course, the current theory of teleological evolutions rehabilitates an ancient notion of natural philosophy $\left[39\right]$. We should therefore be allowed to look for the possibility that there are other relations between this theory and philosophy. As we are going to see, their abundance is a surprising interdisciplinary aspect of the present work.
In effect, the teleological theory in question seems to revamp a full body of ancient philosophical views sort of opposite (and complementary) to the body of philosophical views emerged with the scientific revolution.
Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view555 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops. and the causal view.
The causal view and the mutual causality view seem to have respectively subtended two opposing views throughout the history of philosophy. The causal view would underlie the materialist, mechanistic and reductionist views, while the mutual causality view would underlie the idealist, teleological and monist views respectively.
Consider the opposition between materialism and idealism. The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[31\right]$, the anthropic cosmological principles of Barrow and Tipler $\left[12\right]$, and the present theory is a mutual causality view.
We find a similar situation in the opposition between the mechanistic and teleological views of evolutions, which obviously respond to the causal and mutual causality views of evolutions, respectively.
For what concerns reductionism vs monism, we note that deriving the teleological character of the decohered evolutions of the living systems on Earth from that of the evolution of a quantum universe is an essentially monistic view. It presupposes that what we see around and inside is essentially inseparable from the whole.
Let us mention other theories in the history of philosophy reminiscent of the mutual causality between sentient life today and the physical laws with their physical constants since the beginning of time:
(i) For Plato, “Ideas [of our mind] are objective perfections that exist in themselves and for themselves and, at the same time, are the causes of natural phenomena and constitute their unity” $\left[40\right]$. In modern language, the cause of natural phenomena that constitute their unity are the fundamental physical laws. The identity between the fundamental physical laws since the beginning of time and the ideas of our mind today is formally reminiscent of the identity (and mutual causality) between the number of the drawer with the ball (the problem-setting) randomly selected by Bob’s measurement at the beginning of the quantum algorithm and the number of the drawer with the ball (the solution of the problem) selected by Alice’s measurement at the end of it. It is as if the number of the drawer with the ball was determined by both the problem-setter and the problem-solver.
Also Plato’s notion that ideas – physical laws – are objective perfections could have a correspondence with the fine tuned universe version of the Anthropic Principle, when it asserts that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
The above should also go along with Fritjov Capra’s $\left[41\right]$ notion that there is identity between the fundamental states of consciousness described by eastern theosophies and the fundamental laws of quantum physics.
(ii) We do not need to be anthropocentric. This mutual causality would also be between physical laws and spatial intelligence, often superior to our in non-human animals. The assumption that spatial intelligence is achieved with quantum speedup excludes that it is only the result of “ordinary” natural evolution (obeying the principle of causality). Indeed a quantum speedup in the development of spatial intelligence would imply the mutual causality in question.
(iii) The present notion of teleological evolution, namely of an evolution toward a goal with an attractor in the goal it will reach in the future, could also be seen in terms of a “systematically lucky” evolution, lucky from the standpoint of the objective of reaching its goal, indeed because attracted by it. Interestingly, also the notion of a systematically lucky evolution seems to have a precedent in pre-scientific thinking. It would be the Stoic notion of $\pi\rho o\nu o\iota\alpha$, foreknowledge $\left[42\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge. Foreknowledge, or advanced knowledge, would have pushed the evolution toward its goal in a luckier and thus faster way than what is possible in the case of a causal evolution, exactly as the advanced knowledge about the solution does in quantum algorithms. This notion of $\pi\rho o\nu o\iota\alpha$ would also fit well with the notion of teleological evolution understood as a stochastic processes with transition probabilities time-symmetrically affected by the past and future of the process, with the luckier transitions more probable than in the classical case.
(iv) It is natural to wander whether the present time-symmetric quantum causal loops view of the universe’s evolution toward sentient life, where the sentient life stage is an attractor of the evolution of the universe toward it, could be related to Nietzsche’s intuition of an eternal return consequent to our will for it.
(v) Michelangelo’s “Creation of Adam” on the ceiling of the Sistine Chapel could be the manifesto of the present theory. There is a mirror symmetry between creator and created $\left[43\right]$. Could this be interpreted as the time-symmetry – and the related mutual causation – between the fundamental physical laws with their physical constants since the beginning of time and the existence of sentient life today? Furthermore, the fact that the figures and shapes portrayed behind the creator figure would seem to be an anatomically accurate portrait of the human brain $\left[44\right]$, could it allude to the teleological character of the evolution of our intelligent species: the human brain being an attractor of the search for it? We would not feel like dismissing these analogies as sheer coincidences. As a matter of fact, in idealistic philosophy – in Schelling – aesthetic insight would reproduce the fundamental mechanisms of the production of life.
6 Research prospects
Future work on the subject of the present work could have a broad spectrum.
For what concerns quantum physics:
(i) Look for a synthesis between the present notion of the time-symmetric quantum causal loops and time-symmetric quantum mechanics $\left[20-24\right]$ to which it is obviously inspired.
(ii) Look for toy models that capture the mechanism for which the evolution of life on Earth (a subsystem), although decohered, inherits the quantum speedup and consequently the teleological character of the evolution of the universe (the system) toward life.
(iii) Explore the possibility that other quantum features suspected to be involved in the fundamental processes of life derive top-down from a quantum character of the universe (as, presently, the teleological character of the evolutions of the living).
(iv) Look for the possible impact of the teleological view on quantum cosmology. According to current assumptions, even the inanimate part of it, anyway an evolution toward life, should be endowed with quantum speedup and thus teleological in character.
(v) The Anthropic Principle is limited to stating the correlation between the “value” of life and the value of the fundamental physical constants since the beginning of time. One could also look for a correlation between the value of life and the character of the fundamental physical laws since the beginning of time. Of course, also these would have an attractor in the goal of creating life.
About the natural sciences, looking for:
(I) A reconsideration of the ban on applying the teleological view and language to natural evolutions. Restoring the dismissed notion of teleological evolution would obviously be a paradigm shift for today’s natural sciences.
(II) Toy models of teleological Darwinian processes. These evolutions could be either approximated by pure quantum processes endowed with quantum speedup or as decohered stochastic processes with transition probabilities time-symmetrically affected by the past and the future of the process.
(III) An application of the time-symmetric quantum causal loops to quantum self-organization. An example is studying the teleological formation of a self-replicating system. There seem to be an enormous jump of complexity in the transition from a soup of life-precursor molecules into a self-replicating system. The idea that this transition has an attractor in the full fledged self-replicating system can therefore be attractive.
Finally, opposing schools of thought throughout the history of philosophy could be reviewed in the light of the opposition between the causal and mutual causality views, as exemplified in the previous subsection.
7 Conclusions
The work is, of course, unconventional. It might be useful to summarize in a consequential way the line of reasoning followed.
We have (trivially) shown that the sheer existence of the optimal quantum speedup of Grover quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and used this knowledge to produce the solution with correspondingly fewer computation steps. The number of computation steps needed to solve the problem in an optimal quantum way is that needed to solve it in an optimal logical way using the advanced knowledge in question.
Since quantum algorithms only involve the measurements of commuting observables, they are submitted to the same (classical) logic. Therefore, Grover algorithm must host the quantum superposition of the causal loops in question.
However, being causal in character as that of any other quantum process, its customary quantum description cannot describe causal loops; we ascribe this to its incompleteness.
Before going to the way of remedying this incompleteness, we must eliminate another, trivial, one. The customary quantum description of quantum algorithms is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It is incomplete since it lacks the initial measurement.
Thus, as a first step, we extend the customary quantum description of the process of solving the problem to the upstream process of setting the problem, by adding the initial measurement of the problem-setting in a quantum superposition (or mixture) of all the possible problem-settings. We have thus a unitary evolution between two one-to-one correlated measurement outcomes that does not change the problem-setting selected by the initial measurement. This is a key feature of quantum algorithms, not visible in their customary quantum description666 If the solution is not an invertible function of the problem-setting, to have one-to-one correlation between the two measurement outcomes, it suffices that the problem-solver, at the end, measures both the solution and the problem-setting. . We call this kind of process the ordinary quantum description of the quantum algorithm, ordinary because it is still a causal description that cannot describe causal loops.
The second step is to change the causal structure of the ordinary quantum description of the quantum algorithm by time-symmetrizing the latter. We time-symmetrically share between the initial and final measurements the selection of the information that specifies the sorted out pair of correlated measurement outcomes, in a quantum superposition of all the possible ways.
Each time-symmetrization instance, mathematically, yields the unitary evolution between the two measurement outcomes back again, but the structure of causality along it changes. It becomes a causal loop that changes its own past. Relativized with respect to the problem-solver (who should be concealed from any outcome of the initial measurement), the causal loop in question becomes one of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
Eventually, we should take the quantum superposition of all the causal loops.
Until now, the time-symmetrization in question has been imposed by the necessity of reproducing, in the quantum description of Grover algorithm, the causal loops logically implied by the sheer existence of its (optimal) quantum speedup. However, as we have seen, the same time-symmetrization is also redundantly imposed by a symmetry implicit in any unitary evolution between an initial and a final, one-to-one correlated, measurement outcomes that does not change the eigenvalue selected by the initial measurement. As a consequence, any such evolution is a quantum superposition of time-symmetric causal loops.
If the unitary evolution builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement, like in the case of Grover algorithm, in each time-symmetrization instance the dimension of the Hilbert space in which the search for one-to-one correlation takes place reduces. This creates the room for a quantum speedup. There should be a quantum speedup even in the case that the search for one-to-one correlation is random.
The process is a quantum superposition of time-symmetric causal loops also in the case that the two observables are already maximally entangled at the beginning of it. It is the case of quantum nonlocality, where the quantum process starts with the spatial separation between two subsystems whose respective observables are already maximally entangled. Here the causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space-separate, where they locally tell either observable the state to assume in view of its future measurement. They play the role of the hidden variables envisaged by Einstein and the others to locally explain quantum nonlocality.
The quantum causal loop description of quantum algorithms highlights their teleological character, namely their being evolutions toward a goal (the solution) with an attractor in the goal they will produce in the future (the solution again). This naturally raises the question whether the time-symmetric quantum causal loops we are dealing with could also be the physical basis of the teleological character of the evolutions of the living for which the notion of teleological evolution was conceived.
This immediately raises the problem of explaining a quantum speedup in evolutions of the living that can be the size of the Earth and have durations of billions of years. The fact it is obviously impossible to scale up the quantum features of the microscopic world to such levels suggests the possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe. Of course, at the same time we should adopt quantum cosmology, namely the assumption that the universe evolves in a unitary quantum way.
We should now call into play the “fine-tuned universe” version of the Anthropic Principle. It asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value of the fundamental physical constants since the beginning of time and the “value” of life today, in particular, either “existence of sentient life” or “absence of life”. We assume for simplicity that this correlation is one-to-one
With the evolution of the universe quantum, also the correlation asserted by the Anthropic Principle becomes quantum. We are thus dealing with a unitary evolution that one-to-one correlates two initially unrelated observables (the value of the fundamental constants and the value of life observables) without changing the eigenvalue of the observable selected by the initial measurement. We have seen that this kind of unitary evolution can occur with quantum speedup and does so even in the case that the search for one-to-one correlation is random.
The evolution of the universe toward the value of life should therefore take place with quantum speedup and thus be teleological in character. In the universe we are in, this would apply to the evolution of the universe toward sentient life.
We have also seen that the evolution of life on Earth, in principle obtainable from that of the universe toward sentient life by tracing out all the rest, should inherit the teleological character of the latter evolution. In fact: (i) evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe and (ii) the resources employed by the former evolution cannot exceed those employed by an image that comprises them.
In conclusion, the notion of the time-symmetric quantum causal loops turns out to be the common physical basis of such diverse physical phenomena as the quantum computational speedup, the consequent teleological character of quantum algorithms, quantum nonlocality, and – under a quantum cosmological interpretation of the Anthropic Principle – the teleological character of the evolutions of the living.
Of course, it is a puzzling notion that radically changes a long-standing way of viewing the role of causality in the physical world.
Regarding quantum phenomena in the microscopic world, we think to have provided clear evidence that the notion of the time-symmetric quantum causal loops is a necessary consequence of the notion of quantum computational speedup.
Regarding the teleological character of the macroscopic evolutions of life, the situation becomes even more puzzling. We limit ourselves to noticing that this is a necessary consequence of the quantum cosmological assumption. Seeing the macroscopic evolutions of life as the partial trace over everything else of a quantum evolution of the universe toward life brings into the macroscopic world of everyday experience the strange character of the microscopic quantum world.
Acknowledgments
Thanks are due to Eliahu Cohen, Artur Ekert, Avshalom Elitzur and David Finkelstein for useful comments throughout the whole of the present approach to the quantum speedup, Daniel Sheehan for organizing the San Diego AAAS Pacific Division conferences on retrocausality, a far-looking forum for the discussion of frontier, also unorthodox, physics, and Mario Rasetti for organizing the first series of international workshops on quantum communication and computation, the Elsag Bailey-ISI Turin Villa Gualino workshops of the years 1991-96.

### v4

@@
 We show that, under a quantum cosmological interpretation of the Anthropic Principle, the same quantum superposition of causal loops underlying the teleological character of quantum algorithms becomes the missing physical basis of the teleological character of natural evolutions.
-1 Foreword
-The present work is about causal loops, of course a puzzling notion: in our natural way of seeing the world, we firmly believe that cause always precedes effect in time.
-In a causal loop, the effect, in turn, changes backward in time its previous cause.
-This may seem paradoxical; it is in fact the source of well-known paradoxes such as the grandfather one.
-So, let us say right away that the causal loops we will be dealing with are quantum and of a special time-symmetric kind that should be imposed by quantum mechanics.
-Therefore, as long as the latter is consistent, they should be exempt from the paradoxes that plague the usual kind.
-They would be the physical reason for the quantum computational speedup, the fact that solving a problem in the quantum world can require fewer logical operations (computation steps) than in the classical world222 The notion that computation is possible in the quantum world, with the related notion of quantum bit, was first introduced by Finkelstein $[1]$ in 1969.
-In 1982, Feynman $\left[2\right]$ observed that the classical simulation of a quantum process can require an exponentially higher amount of physical resources.
-In 1985, Deutsch $\left[3\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup.
+1 Introduction
+The notion of teleological evolution is that of an evolution toward a goal with an attractor in the same goal it will produce in the future.
+It is, of course, a puzzling notion since it necessarily implies a causal loop.
+As is well known, it was often applied to the evolutions of the living before the scientific revolution.
+With the advent of the latter, it was dismissed because of its alleged lack of physical basis.
+That was, of course, under the notion of causal evolution that is dominant in science even today.
+We take the opposite position in the present work. We will show that, in quantum mechanics, there is a well defined physical basis to the notion of teleological evolution.
+In the first place, we show that quantum algorithms are teleological evolutions, indeed evolutions toward a goal (the solution) with an attractor in the goal they will produce in the future (the solution again).
+Of course, seeing this requires a quantum description of quantum algorithms different from the ordinary one, which is naturally that of a causal process.
+The latter quantum description, with causality always going in the same time-direction, cannot describe causal loops.
+We show that the ordinary quantum description of quantum algorithms, although complete from the mathematical standpoint, is physically incomplete and is completed by time-symmetrizing it.
+This leaves the ordinary quantum description mathematically unaltered but replaces its single causal process by a quantum superposition of causal loops.
+In each of them, it is as if the problem-solver knew in advance, before beginning her problem-solving action, one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer logical operations (computation steps).
+Of course, we are dealing with an evolution that responds to the notion of teleological evolution: it is indeed toward a goal (the solution) with an attractor in the goal it will produce in the future (the solution again).
+This completion of the quantum description provides a unified quantitative explanation of all the optimal quantum speedups achieved in solving oracle problems.
+Given an oracle problem, the number of computation steps needed to solve it in an optimal quantum way is that needed to solve it in a optimal logical way benefiting of the advanced knowledge of any of the possible halves of the information about the solution.
+Of course, there is no such an explanation in current quantum computing science.
+Obviously, we have a causal quantum description for each quantum algorithm, but there is no known way of unifying them.
+This part of the work is a further development of the retrocausal approach to the quantum computational speedup $\left[1-5\right]$.
+Of course, both the approach in question and the part of the work discussed so far are unconventional and potentially controversial.
+The notion of causal loop is puzzling and mostly extraneous to today’s physics.
+However, we believe that this time we have given a very simple and easily checkable demonstration that the quantum computational speedup is physically due to quantum causal loops.
+Let us outline it.
+We consider the following problem.
+Bob, the problem-setter, hides a ball in a chest of four drawers – we call the number of the drawer with the ball the problem-setting.
+Alice, the problem-solver, is to locate the ball by opening drawers.
+In the classical case, she has to open up to three drawers.
+If the ball is only in the third drawer opened, she has solved the problem.
+If it is not even in that drawer, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well.
+In the quantum case – by the optimal quantum search algorithm of Grover $\left[6\right]$ – she always solves the problem by opening just one drawer222 In the four drawers case, Grover algorithm yields the solution with certainty.
+In the following, by Grover algorithm we will always mean Long’s $\left[7\right]$ version of it, which does that always.
 .
-More in general, they would always be present in quantum correlations.
-The present work is divided into two parts.
-The first is a review and further clarification of the retrocausal approach to the quantum computational speedup $\left[4-8\right]$.
-It concerns a (time-symmetric) quantum causal loops description of quantum algorithms that provides a unified quantitative explanation of their quantum speedups.
-According to it, quantum algorithms are quantum superpositions of causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future.
-Then, she could use this knowledge to produce the solution with fewer computation steps.
-In equivalent terms, the computational complexity of the problem to be solved, or the dimension of the Hilbert space in which the search for the solution takes place, correspondingly reduces.
-The number of computation steps needed to solve the problem in an optimal quantum way is the number needed to solve it in an optimal logical way benefiting from the advanced knowledge in question.
-We call this the advanced knowledge rule.
-The first part of the work ends with the observation that, under their time-symmetric quantum causal loops description, quantum algorithms turn out to be teleological evolutions, namely evolutions toward a goal (the solution of the problem) with an attractor in the goal they will produce in the future (the solution again) – a causal loop in fact.
-As is well known, the evolutions of the living were often thought to be teleological in character.
-The teleological notion was dismissed with the advent of modern science for the alleged absence of a physical basis.
-That was, of course, under the notion of causal evolution that is dominant in science also today and that the present work challenges.
-The second part of the work is about the possibility that the time-symmetric quantum causal loops at the physical basis of the teleological character of quantum algorithms are also the physical basis of the teleological character of the natural evolutions for which the notion of teleological evolution was conceived.
-We will see that, under a quantum cosmological interpretation of the Anthropic Principle, this possibility turns out to be plausible.
+This would obviously be impossible in the classical case, it seems also impossible to our intuitive way of viewing it.
+It should already be clear from the above that that of the quantum speedup is a fundamental discovery that changes a long standing way of viewing things333 The notion that computation is possible in the quantum world, with the related notion of quantum bit, was introduced by Finkelstein $[8]$ in 1969.
+In 1982, Feynman $\left[9\right]$ observed that the classical simulation of a quantum process can require an exponentially higher amount of physical resources.
+In 1985, Deutsch $\left[10\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup.
+.
+We can go now to our demonstration, which is surprisingly simple. Of course, we must adopt a way of thinking open to retrocausality.
+The fact that Alice always solves the problem of locating the ball by opening just one drawer is logically equivalent to her knowing in advance, before beginning her problem-solving action, that the ball is in a given pair of drawers, in a mutually exclusive way one of the three pairs obtained by pairing the drawer with the ball with another drawer.
+In fact, if the ball is in the drawer opened, she has solved the problem.
+If not, it must be in the other drawer and she has solved the problem of locating the ball as well, having opened just one drawer.
+Then the question becomes: from where does Alice always get the information that the ball is in a given pair of drawers?
+Since, to her, the only possible source of it is the solution of the problem she will produce and measure in the future, the only possible answer is that it comes to her from her future measurement of the solution.
+This is obviously a causal loop.
+Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution by opening correspondingly fewer drawers.
+This, in the mutually exclusive $or$ of all the possible ways of taking half of the information about the solution.
+We would argue that, with the above elementary logic, we have, so to speak, broken the sound barrier: the quantum speedup does logically imply causal loops.
+Then, one gos from logic to physics as follows.
+Involving only the measurements of commuting observables (the problem-setting and the corresponding solution), quantum algorithms are submitted to the same (classical) logic.
+Therefore they must host the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of an optimal quantum speedup; of course, logical mutual exclusivity of possibilities, moving to the quantum world, becomes their quantum superposition.
+We have already seen that the quantum description of this quantum superposition is generated by time-symmetrizing the ordinary quantum description of the quantum algorithm.
+This time-symmetrization must therefore be performed because, as we will see, is the only way of generating the causal loops that must be hosted in the quantum algorithm.
+Let us add a few comments to the above.
+By ordinary quantum description of quantum algorithms, we do not mean their usual quantum description.
+The latter is limited to the unitary process that solves the problem, where the quantum speedup resides, and the final measurement of the solution.
+It is trivially incomplete since it lacks the initial measurement.
+Since we need to complete the quantum description, in the first place we should emend this trivial incompleteness.
+We do this by extending the usual quantum description to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition (or, indifferently, mixture) of all possible problem-settings.
+In the case of Grover algorithm, we obtain a unitary evolution between an initial and a final one-to-one correlated measurement outcomes encoding respectively the problem-setting (the number of the drawer with the ball selected by the initial measurement) and the corresponding solution of the problem.
+We have called it the ordinary quantum description of the quantum algorithm, ordinary since it is still a causal description that cannot describe causal loops.
+Note that any quantum algorithm that produces the solution with certainty can be put in this form.
+If the solution is not an invertible function of the problem-setting, it suffices that Alice, at the end, measures both the solution and the problem setting.
+This is a trivial but important step: of course, the usual quantum description of quantum algorithms would completely obscure the approach we are following.
+Let us eventually note that we are dealing with the same kind of incompleteness of the ordinary quantum description that was ascribed by Einstein, Podolsky and Rosen $\left[11\right]$ to the quantum description of quantum nonlocality.
+Also this quantum description can be seen as a unitary evolution (the process of spatial separations of the two subsystems) between two one-to-one correlated measurement outcomes (those of measuring in the same basis the two respective maximally entangled observables) – we must think that this process goes from one to the other measurement outcome via the time the two subsystems were not spatially separate.
+Time symmetrized, it generates a quantum superposition of causal loops that play the role of the hidden variables envisaged by those authors to physically (locally) explain quantum non-locality $\left[5\right]$.
+This ends the first part of the work, a review and further clarification of the retrocausal approach to the quantum speedup.
+Of course, under the above explanation of their quantum speedup, quantum algorithms turn out to be teleological evolutions.
+The second part of the work is about the possibility that the quantum causal loops that stand at the basis of the teleological character of quantum algorithms are the missing physical basis of the teleological character of the evolutions of the living.
+This immediately raises two problems.
+One is explaining the presence of quantum features (the quantum speedup) in evolutions that can be the size of the Earth and have durations of billions of years.
+The obvious impossibility of scaling up the quantum features of the microscopic world to such levels points to the only possible solution: to scale them down from the largest and longest existing evolution there is, that of the universe.
+Of course, at the same time, we should adopt quantum cosmology, that is, the assumption that the universe – a perfectly isolated system since by definition there is nothing outside of it – evolves in a unitary quantum way.
+The other is to find a formal analogy between quantum algorithms and, under quantum cosmology, natural evolutions.
+Obviously, it should justify the fact that also natural evolutions can occur with a quantum speedup and are consequently teleological in character.
+Interestingly, the Anthropic Principle comes to our rescue.
+The Fine-tuned Universe version of it asserts that the unitary evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental constants since the beginning of time and the “value” of life today, in particular (“sentient life” or “no life”).
+The reason for this correlation is that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
+Having adopted quantum cosmology, of course, also implies a quantum interpretation of the Anthropic Principle.
+Simply, the correlation between the value of the physical constants since the beginning of time and the value of life today becomes quantum and its building up is performed by the universe’s unitary quantum evolution.
+This is what we were looking for.
+In fact, it is exactly what quantum algorithms do.
+They unitarily build up a quantum correlation between the problem-setting since the beginning of them and the solution of the problem at their end.
+Under this formal analogy, the evolution of the universe toward the value of life could be endowed with quantum speedup and would even in the case that the quantum search for the correlation in question is random, exactly as in random quantum search algorithms $\left[12\right]$.
+Of course, being endowed with quantum speedup, under the quantum causal loops explanation of it, the evolution in question is teleological in character.
+In the universe we are in, what becomes teleological in character is the evolution of the universe toward sentient life.
+Eventually, as we shall see, the evolution of life on Earth, in principle obtainable from that of the universe toward life by tracing out everything else, even if decohered, should inherit the quantum speed and thus the teleological character of the latter evolution.
+The same would consequently hold for any evolution of life on Earth, e.
+g.
+that of species.
+The notion of a teleological evolution of life is of course puzzling.
+For example, the Darwinian theory of evolution would change as as follow: those random DNA mutations that, in the future, will survive natural selection would become more probable.
+With the past influenced by the future, this is a completely new way of viewing the physical world.
+No wonder the teleological notion has been rejected by science.
+In conclusion, the work is highly unconventional, but its findings were obtained in a surprisingly simple and well-verifiable way.
+If it were correct, as we think, it would be a paradigm shift not only for the natural sciences, but also for the way we see physical reality.
 2 Rationale
-As the present work implies a fundamentally different way of seeing the role of causality in quantum correlation, we replace the usual introduction by a more extended rationale that describes this way of seeing in sufficient detail.
-Quantum correlation is of course the common feature of quantum algorithms (quantum correlation between the problem-setting and the corresponding solution), quantum nonlocality (quantum correlation between the outcomes of measuring two maximally entangled observables in a state of spatial separation of the respective subsystems), and a quantum cosmological interpretation of the Anthropic Principle (quantum correlation between the value of the fundamental constants since the beginning of time and the “value” of life today, in particular “sentient life” or “no life”).
-The rationale is divided into two parts.
-The first provides the concept of the time-symmetric quantum causal loops, showing that they stand at the physical basis of the quantum computational speedup and that, consequently, quantum algorithms respond to the notion of teleological evolution.
-This naturally raises the question whether the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithms could also be the missing physical basis of a teleological character of the natural evolutions for which the teleological notion was conceived.
-The second part of the rationale proposes an affirmative answer to this question.
-It is, of course, a speculation.
-It will imply seeing quantum features – the quantum speedup – in evolutions that can be the size of the Earth and last billions of years.
-We will show how this becomes possible under a quantum cosmological interpretation of the Anthropic Principle.
+As the present work is about a fundamentally different way of viewing the role of causality in quantum (correlation) processes, before going to the formalizations we provide a Rationale that describes it conceptually.
+This Rationale is divided into two parts corresponding to the two parts in which the work is divided.
+The first provides the concept of the time-symmetric quantum causal loops, showing how they stand at the physical basis of the quantum computational speedup and quantum nonlocality.
+In particular, the time-symmetric quantum causal loop description of quantum algorithms highlights their teleological character, their being evolutions toward a goal (the solution of the problem) with an attractor in the goal they will produce in the future (the solution of the problem again).
+The second addresses the question whether the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithms could also be the missing physical basis of a teleological character of the natural evolutions for which the teleological notion was conceived.
 2.1 The time-symmetric quantum causal loops – the concept
 We go now to the first part of the work.
-We begin by showing, by the way very simply, that the sheer existence of the optimal quantum speedup delivered by Grover $\left[9\right]$ quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops.
-We start with the simplest instance of the quantum algorithm.
-The problem to be solved is as follows.
-Bob, the problem-setter, hides a ball in a chest of four drawers; the number of the drawer in which to hide the ball – the problem-setting – is selected at random.
-Alice, the problem-solver, is to locate the ball by opening drawers.
-In the classical case, she has to open from one to three drawers.
-If the ball is only in the third drawer opened, she has solved the problem; if not even in that, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well.
-In the quantum case, with the optimal quantum search algorithm of Grover, she always locates the ball by opening just one drawer, in a quantum superposition of the four possible drawer numbers.
-Always solving the problem by opening just one drawer would obviously be impossible in the classical case.
-The fact that Alice always solves the problem of locating the ball by opening just one drawer logically implies that it is as if she always knew in advance, before beginning her problem-solving action, that the ball is in one of the three pairs of drawers formed by pairing the drawer with the ball with another drawer.
+We begin by showing, by the way very simply, that the sheer existence of the optimal quantum speedup delivered by Grover $\left[6\right]$ quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops.
+Of course, we must ignore the ban on the use of the notion of retrocausality.
+The fact that Alice (the problem-solver) always solves the problem of locating the ball by opening just one drawer logically implies that it is as if she always knew in advance, before beginning her problem-solving action, that the ball is in one of the three pairs of drawers formed by pairing the drawer with the ball with another drawer.
 Indeed, in any of these three mutually exclusive cases, she opens either drawer; if the ball is in it, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well.
@@
 We consider now $N$ drawers.
-We should use to Long’s $\left[10\right]$ version of Grover algorithm that always provides the solution with certainty.
 Grover algorithm (also in Long’s version) yields a quadratic quantum speedup, that is, always solves the problem by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers against the $\operatorname*{O}\left(N\right)$ of the best classical case.
 Also now, this quantum speedup is logically justified by the fact that Alice knows in advance, in a mutually exclusive way, one of the possible halves of the information about the number of the drawer with the ball.
-This means that she knows in advance (in a mutually exclusive way) that the ball is in one of the possible tuples of $\sqrt{N}$ drawers.
+This means that she knows in advance that the ball is in one of the possible tuples of $\sqrt{N}$ drawers.
 She then solves the problem, in the best possible logical way, by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers, which is indeed what Grover algorithm does.
@@
 In other words, causality always goes in the same time direction.
-Obviously, it is a description that cannot describe causal loops; we ascribe this to its incompleteness and show that it is completed by time-symmetrizing it.
-First, we should emend a trivial incompleteness of the usual quantum description of quantum algorithms.
+Obviously, it is a description that cannot describe causal loops; we tentatively ascribe this to its incompleteness.
+Before going to this assumed incompleteness, of a fundamental kind, we should emend a trivial incompleteness of the current quantum description of quantum algorithms.
 It is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution.
@@
 Then, we extend it to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition (or, indifferently, mixture) of all possible problem-settings.
-In the case of Grover algorithm, where the solution is an invertible function of the problem-setting, we obtain a unitary evolution (encoding the process of solving the problem) between an initial and a final one-to-one correlated measurement outcomes (encoding respectively the problem-setting selected by the initial measurement and the corresponding solution read by the final measurement).
+In the case of Grover algorithm, where the solution is an invertible function of the problem-setting (both are the number of the drawer with the ball), we obtain a unitary evolution (encoding the process of solving the problem) between an initial and a final one-to-one correlated measurement outcomes (encoding respectively the problem-setting selected by the initial measurement and the corresponding solution read by the final measurement).
 Of course, this unitary evolution never changes the problem-setting selected by the initial measurement.
-In view of going beyond Grover algorithm, we note that the one-to-one correlation between the two measurement outcomes is no restriction: if the solution is not an invertible function of the problem-setting, to have this correlation it suffices that, at the end, the problem-solver measures both the solution and the problem setting.
+By the way, seeing quantum algorithms as unitary evolutions between an initial and a final one-to-one correlated measurement outcomes is a key step of the present work.
+In view of going beyond Grover algorithm, we note that, if the solution is not an invertible function of the problem-setting, we can have one-to-one correlation between the initial and final measurement outcomes as well by requiring that, at the end, the problem-solver measures both the solution and the problem setting.
 The only constraint is that the quantum algorithm determines the solution with certainty.
-We call the above quantum description of the quantum algorithm its ordinary quantum description, ordinary because it is still a causal description that cannot describe causal loops.
-It is the “less incomplete” quantum description that we should complete by the time-symmetrizing it.
-In order that it generates the causal loops logically implied by the sheer existence of the quantum speedup, the time-symmetrization must be as follows.
+We call the above, less incomplete, quantum description of the quantum algorithm its ordinary quantum description, ordinary because it is still a causal description that cannot describe causal loops.
+We should complete this ordinary quantum description in a way that leaves it mathematically unaltered but changes its causality structure – describes the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup.
+This is simply done time-symmetrizing it as follows.
 Instead of ascribing the selection of the pair of one-to-one correlated measurement outcomes all to the initial measurement, as customary, we require that it time-symmetrically shares between the initial and final measurements, of course in a quantum superposition of all the possible ways.
@@
 Note that, mathematically, i.e.
-disregarding the time-direction of causality, it is its ordinary quantum description back again.
+disregarding the time-direction of causality, it is the ordinary quantum description back again.
 Also note that, together, the forward and backward in time propagations describe a causal loop that changes its own past (we call it time-symmetric since it is generated by a time-symmetrization procedure).
-As seen by the problem-solver (see Subsection 4.2.), from whom the outcome of the initial measurement should always be concealed (would give her information about the solution), it is one of the causal loops logically implied by the sheer existence of the optimal quantum speedup.
+As seen by the problem-solver, from whom the outcome of the initial measurement should always be concealed (would give her information about the solution – see Subsection 4.2.), it is one of the causal loops logically implied by the sheer existence of the optimal quantum speedup.
 Eventually, we should take the quantum superposition of all the causal loops.
-As we can see, this time-symmetrization leaves the unitary evolution between the two measurement outcomes mathematically unaltered, but changes the structure of causality along it.
+Of course, the backward in time propagations of all time-symmetrization instances are identical with each other.
+Therefore, as required, this time-symmetrization leaves the unitary evolution between the two measurement outcomes mathematically unaltered.
+However, it changes the structure of causality along it.
 As seen by the problem-solver, the single causal process of the ordinary quantum description is replaced by the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
-As already said, in each causal loop it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future.
-Then, she can use this knowledge to produce the solution with fewer computation steps.
-When she exploits it entirely, the quantum speedup is optimal.
+As already said, in each causal loop it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future; then, she can use this knowledge to produce the solution with fewer computation steps.
+If she exploits it entirely, the quantum speedup is optimal.
 In equivalent terms, in each causal loop, the computational complexity of the problem to be solved reduces – this will become tangible in the formalizations of Section 4.
@@
 In the present work, we will always know that there is such a unitary evolution.
-It is Grover algorithm in the present case, but it could be any other quantum algorithm that provides the solution with certainty – see also $\left[7\right]$.
+It is Grover algorithm in the present case, but it could be any other quantum algorithm that provides the solution with certainty – see also $\left[4\right]$.
 Under the assumption that between two one-to-one correlated measurement outcomes there can always be a unitary evolution of the kind in question, the advanced knowledge rule holds for any oracle problem.
@@
 Summarizing, logic says that the sheer existence of an optimal quantum speedup logically implies the mutually exclusive $or$ of well defined causal loops.
-Of course, these causal loop must appear in the quantum description of the quantum algorithm that delivers the quantum speedup, which is submitted to the same (classical) logic.
-In turn, quantum mechanics says that the only way of obtaining this is by time-symmetrizing the ordinary quantum description of the quantum algorithm.
-This, in turn, generates the description of the quantum superposition of causal loops in question.
-Therefore, as long as logic and quantum mechanics are consistent, these time-symmetric quantum causal loops should be exempt from the paradoxes that plague the usual kind.
+Of course, these causal loop must appear in the quantum description of the quantum algorithm that delivers the quantum speedup that, involving only the measurements of commuting observables, is submitted to the same (classical) logic.
+In turn, it is easily seen that the only way of obtaining these causal loops is by time-symmetrizing the ordinary quantum description of the quantum algorithm.
+In fact, to have a causal loop, we must properly share the selection of the pair problem-setting and corresponding solution between the initial and final measurements.
+Because of the time-symmetry of the unitary evolution between the two one-to-one correlated measurement outcomes, we should share it evenly.
+This exactly specifies the time-symmetrization we are dealing with.
+Therefore, as long as logic and quantum mechanics are consistent, the time-symmetric quantum causal loops generated by the time-symmetrization operation should be exempt from the paradoxes that plague the usual kind.
 Interestingly, the time-symmetrization we are dealing with turns out to be mandatory also to satisfy a symmetry implicit in any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement.
 In fact, in this kind of process, we can postpone the initial measurement at will – all results must remain unaltered.
-When the two measurements become simultaneous, the time-symmetrization in question must already be there to become an atemporal symmetrization mandatory to represent the perfect symmetry between the simultaneous measurements of two maximally entangled observables.
+When the two measurements become simultaneous, the time-symmetrization in question must already be there to become the atemporal symmetrization mandatory to represent the perfect symmetry between the simultaneous measurements of two maximally entangled observables.
 This further motivation for time-symmetrizing the unitary evolutions we are dealing with, besides further justifying the previous motivation, makes things much more general.
@@
 After the time-symmetrization, this process becomes a quantum superposition of time-symmetric quantum causal loops locally going from one to the other measurement outcome via the time the two subsystems were not yet spatially separate.
-Here, always locally, these causal loops tell either observable the state to assume in view of its future measurement – see $\left[8\right]$.
-These causal loops play the role of the hidden variables envisaged by Einstein and the others $\left[11\right]$ to locally (i.e.
+Here, always locally, these causal loops tell either observable the state to assume in view of its future measurement – see $\left[5\right]$.
+We can see that these causal loops play the role of the hidden variables envisaged by Einstein and the others $\left[11\right]$ to locally (i.e.
 also physically) describe quantum nonlocality.
@@
 The obvious impossibility of scaling up the quantum features of the microscopic world to such levels suggests the possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe.
-Naturally, at the same time we should adopt quantum cosmology $\left[12,13\right]$, namely the assumption that the evolution of the universe, a perfectly isolated system since by definition there is nothing outside it, is quantum and unitary in character.
-An immediate consequence is having to also adopt a quantum cosmological interpretation of the Anthropic Principle $\left[14-18\right]$, the “fine-tuned universe” version of it.
+Naturally, at the same time we should adopt quantum cosmology $\left[13,14\right]$, namely the assumption that the evolution of the universe, a perfectly isolated system since by definition there is nothing outside it, is quantum and unitary in character.
+An immediate consequence is having to also adopt a quantum cosmological interpretation of the Anthropic Principle $\left[15-19\right]$, the “Fine-tuned Universe” version of it.
 This version of the Principle asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental physical constants since the beginning of time and the “value” of life today, in particular “sentient life” or “no life”.
@@
 Today, there would be a quantum superposition of tensor products, each a possible value of the fundamental constants multiplying the corresponding value of life.
-The measurement of either observable today would select the eigenvalues of both observables.
-All measurement outcomes are as if the value of the fundamental constants was selected at the beginning of time.
+The measurement of either observable today would select the eigenvalues of both maximally entangled observables.
+Everything is as if the value of the fundamental constants was selected at the beginning of time.
 Summing up, we are dealing with a unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue selected by the initial measurement (here the value of the fundamental constants).
-We have seen that this unitary evolution can always occur with quantum speedup.
-Being a quantum search (for one-to-one correlation), it should occur with quantum speedup even if the search (by the universe’s evolution under the fundamental physical laws and constants) is random $\left[19\right]$.
+We have seen that this unitary evolution can always occur with quantum speedup – there is the room for it.
+Being a quantum search (for one-to-one correlation), it should occur with quantum speedup even if the search (by the universe’s evolution under the fundamental physical laws and constants) is random $\left[12\right]$.
 It should consequently be teleological in character.
@@
 For the time being, the fact that quantum algorithms are certainly teleological evolutions, and that, under a quantum cosmological interpretation of the Anthropic Principle, natural evolutions would be teleological in character too, may shake the current belief that such a physical basis cannot exist.
-That quantum algorithms are teleological evolutions might even be acceptable from the psychological standpoint: it would be the strangeness of the microscopic quantum world and its short lived evolutions soon killed by decoherence.
+That quantum algorithms are teleological evolutions might even be acceptable from the psychological standpoint: it would be the usual strangeness of the microscopic quantum world and its short lived evolutions soon killed by decoherence.
 Things become quite different in the case of the evolutions of the living.
 The fact that what happens around and within us would be driven not only by the past but also by a future even billions of years ahead is more than puzzling.
-It might go along with ancient beliefs, but of a kind deliberately ignored by science.
+It might also go along with ancient beliefs, but of a kind deliberately ignored by science.
 Once science was accused of heresy for denying a physical basis to beliefs of the kind in question.
@@
 Of course, we are not alone in challenging the predominant idea that causality only goes forward in time.
-Besides the retrocausal approach to the quantum computational speedup $\left[4-8\right]$, we should mention time-symmetric quantum mechanics $\left[20-24\right]$, in particular its study of the possibility that a future choice changes a past quantum measurement outcome $\left[23\right]$, and ongoing attempts to study quantum processes without the need to preassume a definite causal structure $\left[25-28\right]$.
+Besides the retrocausal approach to the quantum computational speedup $\left[1-5\right]$, we should mention time-symmetric quantum mechanics $\left[20-24\right]$, in particular its study of the possibility that a future choice changes a past quantum measurement outcome $\left[23\right]$, and ongoing attempts to study quantum processes without the need to preassume a definite causal structure $\left[25-28\right]$.
 The present approach is naturally inspired by time-symmetric quantum mechanics and has in common with the attempts in question the fact of being open to the violation of the ordinary notion of causality, not in common the way of looking at the quantum description of quantum algorithms.
@@
 By the way, despite the importance of the author’s name, Costa de Beauregard’s explanation of quantum nonlocality never took off.
-Other precedents are Wheeler’s Participatory Anthropic Principle $\left[31,32\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[12\right]$.
+Other precedents are Wheeler’s Participatory Anthropic Principle $\left[31,32\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[13\right]$.
 These two principles share with the present work the notion of mutual causality between the observation today and the value of the fundamental physical laws and constants since the beginning of time mediated by a quantum evolution of the universe in between.
@@
 4 The time-symmetric quantum causal loops – formalization
-We formalize the notion of the time-symmetric quantum causal loops on the four drawer instance of Grover $\left[9\right]$quantum search algorithm.
-As anticipated in the Rationale, this immediately generalizes first to any number of drawers (under Long’s $\left[10\right]$ version of Grover algorithm that always provides the solution with certainty), then to any unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement.
+We formalize the notion of the time-symmetric quantum causal loops on the four drawer instance of Grover $\left[6\right]$ quantum search algorithm.
+As anticipated in the Rationale, this immediately generalizes first to any number of drawers, then to any unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement.
 The formalization consists of the following steps:
@@
 Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors.
-By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow333 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball.
+By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow444 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball.
 We omit this operantion that changes nothing for us.
@@
 In fact, it would tell her, immediately after the initial measurement of $\hat{B}$, the number of the drawer with the ball selected by Bob, which is also the solution of the problem – see the state of register $B$ in the bottom-left corner of Table I diagram.
-Of course, in the quantum description of the quantum algorithm with respect to Alice, she must be concealed from the outcome of the initial measurement444 More in general, one should assume that the problem-solver has a complete knowledge of the problem, thus also of the relation between problem-setting and solution.
+Of course, in the quantum description of the quantum algorithm with respect to Alice, she must be concealed from the outcome of the initial measurement555 More in general, one should assume that the problem-solver has a complete knowledge of the problem, thus also of the relation between problem-setting and solution.
 Knowledge of the problem-setting selected by the initial measurement would tell her also the solution.
@@
 5 The physical basis of the notion of teleological evolution
-We have seen, rather simply, that the complete quantum description of quantum algorithms is a quantum superposition of time-symmetric causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps.
+We have seen, by the way rather simply, that the complete quantum description of quantum algorithms is a quantum superposition of time-symmetric causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps.
 In the optimal quantum case, the number of computation steps is that needed to solve the problem in the optimal logical way benefiting of the advanced knowledge in question.
@@
 Consequently, the evolution of a universe toward the value of life turns out to be teleological in character.
-In the universe we are in, what turns out to be teleological in character is the evolution of the universe toward sentient life.
-Eventually, the evolution of life on earth, in principle obtainable from that of the universe by tracing out all the rest, even if decohered, would inherit the quantum speedup and the consequent teleological character of the evolution of the universe toward life.
-In the end, the notion that the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithm are also the physical basis of the teleological character of natural evolutions would seem to be physically plausible, this, of course, within a quantum cosmological view.
+Of course, in the universe we are in, what turns out to be teleological in character is the evolution of the universe toward sentient life.
+Eventually, as we have seen, the evolution of life on earth, in principle obtainable from that of the universe by tracing out all the rest, even if decohered, should inherit the quantum speedup and the consequent teleological character of the evolution of the universe toward life.
+In the end, the notion that the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithm are also the physical basis of the teleological character of natural evolutions would seem to be physically plausible, this, of course, within the quantum cosmological view.
 This section is a discussion of the many impacts that the present teleological view of natural evolutions might have on different areas of science and knowledge; of course, not forgetting the natural philosophy that gave rise to the notion of teleological evolution.
@@
 By the way, since the consequent teleological character of the evolution would be hidden in transition probabilities, the process would be difficult to distinguish from a classical stochastic process.
-Let us call time-symmetric also the kind of stochastic evolution that we have just outlined.
+Since also in this case the time-symmetric causal loops we are dealing with should be involved, let us call time-symmetric also the just outlined kind of stochastic evolution.
 We mention what could be an emblematic example of time-symmetric stochastic evolution.
@@
 So, why anticipating something designed for flight millions of years in advance?
-The usual explanation is of course a causal one: at first, feathers appeared at random and survived natural selection as good thermal insulators, later they proved useful for speeding escape from predators, perhaps for gliding from trees to the ground, and only eventually they turned out to be what was needed to conquer the air.
+The usual explanation is of course a causal one: at first, feathers appeared through random mutations and survived natural selection as good thermal insulators, later they proved useful for speeding escape from predators, perhaps for gliding from trees to the ground, and only eventually they turned out to be what was needed to conquer the air.
 According to the present teleological view of the evolutions of the living, even during the flightless period of the feathered dinosaurs, those random DNA mutations that would have led in the future to the conquest of the air – obviously an extremely successful outcome from the point of view of survival – would have more likely occurred.
@@
 If it were completely independent, it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup.
-Therefore, an obvious, approximate, way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact it is decohered and model it in a quantum way.
+Therefore, an obvious (approximate) way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact it is decohered and model it in a quantum way.
 There are encouraging precedents to this.
@@
 5.5 Impact on philosophy
-Of course, the current theory of teleological evolutions rehabilitates an ancient notion of natural philosophy $\left[39\right]$.
+Of course, the current theory of teleological evolutions would rehabilitate an ancient notion of natural philosophy $\left[39\right]$.
 We should therefore be allowed to look for the possibility that there are other relations between this theory and philosophy.
@@
 In effect, the teleological theory in question seems to revamp a full body of ancient philosophical views sort of opposite (and complementary) to the body of philosophical views emerged with the scientific revolution.
-Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view555 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops.
+Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view666 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops.
 and the causal view.
@@
 Consider the opposition between materialism and idealism.
-The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[31\right]$, the anthropic cosmological principles of Barrow and Tipler $\left[12\right]$, and the present theory is a mutual causality view.
+The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[31\right]$, the anthropic cosmological principles of Barrow and Tipler $\left[13\right]$, and the present theory is a mutual causality view.
 We find a similar situation in the opposition between the mechanistic and teleological views of evolutions, which obviously respond to the causal and mutual causality views of evolutions, respectively.
 For what concerns reductionism vs monism, we note that deriving the teleological character of the decohered evolutions of the living systems on Earth from that of the evolution of a quantum universe is an essentially monistic view.
-It presupposes that what we see around and inside is essentially inseparable from the whole.
+It presupposes that what we see around and within us is essentially inseparable from the whole.
 Let us mention other theories in the history of philosophy reminiscent of the mutual causality between sentient life today and the physical laws with their physical constants since the beginning of time:
@@
 (i) Look for a synthesis between the present notion of the time-symmetric quantum causal loops and time-symmetric quantum mechanics $\left[20-24\right]$ to which it is obviously inspired.
-(ii) Look for toy models that capture the mechanism for which the evolution of life on Earth (a subsystem), although decohered, inherits the quantum speedup and consequently the teleological character of the evolution of the universe (the system) toward life.
-(iii) Explore the possibility that other quantum features suspected to be involved in the fundamental processes of life derive top-down from a quantum character of the universe (as, presently, the teleological character of the evolutions of the living).
+(ii) Look for toy models that capture the mechanism for which the evolution of life on Earth (the mostly autonomous evolution of a subsystem), although decohered, inherits the quantum speedup and consequently the teleological character of the evolution of the universe (the system) toward life.
+(iii) The mechanism described in (ii) might apply to other life processes for which the presence of a quantum behavior is suspected, for example, to the possible quantum behavior of the brain.
+Although mostly autonomous from the rest of the universe, this behavior could derive from that of a quantum universe by tracing out everything external to the brain.
+Note the radical difference between this view and the usual one where people looks for an autonomous quantum behavior in the macroscopic, warm and wet brain.
+By the way, we could introduce a distinction between two kinds of quantum behaviors: autonomous, like those of microscopic systems sufficiently isolated from all the rest, and derived, namely quantum behaviors of even macroscopic subsystems obtained from that of a quantum universe – the system – by tracing out all the rest.
 (iv) Look for the possible impact of the teleological view on quantum cosmology.
-According to current assumptions, even the inanimate part of it, anyway an evolution toward life, should be endowed with quantum speedup and thus teleological in character.
+According to current assumptions, even the inanimate part of the latter, always an evolution toward life, should be endowed with quantum speedup and thus have teleological character.
 (v) The Anthropic Principle is limited to stating the correlation between the “value” of life and the value of the fundamental physical constants since the beginning of time.
@@
 (II) Toy models of teleological Darwinian processes.
-These evolutions could be either approximated by pure quantum processes endowed with quantum speedup or as decohered stochastic processes with transition probabilities time-symmetrically affected by the past and the future of the process.
+These evolutions could be either approximated as pure quantum processes endowed with quantum speedup or described as decohered stochastic processes with transition probabilities time-symmetrically affected by the past and the future of the process.
 (III) An application of the time-symmetric quantum causal loops to quantum self-organization.
@@
 7 Conclusions
-The work is, of course, unconventional. It might be useful to summarize in a consequential way the line of reasoning followed.
+The work is, of course, highly unconventional. It might be useful to summarize in a consequential way the line of reasoning followed.
 We have (trivially) shown that the sheer existence of the optimal quantum speedup of Grover quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops.
@@
 We have thus a unitary evolution between two one-to-one correlated measurement outcomes that does not change the problem-setting selected by the initial measurement.
-This is a key feature of quantum algorithms, not visible in their customary quantum description666 If the solution is not an invertible function of the problem-setting, to have one-to-one correlation between the two measurement outcomes, it suffices that the problem-solver, at the end, measures both the solution and the problem-setting.
+This is a key feature of quantum algorithms, not visible in their customary quantum description777 If the solution is not an invertible function of the problem-setting, to have one-to-one correlation between the two measurement outcomes, it suffices that the problem-solver, at the end, measures both the solution and the problem-setting.
 .
@@
 Eventually, we should take the quantum superposition of all the causal loops.
+Of course, the time-symmetrization leaves the ordinary quantum description mathematically unaltered but changes its single causal process into a quantum superposition of causal loops.
 Until now, the time-symmetrization in question has been imposed by the necessity of reproducing, in the quantum description of Grover algorithm, the causal loops logically implied by the sheer existence of its (optimal) quantum speedup.
@@
 This creates the room for a quantum speedup.
-There should be a quantum speedup even in the case that the search for one-to-one correlation is random.
+There is a quantum speedup even in the case that the search for one-to-one correlation is random.
 The process is a quantum superposition of time-symmetric causal loops also in the case that the two observables are already maximally entangled at the beginning of it.
@@
 This immediately raises the problem of explaining a quantum speedup in evolutions of the living that can be the size of the Earth and have durations of billions of years.
-The fact it is obviously impossible to scale up the quantum features of the microscopic world to such levels suggests the possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe.
+The fact it is obviously impossible to scale up the quantum features of the microscopic world to such levels suggests the only possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe.
 Of course, at the same time we should adopt quantum cosmology, namely the assumption that the universe evolves in a unitary quantum way.
@@
 It asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value of the fundamental physical constants since the beginning of time and the “value” of life today, in particular, either “existence of sentient life” or “absence of life”.
-We assume for simplicity that this correlation is one-to-one
+We assume for simplicity that this correlation is one-to-one.
 With the evolution of the universe quantum, also the correlation asserted by the Anthropic Principle becomes quantum.
@@
 In conclusion, the notion of the time-symmetric quantum causal loops turns out to be the common physical basis of such diverse physical phenomena as the quantum computational speedup, the consequent teleological character of quantum algorithms, quantum nonlocality, and – under a quantum cosmological interpretation of the Anthropic Principle – the teleological character of the evolutions of the living.
-Of course, it is a puzzling notion that radically changes a long-standing way of viewing the role of causality in the physical world.
-Regarding quantum phenomena in the microscopic world, we think to have provided clear evidence that the notion of the time-symmetric quantum causal loops is a necessary consequence of the notion of quantum computational speedup.
+Of course, that of teleological evolution is a puzzling notion that radically changes a long-standing way of viewing the role of causality in the physical world.
+Regarding the quantum phenomena in the microscopic world, we think to have provided clear evidence that the notion of the time-symmetric quantum causal loops is a necessary consequence of the notion of quantum computational speedup.
 Regarding the teleological character of the macroscopic evolutions of life, the situation becomes even more puzzling.
 We limit ourselves to noticing that this is a necessary consequence of the quantum cosmological assumption.
-Seeing the macroscopic evolutions of life as the partial trace over everything else of a quantum evolution of the universe toward life brings into the macroscopic world of everyday experience the strange character of the microscopic quantum world.
+Seeing the macroscopic evolutions of life as the partial trace over everything else of a quantum evolution of the universe toward life brings into the macroscopic world of everyday experience some of the strange characters of the microscopic quantum world.
 Acknowledgments

### v6

@@
 1 Introduction
-The notion of teleological evolution is that of an evolution toward a goal with an attractor in the same goal it will produce in the future.
-It is, of course, a puzzling notion since it necessarily implies a causal loop.
-As is well known, it was often applied to the evolutions of the living before the scientific revolution.
-With the advent of the latter, it was dismissed because of its alleged lack of physical basis.
-That was, of course, under the notion of causal evolution that is dominant in science even today.
-We take the opposite position in the present work. We will show that, in quantum mechanics, there is a well defined physical basis to the notion of teleological evolution.
-In the first place, we show that quantum algorithms are teleological evolutions, indeed evolutions toward a goal (the solution) with an attractor in the goal they will produce in the future (the solution again).
-Of course, seeing this requires a quantum description of quantum algorithms different from the ordinary one, which is naturally that of a causal process.
-The latter quantum description, with causality always going in the same time-direction, cannot describe causal loops.
-We show that the ordinary quantum description of quantum algorithms, although complete from the mathematical standpoint, is physically incomplete and is completed by time-symmetrizing it.
-This leaves the ordinary quantum description mathematically unaltered but replaces its single causal process by a quantum superposition of causal loops.
-In each of them, it is as if the problem-solver knew in advance, before beginning her problem-solving action, one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer logical operations (computation steps).
-Of course, we are dealing with an evolution that responds to the notion of teleological evolution: it is indeed toward a goal (the solution) with an attractor in the goal it will produce in the future (the solution again).
-This completion of the quantum description provides a unified quantitative explanation of all the optimal quantum speedups achieved in solving oracle problems.
-Given an oracle problem, the number of computation steps needed to solve it in an optimal quantum way is that needed to solve it in a optimal logical way benefiting of the advanced knowledge of any of the possible halves of the information about the solution.
-Of course, there is no such an explanation in current quantum computing science.
-Obviously, we have a causal quantum description for each quantum algorithm, but there is no known way of unifying them.
-This part of the work is a further development of the retrocausal approach to the quantum computational speedup $\left[1-5\right]$.
-Of course, both the approach in question and the part of the work discussed so far are unconventional and potentially controversial.
-The notion of causal loop is puzzling and mostly extraneous to today’s physics.
-However, we believe that this time we have given a very simple and easily checkable demonstration that the quantum computational speedup is physically due to quantum causal loops.
-Let us outline it.
-We consider the following problem.
+In this work we propose a physical basis for the notion of teleological evolution, that is, an evolution toward a goal with an attractor in the same goal it will produce in the future.
+It is, of course, a puzzling notion since it implies a causal loop.
+For centuries it has been applied by natural philosophy to the evolutions of the living; it was then dismissed with the advent of modern science because of its alleged lack of a physical basis.
+What we think should change that (also centuries old) position of science is the relatively recent discovery of the quantum computational speedup.
+In this work, by quantum (computational) speedup we mean its strict form, that is, the fact that quantum algorithms solve problems with fewer logical operations (computation steps) than classically possible.
+As we try to be interdisciplinary, let us provide an example of quantum speedup right away.
+Consider the following problem.
 Bob, the problem-setter, hides a ball in a chest of four drawers – we call the number of the drawer with the ball the problem-setting.
@@
 If it is not even in that drawer, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well.
-In the quantum case – by the optimal quantum search algorithm of Grover $\left[6\right]$ – she always solves the problem by opening just one drawer222 In the four drawers case, Grover algorithm yields the solution with certainty.
-In the following, by Grover algorithm we will always mean Long’s $\left[7\right]$ version of it, which does that always.
+In the quantum case – by the optimal quantum search algorithm of Grover $\left[1\right]$ – she always solves the problem by opening just one drawer.
+This would obviously be impossible in the classical case, it seems also impossible to our intuitive way of seeing the thing.
+Let us mention priorities along the path that led to this fundamental discovery.
+The notion that computation is possible in the quantum world, with the related notion of quantum bit, was introduced by Finkelstein $[2]$ in 1969.
+In 1982, Feynman $\left[3\right]$ highlighted the higher computational efficiency of the quantum world, showing that the classical simulation of a quantum process can require an amount of physical resources exponentially higher than the amount involved in the quantum process itself.
+In 1985, Deutsch $\left[4\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup.
+The reason we believe that that position of science should be changed is simply the fact that quantum algorithms turn out to be teleological evolutions, that is, evolutions toward a goal (the solution of the problem) with an attractor in the same goal they will produce in the future (the solution again).
+We will show that this is the immediate consequence of a retrocausal approach to the quantum speedup of ours and co-authors $\left[5-9\right]$.
+The present work is divided into two parts.
+The first is a review and further clarification of that retrocausal approach to the quantum speedup: it will justify the teleological character of quantum algorithms.
+The second is whether the same justification could constitute the physical basis of the teleological character of the evolutions of the living for which the teleological notion was conceived.
+The present introduction is divided into two corresponding parts.
+1.1 The teleological character of quantum algorithms
+Let us focus on Grover quantum search algorithm, by the way a most universal quantum algorithm that can solve with a quadratic quantum speedup any NP problem.
+We show that just opening to retroausality allows to produce a simple quantitative explanation of the optimal quantum speedup of Grover algorithm that will also apply to a very large class of quantum speedups.
+Let us begin with the four drawers case. Our argument goes as follows.
+The fact that Alice always solves the problem of locating the ball by opening just one drawer is logically equivalent to her knowing in advance, before she begins her problem-solving action, that the ball is in a given pair of drawers and benefiting of that knowledge to solve the problem by opening just one drawer.
+In fact, if the ball is in the drawer opened, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well, in either case having opened just one drawer.
+Of course, what Alice should know in advance is the mutually exclusive $or$ of the three possible ways of pairing the number of the drawer with the ball with that of another drawer.
+Note that this also means knowing in advance half of the information that specifies the solution of the problem.
+Then the question becomes: from where does Alice, before beginning her problem-solving action, always get that information about the solution?
+Since, to her, the only possible source of it is the solution of the problem she will produce and measure in the future, the only possible answer is that it comes to her from her future measurement of the solution.
+Of course, this implies retrocausality, more specifically a causal loop.
+The above immediately generalizes to any number, $N$, of drawers.
+Grover algorithm yields a quadratic quantum speedup, that is, always solves the problem by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers against the $\operatorname*{O}\left(N\right)$ of the best classical case.
+Also now, this quantum speedup can be logically justified by the fact that Alice knows in advance, in a mutually exclusive way, one of the possible halves of the information about the number of the drawer with the ball.
+This means that she knows in advance that the ball is (in a mutually exclusive way) in one of the possible tuples of $\sqrt{N}$ drawers.
+She then solves the problem in the best possible logical way by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers, which is indeed what Grover algorithm does.
+We can conclude that the optimal quantum speedup of Grover algorithm logically implies the mutually exclusive $or$ of well defined causal loops, each for each possible way of taking half of the information about the solution.
+In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution by opening fewer drawers.
+We can formulate the following advanced knowledge rule that, as checked in $\left[8\right]$, also holds for all the major quantum algorithms for the solution of oracle problems:
+The number of computation steps (oracle queries) needed to solve the (oracle) problem in an optimal quantum way is the number needed to solve it in an optimal logical way benefiting of the advanced knowledge of one (any one) of the possible halves of the information about the solution.
+Later we will show why this rule applies to any problem for which a problem setting and a corresponding solution can be identified.
+By the way, it is a rule that does not exist in mainstream quantum computing science.
+There, the only way of knowing the number of computation steps needed to solve a problem with a quantum speedup is finding a quantum algorithm that does that.
+As is well known, this is a matter of high ingenuity.
+Of course, this goes along with the fact that mainstream quantum computing science, as mainstream quantum mechanics, relies on the notion of causal evolution that is incompatible with that of causal loop.
+Let us make a digression.
+As the work goes on, we will naturally go into more and more detail, with the consequent risk of masking the forest with the trees.
+So, we would like to emphasize that the forest is right now in front of our eyes.
+The notion of causal loops is obviously controversial; probably most people believe that causal loops cannot exist except perhaps in very exotic physical situations such as the closed timelike curves allowed by general relativity.
+Now we would have, moreover trivially, demonstrated that the quantum speedup of a quantum algorithm logically implies their existence.
+This is the forest!
+As we will venture into very exotic situations ourselves, but in ordinary non-relativistic quantum mechanics, we would like to suggest that we keep this basic fact in mind: it is the source of all exoticism.
+Back to our line of thinking, of course we should ask ourselves what is the impact of the fact that a quantum speed up logically implies the mutually exclusive $or$ of causal loops on our way of seeing the quantum algorithm, that is, on the quantum description that we make of it.
+To this end, we should make reference to a quantum description of the quantum algorithm less incomplete than the usual one, which is limited to the process of solving the problem (where the quantum speedup resides) and lacks the initial measurement.
+Of course, the canonical quantum description of a quantum process must include the initial and final measurements and the unitary evolution in between.
+Of course, this problem is easily adjusted.
+It suffices to extend the usual quantum description of the process of solving the problem to the upstream process of setting the problem.
+The extended quantum description is thus as follows.
+An initial measurement, in a quantum superposition (or mixture) of all the possible problem settings, selects one of them at random222 The problem-setter would be free to unitarily change the randomly selected problem-setting into a wanted one.
+We disregard this possibility that, in the present context, changes nothing.
 .
-This would obviously be impossible in the classical case, it seems also impossible to our intuitive way of viewing it.
-It should already be clear from the above that that of the quantum speedup is a fundamental discovery that changes a long standing way of viewing things333 The notion that computation is possible in the quantum world, with the related notion of quantum bit, was introduced by Finkelstein $[8]$ in 1969.
-In 1982, Feynman $\left[9\right]$ observed that the classical simulation of a quantum process can require an exponentially higher amount of physical resources.
-In 1985, Deutsch $\left[10\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup.
+This is followed by the unitary computation of the corresponding solution and by the final measurement that reads it without changing the quantum state333 With four drawers, Grover algorithm yields the solution of the problem with certainty; when the drawers are more than four, by Grover algorithm we will always mean Long’s version of it $\left[10\right]$, which always provides the solution with certainty.
 .
-We can go now to our demonstration, which is surprisingly simple. Of course, we must adopt a way of thinking open to retrocausality.
-The fact that Alice always solves the problem of locating the ball by opening just one drawer is logically equivalent to her knowing in advance, before beginning her problem-solving action, that the ball is in a given pair of drawers, in a mutually exclusive way one of the three pairs obtained by pairing the drawer with the ball with another drawer.
-In fact, if the ball is in the drawer opened, she has solved the problem.
-If not, it must be in the other drawer and she has solved the problem of locating the ball as well, having opened just one drawer.
-Then the question becomes: from where does Alice always get the information that the ball is in a given pair of drawers?
-Since, to her, the only possible source of it is the solution of the problem she will produce and measure in the future, the only possible answer is that it comes to her from her future measurement of the solution.
-This is obviously a causal loop.
-Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution by opening correspondingly fewer drawers.
-This, in the mutually exclusive $or$ of all the possible ways of taking half of the information about the solution.
-We would argue that, with the above elementary logic, we have, so to speak, broken the sound barrier: the quantum speedup does logically imply causal loops.
-Then, one gos from logic to physics as follows.
-Involving only the measurements of commuting observables (the problem-setting and the corresponding solution), quantum algorithms are submitted to the same (classical) logic.
-Therefore they must host the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of an optimal quantum speedup; of course, logical mutual exclusivity of possibilities, moving to the quantum world, becomes their quantum superposition.
-We have already seen that the quantum description of this quantum superposition is generated by time-symmetrizing the ordinary quantum description of the quantum algorithm.
-This time-symmetrization must therefore be performed because, as we will see, is the only way of generating the causal loops that must be hosted in the quantum algorithm.
-Let us add a few comments to the above.
-By ordinary quantum description of quantum algorithms, we do not mean their usual quantum description.
-The latter is limited to the unitary process that solves the problem, where the quantum speedup resides, and the final measurement of the solution.
-It is trivially incomplete since it lacks the initial measurement.
-Since we need to complete the quantum description, in the first place we should emend this trivial incompleteness.
-We do this by extending the usual quantum description to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition (or, indifferently, mixture) of all possible problem-settings.
-In the case of Grover algorithm, we obtain a unitary evolution between an initial and a final one-to-one correlated measurement outcomes encoding respectively the problem-setting (the number of the drawer with the ball selected by the initial measurement) and the corresponding solution of the problem.
-We have called it the ordinary quantum description of the quantum algorithm, ordinary since it is still a causal description that cannot describe causal loops.
-Note that any quantum algorithm that produces the solution with certainty can be put in this form.
-If the solution is not an invertible function of the problem-setting, it suffices that Alice, at the end, measures both the solution and the problem setting.
-This is a trivial but important step: of course, the usual quantum description of quantum algorithms would completely obscure the approach we are following.
-Let us eventually note that we are dealing with the same kind of incompleteness of the ordinary quantum description that was ascribed by Einstein, Podolsky and Rosen $\left[11\right]$ to the quantum description of quantum nonlocality.
-Also this quantum description can be seen as a unitary evolution (the process of spatial separations of the two subsystems) between two one-to-one correlated measurement outcomes (those of measuring in the same basis the two respective maximally entangled observables) – we must think that this process goes from one to the other measurement outcome via the time the two subsystems were not spatially separate.
-Time symmetrized, it generates a quantum superposition of causal loops that play the role of the hidden variables envisaged by those authors to physically (locally) explain quantum non-locality $\left[5\right]$.
-This ends the first part of the work, a review and further clarification of the retrocausal approach to the quantum speedup.
-Of course, under the above explanation of their quantum speedup, quantum algorithms turn out to be teleological evolutions.
-The second part of the work is about the possibility that the quantum causal loops that stand at the basis of the teleological character of quantum algorithms are the missing physical basis of the teleological character of the evolutions of the living.
-This immediately raises two problems.
-One is explaining the presence of quantum features (the quantum speedup) in evolutions that can be the size of the Earth and have durations of billions of years.
-The obvious impossibility of scaling up the quantum features of the microscopic world to such levels points to the only possible solution: to scale them down from the largest and longest existing evolution there is, that of the universe.
-Of course, at the same time, we should adopt quantum cosmology, that is, the assumption that the universe – a perfectly isolated system since by definition there is nothing outside of it – evolves in a unitary quantum way.
-The other is to find a formal analogy between quantum algorithms and, under quantum cosmology, natural evolutions.
-Obviously, it should justify the fact that also natural evolutions can occur with a quantum speedup and are consequently teleological in character.
-Interestingly, the Anthropic Principle comes to our rescue.
-The Fine-tuned Universe version of it asserts that the unitary evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental constants since the beginning of time and the “value” of life today, in particular (“sentient life” or “no life”).
-The reason for this correlation is that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
-Having adopted quantum cosmology, of course, also implies a quantum interpretation of the Anthropic Principle.
-Simply, the correlation between the value of the physical constants since the beginning of time and the value of life today becomes quantum and its building up is performed by the universe’s unitary quantum evolution.
-This is what we were looking for.
-In fact, it is exactly what quantum algorithms do.
-They unitarily build up a quantum correlation between the problem-setting since the beginning of them and the solution of the problem at their end.
-Under this formal analogy, the evolution of the universe toward the value of life could be endowed with quantum speedup and would even in the case that the quantum search for the correlation in question is random, exactly as in random quantum search algorithms $\left[12\right]$.
-Of course, being endowed with quantum speedup, under the quantum causal loops explanation of it, the evolution in question is teleological in character.
-In the universe we are in, what becomes teleological in character is the evolution of the universe toward sentient life.
-Eventually, as we shall see, the evolution of life on Earth, in principle obtainable from that of the universe toward life by tracing out everything else, even if decohered, should inherit the quantum speed and thus the teleological character of the latter evolution.
-The same would consequently hold for any evolution of life on Earth, e.
-g.
-that of species.
-The notion of a teleological evolution of life is of course puzzling.
-For example, the Darwinian theory of evolution would change as as follow: those random DNA mutations that, in the future, will survive natural selection would become more probable.
-With the past influenced by the future, this is a completely new way of viewing the physical world.
-No wonder the teleological notion has been rejected by science.
-In conclusion, the work is highly unconventional, but its findings were obtained in a surprisingly simple and well-verifiable way.
-If it were correct, as we think, it would be a paradigm shift not only for the natural sciences, but also for the way we see physical reality.
-2 Rationale
-As the present work is about a fundamentally different way of viewing the role of causality in quantum (correlation) processes, before going to the formalizations we provide a Rationale that describes it conceptually.
-This Rationale is divided into two parts corresponding to the two parts in which the work is divided.
-The first provides the concept of the time-symmetric quantum causal loops, showing how they stand at the physical basis of the quantum computational speedup and quantum nonlocality.
-In particular, the time-symmetric quantum causal loop description of quantum algorithms highlights their teleological character, their being evolutions toward a goal (the solution of the problem) with an attractor in the goal they will produce in the future (the solution of the problem again).
-The second addresses the question whether the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithms could also be the missing physical basis of a teleological character of the natural evolutions for which the teleological notion was conceived.
-2.1 The time-symmetric quantum causal loops – the concept
-We go now to the first part of the work.
-We begin by showing, by the way very simply, that the sheer existence of the optimal quantum speedup delivered by Grover $\left[6\right]$ quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops.
-Of course, we must ignore the ban on the use of the notion of retrocausality.
-The fact that Alice (the problem-solver) always solves the problem of locating the ball by opening just one drawer logically implies that it is as if she always knew in advance, before beginning her problem-solving action, that the ball is in one of the three pairs of drawers formed by pairing the drawer with the ball with another drawer.
-Indeed, in any of these three mutually exclusive cases, she opens either drawer; if the ball is in it, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well.
-The question is then: from where does Alice always get (in a mutually exclusive way) this advanced knowledge of one of the possible halves of the information that specifies the solution?
-Since, to her, the only possible source of it is the solution she will produce and measure in the future, the only possible answer is that it comes to her from her future measurement of the solution.
-This is, of course, a causal loop that allows her to produce the solution by opening just one drawer.
-There must be, in a mutually exclusive way, one causal loop for each of the three possible ways of knowing in advance one of the possible halves of the information that specifies the solution.
-We consider now $N$ drawers.
-Grover algorithm (also in Long’s version) yields a quadratic quantum speedup, that is, always solves the problem by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers against the $\operatorname*{O}\left(N\right)$ of the best classical case.
-Also now, this quantum speedup is logically justified by the fact that Alice knows in advance, in a mutually exclusive way, one of the possible halves of the information about the number of the drawer with the ball.
-This means that she knows in advance that the ball is in one of the possible tuples of $\sqrt{N}$ drawers.
-She then solves the problem, in the best possible logical way, by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers, which is indeed what Grover algorithm does.
-We can therefore formulate the following advanced knowledge rule:
-The number of computation steps (drawer openings) needed to solve Grover’s problem in an optimal quantum way is that needed to solve it in an optimal logical way given the advanced knowledge of one of the possible halves of the information about the solution.
-Of course, involving only the measurements of commuting observables (the problem-setting and the corresponding solution), Grover algorithm, as any other quantum algorithm, is submitted to the same (classical) logic.
-Its quantum description should therefore describe the quantum superposition of the causal loops in question (logical mutual exclusivity of possibilities becomes their quantum superposition in the quantum world).
-However, the above immediately raises a problem.
-The current quantum description of quantum algorithms, as that of any quantum process, is causal in character.
-In other words, causality always goes in the same time direction.
-Obviously, it is a description that cannot describe causal loops; we tentatively ascribe this to its incompleteness.
-Before going to this assumed incompleteness, of a fundamental kind, we should emend a trivial incompleteness of the current quantum description of quantum algorithms.
-It is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution.
-It is obviously incomplete because it lacks the initial measurement.
-Then, we extend it to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition (or, indifferently, mixture) of all possible problem-settings.
-In the case of Grover algorithm, where the solution is an invertible function of the problem-setting (both are the number of the drawer with the ball), we obtain a unitary evolution (encoding the process of solving the problem) between an initial and a final one-to-one correlated measurement outcomes (encoding respectively the problem-setting selected by the initial measurement and the corresponding solution read by the final measurement).
-Of course, this unitary evolution never changes the problem-setting selected by the initial measurement.
-By the way, seeing quantum algorithms as unitary evolutions between an initial and a final one-to-one correlated measurement outcomes is a key step of the present work.
-In view of going beyond Grover algorithm, we note that, if the solution is not an invertible function of the problem-setting, we can have one-to-one correlation between the initial and final measurement outcomes as well by requiring that, at the end, the problem-solver measures both the solution and the problem setting.
-The only constraint is that the quantum algorithm determines the solution with certainty.
-We call the above, less incomplete, quantum description of the quantum algorithm its ordinary quantum description, ordinary because it is still a causal description that cannot describe causal loops.
-We should complete this ordinary quantum description in a way that leaves it mathematically unaltered but changes its causality structure – describes the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup.
-This is simply done time-symmetrizing it as follows.
-Instead of ascribing the selection of the pair of one-to-one correlated measurement outcomes all to the initial measurement, as customary, we require that it time-symmetrically shares between the initial and final measurements, of course in a quantum superposition of all the possible ways.
-In each time-symmetrization instance, the initial measurement selects one of the possible halves of the information that specifies the sorted out pair of correlated measurement outcomes (it selects a quantum superposition of problem-settings instead of the customary single one).
-Its measurement outcome unitarily propagates forward in time until becoming the state immediately before the final measurement, a quantum superposition of tensor products each a problem-setting multiplying the corresponding solution.
-The final measurement completes the selection, of course by selecting a single pair problem-setting and corresponding solution.
-Its measurement outcome, obviously the same of the ordinary quantum description, time-symmetrically propagates backward in time, by the inverse of the time-forward unitary transformation, until replacing the previous initial measurement outcome.
-This backward in time propagation, which inherits both selections, is an instance of the time-symmetrization of the ordinary quantum description of the quantum algorithm.
-Note that, mathematically, i.e.
-disregarding the time-direction of causality, it is the ordinary quantum description back again.
-Also note that, together, the forward and backward in time propagations describe a causal loop that changes its own past (we call it time-symmetric since it is generated by a time-symmetrization procedure).
-As seen by the problem-solver, from whom the outcome of the initial measurement should always be concealed (would give her information about the solution – see Subsection 4.2.), it is one of the causal loops logically implied by the sheer existence of the optimal quantum speedup.
-Eventually, we should take the quantum superposition of all the causal loops.
-Of course, the backward in time propagations of all time-symmetrization instances are identical with each other.
-Therefore, as required, this time-symmetrization leaves the unitary evolution between the two measurement outcomes mathematically unaltered.
-However, it changes the structure of causality along it.
-As seen by the problem-solver, the single causal process of the ordinary quantum description is replaced by the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
-As already said, in each causal loop it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future; then, she can use this knowledge to produce the solution with fewer computation steps.
-If she exploits it entirely, the quantum speedup is optimal.
-In equivalent terms, in each causal loop, the computational complexity of the problem to be solved reduces – this will become tangible in the formalizations of Section 4.
-The number of computation steps needed to solve the reduced complexity problem in an optimal quantum way is that needed to solve it in an optimal logical way.
-In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place correspondingly reduces.
-In view of what will follow, it is important to note that these results have been obtained independently of the distinctive character of the unitary evolution between the two correlated measurement outcomes (i.e.
-the actual quantum algorithm).
-It sufficed to know that it is a unitary evolution that one-to-one correlates two initially unrelated observables (the problem-setting observable and the solution observable) never changing the eigenvalue of the observable selected by the initial measurement.
-In the present work, we will always know that there is such a unitary evolution.
-It is Grover algorithm in the present case, but it could be any other quantum algorithm that provides the solution with certainty – see also $\left[4\right]$.
-Under the assumption that between two one-to-one correlated measurement outcomes there can always be a unitary evolution of the kind in question, the advanced knowledge rule holds for any oracle problem.
-Eventually, we note that the time-symmetrization in question, and the consequent time-symmetric quantum causal loops, are imposed by logic and quantum mechanics.
-Summarizing, logic says that the sheer existence of an optimal quantum speedup logically implies the mutually exclusive $or$ of well defined causal loops.
-Of course, these causal loop must appear in the quantum description of the quantum algorithm that delivers the quantum speedup that, involving only the measurements of commuting observables, is submitted to the same (classical) logic.
-In turn, it is easily seen that the only way of obtaining these causal loops is by time-symmetrizing the ordinary quantum description of the quantum algorithm.
-In fact, to have a causal loop, we must properly share the selection of the pair problem-setting and corresponding solution between the initial and final measurements.
-Because of the time-symmetry of the unitary evolution between the two one-to-one correlated measurement outcomes, we should share it evenly.
-This exactly specifies the time-symmetrization we are dealing with.
-Therefore, as long as logic and quantum mechanics are consistent, the time-symmetric quantum causal loops generated by the time-symmetrization operation should be exempt from the paradoxes that plague the usual kind.
-Interestingly, the time-symmetrization we are dealing with turns out to be mandatory also to satisfy a symmetry implicit in any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement.
-In fact, in this kind of process, we can postpone the initial measurement at will – all results must remain unaltered.
-When the two measurements become simultaneous, the time-symmetrization in question must already be there to become the atemporal symmetrization mandatory to represent the perfect symmetry between the simultaneous measurements of two maximally entangled observables.
-This further motivation for time-symmetrizing the unitary evolutions we are dealing with, besides further justifying the previous motivation, makes things much more general.
-Any such unitary evolution must be time-symmetrized, what leaves it mathematically unaltered but changes its ordinary, causal, quantum description into that a quantum superposition of time-symmetric causal loops.
-If the unitary evolution is a quantum algorithm that determines the solution with certainty, in each causal loop it is as if the problem solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future.
-However, this just creates the room for its occurring with quantum speedup; the quantum algorithm does not need any more to be optimal like in the case of Grover algorithm.
-How much of this room is exploited depends on the problem-solver, i.e.
-on the quantum algorithm.
-If the unitary evolution in question is not a quantum algorithm, but anyhow creates one-to-one quantum correlation between two initially unrelated observables (without changing the eigenvalue of the observable selected by the initial measurement), it can always occur with quantum speedup.
-In this more general case, the problem-solver advanced knowledge might loose any meaning but, of course, can be replaced by the reduction of the dimension of the Hilbert space in which the search for one-to-one correlation takes place.
-This reduction, in turn, creates the room for a quantum speedup.
-There is also the case that the unitary evolution does not build up any quantum correlation.
-It is the case of quantum nonlocality, namely the measurement in the same basis of two maximally entangled observables whose respective subsystems are in a state of spatial separation.
-Here maximal entanglement between the two observables is already there before the space-separation process.
-After the time-symmetrization, this process becomes a quantum superposition of time-symmetric quantum causal loops locally going from one to the other measurement outcome via the time the two subsystems were not yet spatially separate.
-Here, always locally, these causal loops tell either observable the state to assume in view of its future measurement – see $\left[5\right]$.
-We can see that these causal loops play the role of the hidden variables envisaged by Einstein and the others $\left[11\right]$ to locally (i.e.
-also physically) describe quantum nonlocality.
-Note that, as a consequence, the incompleteness we are currently ascribing to the quantum description of quantum algorithms is the same that Einstein and the others ascribed to the quantum description of quantum nonlocality.
-In both cases, what is incomplete is the quantum description of a quantum correlation.
-With this, we end the first part of the work.
-2.2 Teleological character of natural evolutions – a possible physical basis
-The second part of the work examines the possibility that the time-symmetric quantum causal loops underlying the teleological character of quantum algorithms are also the missing physical basis of the teleological character of the natural evolutions for which the teleological notion was conceived.
-This, of course, immediately raises the problem of explaining the existence of quantum features – indeed the quantum speedup – in evolutions that can be the size of the Earth and have a duration of billions of years.
-The obvious impossibility of scaling up the quantum features of the microscopic world to such levels suggests the possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe.
-Naturally, at the same time we should adopt quantum cosmology $\left[13,14\right]$, namely the assumption that the evolution of the universe, a perfectly isolated system since by definition there is nothing outside it, is quantum and unitary in character.
-An immediate consequence is having to also adopt a quantum cosmological interpretation of the Anthropic Principle $\left[15-19\right]$, the “Fine-tuned Universe” version of it.
-This version of the Principle asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental physical constants since the beginning of time and the “value” of life today, in particular “sentient life” or “no life”.
-Let us assume, for simplicity, that this correlation today is one-to-one.
-The significance of this principle lies in the fact that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
-Under a quantum cosmological interpretation of it, the correlation addressed by the Anthropic Principle becomes quantum.
-We should think of a value of the fundamental constants observable and a value of life observable.
-At the beginning, there should be a quantum superposition (or mixture) of all the possible universes, each with its value of the fundamental constants.
-An initial measurement of the value of the fundamental constants observable would select one value at random.
-The subsequent unitary evolution of the universe would produce, at the present time, the corresponding value of life, which would then be measured.
-In Subsection 5.2., after the formalization of the notion of the time-symmetric quantum causal loops, we will provide a formalized toy model of the building up of the correlation in question on the part of the universe’s evolution.
-By the way, since the value of the fundamental constants never changes along the universe’s evolution, in the universe we are in we can do without the initial measurement (when there were no observers).
-Today, there would be a quantum superposition of tensor products, each a possible value of the fundamental constants multiplying the corresponding value of life.
-The measurement of either observable today would select the eigenvalues of both maximally entangled observables.
-Everything is as if the value of the fundamental constants was selected at the beginning of time.
-Summing up, we are dealing with a unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue selected by the initial measurement (here the value of the fundamental constants).
-We have seen that this unitary evolution can always occur with quantum speedup – there is the room for it.
-Being a quantum search (for one-to-one correlation), it should occur with quantum speedup even if the search (by the universe’s evolution under the fundamental physical laws and constants) is random $\left[12\right]$.
-It should consequently be teleological in character.
-In the universe we are in, the search for one-to-one correlation between the value of the fundamental constants and the value of life becomes search for sentient life.
-It is thus the universe’s evolution toward sentient life that occurs with quantum speedup and is consequently teleological in character.
-In principle, the evolution of life on Earth can be obtained from that of the universe toward sentient life by tracing out all the rest.
-Evolution times would be identical and the amount of the resources employed by the evolution of life on Earth could not exceed those employed by its image in the evolution of the universe that comprises them.
-In the end, the evolution of life on Earth, even if decohered, would occur with quantum speedup and would thus be teleological in character.
-This should also hold for the evolutions of species, etc..
-Of course, all the above is puzzling.
-A physical basis to the dismissed notion of teleological evolution would be a paradigm shift for the natural sciences.
-As is well known, empirically, the teleological notion better fits with the functional structure and the evolutions of the living.
-If it has been discarded, it is only because of the assumed absence of a physical basis.
-For the time being, the fact that quantum algorithms are certainly teleological evolutions, and that, under a quantum cosmological interpretation of the Anthropic Principle, natural evolutions would be teleological in character too, may shake the current belief that such a physical basis cannot exist.
-That quantum algorithms are teleological evolutions might even be acceptable from the psychological standpoint: it would be the usual strangeness of the microscopic quantum world and its short lived evolutions soon killed by decoherence.
-Things become quite different in the case of the evolutions of the living.
-The fact that what happens around and within us would be driven not only by the past but also by a future even billions of years ahead is more than puzzling.
-It might also go along with ancient beliefs, but of a kind deliberately ignored by science.
-Once science was accused of heresy for denying a physical basis to beliefs of the kind in question.
-We could be facing the reverse problem now: proposing a physical basis to the notion of teleological evolution could well be seen as a scientific heresy.
+Note that there is one-to-one correlation between the initial measurement outcome, which encodes the problem-setting, and the final one, which encodes the corresponding solution – in Grover algorithm, they are both the number of the drawer with the ball.
+Furthermore, there is a unitary evolution connecting the two measurement outcomes, despite the fact it comprises the final measurement (which indeed does not change the quantum state).
+This extension of the usual quantum description of quantum algorithms to the process of setting the problem will play a key role in the present work.
+We will call it the ordinary quantum description of the quantum algorithm, ordinary because it is still the ordinary causal description of mainstream quantum mechanics.
+We can go now to the impact of the logical implication in question on our way of describing the quantum algorithm.
+The first thing to observe is that, involving only the measurements of commuting observables (the problem-setting and the solution observables), the quantum algorithm is subject to the same (classical) logic of the logical implication in question; in other words, the non-classical quantum logic of non-commuting observables is not involved here.
+The quantum algorithm that delivers the optimal quantum speedup must therefore accommodate the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of its quantum speedup (in the quantum world, the logical mutual exclusive $or$ of possibilities becomes their quantum superposition).
+However, being causal in character, its ordinary quantum description cannot describe causal loops.
+According to the present line of thinking, we must conclude that the ordinary quantum description of the quantum algorithm is either wrong or incomplete, since it does not and, being causal, cannot describe causal loops.
+We will see that it is incomplete as it is completed by time-symmetrizing it.
+We should require that the initial and final measurements time-symmetrically contribute to the selection of their pair of one-to-one correlated measurement outcomes – the problem-setting and the corresponding solution.
+The selection performed by the initial measurement should propagate forward in time until becoming the state immediately before the final measurement, that performed by the final measurement, time-symmetrically, should propagate backward in time until replacing the previous initial measurement outcome.
+This is a causal loop that changes its own past; it is, indeed, one of the causal loops logically implied by the sheer existence of the quantum speedup.
+Of course, we should then take the quantum superposition of all the possible ways of evenly sharing.
+We will see that all this, mathematically, leaves the unitary part of the ordinary quantum description of the quantum algorithm unaltered but changes the description of the behavior of causality along that part.
+The single, unitary, causal process of the ordinary quantum description of the quantum algorithm changes into the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup.
+We will (easily) show that this time-symmetrization is the only possible way to generate the quantum superposition of the causal loops in question, which in turn must be hosted within the quantum algorithm.
+Therefore, according to the current line of thinking, it is imposed by logic and quantum mechanics.
+Most importantly to present ends, the fact that, in each causal loop, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and used this knowledge to produce the solution with fewer computation steps highlights the teleological character of quantum algorithms.
+They are indeed evolutions toward a goal (the solution) with an attractor in the very goal they will produce in the future (the solution again).
+Until now we have focused on Grover quantum search algorithm, but the results obtained go well beyond it.
+In the first place, we never needed to know the unitary evolution between the two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement (i.e., Grover algorithm).
+It sufficed to know that such an evolution exists.
+In the second place, until now, the time-symmetrization of the ordinary quantum description of the quantum algorithm has been imposed by the necessity of generating the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of its quantum speedup.
+However, it turns out that there is another reason for time-symmetrizing, independent of this.
+Let us consider the unitary (thus reversible) part of the quantum algorithm.
+It is a unitary evolution interconnecting two one-to-one correlated measurement outcomes that never changes the eigenvalue selected by the initial measurement.
+This allows to postpone the initial measurement at will.
+When made simultaneous with the final measurement, the time-symmetrization in question must already be in place to become the perfect atemporal symmetry between the simultaneous measurements of two maximally entangled observables.
+The two things together make the results obtained in the case of Grover algorithm applicable to a very general class of quantum processes involving quantum correlation. In fact, they imply that:
+Any unitary evolution between an initial and final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is the quantum superposition of the causal loops generated by time-symmetrizing its ordinary quantum description.
+We call the above the causal loop rule of quantum correlation.
+We can see that it is a generalization also when the process is still a quantum algorithm.
+In fact, until now we have remained committed to Grover algorithm that is an optimal one.
+However, also the time-symmetrization of the ordinary quantum description of non-optimal quantum algorithms yields a quantum superposition of causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution.
+The point is that this time she may not exploit this information at all, or not fully exploit it.
+For example, the latter is the case of the random quantum search algorithms $\left[11\right]$, which provides a suboptimal quantum speedup.
+We will be interested in this kind of quantum algorithms further on.
+In view of applying the causal loop rule of quantum correlation to processes that have the form of quantum algorithms but are not that, let us note that we can do without the anthropomorphic notion of a problem-solver that produces the solution with fewer computation steps because she knows in advance half of the information about it.
+We can also avoid to speak of a “quantum algorithm”.
+We can say that, in each time-symmetrization instance, there is a reduction of the dimension of the Hilbert space in which the search for the one-to-one correlation between the initial and final measurement outcomes takes place and consequently the correlation in question can be achieved with a quantum speedup.
+Of course, the process does not even need to have the form of a quantum algorithm.
+In other words, it does not need to create one-to-one correlation between two initially unrelated observables.
+The two observables can be maximally entangled before the process begins.
+As we shall see, quantum nonlocality is exactly a process of this kind and the quantum superposition of causal loops generated by its time-symmetrization explains the EPR (Einstein, Podolsky, and Rosen) paradox.
+1.2 The possible teleological character of the evolutions of the living
+The teleological character of quantum algorithms naturally raises the question of whether the quantum superposition of causal loops that underlie it can also be the missing physical basis of the teleological character of the evolutions of the living for which the teleological notion was conceived.
+Having established the teleological character of quantum algorithms, probing this possibility would even seem to be a duty of the physical science to the natural sciences.
+One difficulty, of course, would be to explain the presence of quantum features – the quantum speedup – in the evolutions of life that can be Earth-sized and last billions of years.
+However, the obvious impossibility that the quantum features of the microscopic world scale up to such levels opens up another possibility: that they scale down from the largest and longest existing evolution, that of the universe.
+Of course, at the same time, we should adopt quantum cosmology $\left[12,13\right]$, i.e., the assumption that the universe – a perfectly isolated system since by definition there is nothing outside it – evolves in a unitary quantum way.
+Interestingly, the corresponding quantum cosmological interpretation of the Anthropic Principle $[14-18]$ brings all things together.
+The Fine-tuned Universe version of this principle asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental constants since the beginning of time and the “value” of life today (in particular: “presence of sentient life” or “absence of life”).
+The reason for this correlation is that the slightest variation in the value that the fundamental constants have in the universe we are in would produce a radically different universe, most likely incapable of hosting life.
+Let us assume, for simplicity, that this correlation is one-to-one.
+The quantum interpretation of the Anthropic Principle produces a formal identity between the building up, by the unitary evolution of the universe under the fundamental physical laws and constants, of the quantum correlation between the value of the fundamental constants since the beginning of time and the value of life today on the one side and, on the other, the building up, by a quantum search algorithm, of the quantum correlation between the value of the problem-setting since its beginning and the value of the solution at its end.
+In other words, the unitary evolution of the universe toward the value of life, formally, turns out to be a quantum search algorithm.
+These quantum algorithms yield a quantum speedup even if the search for the correlation is random $\left[11\right]$.
+It is indeed reasonable (and conservative) to assume that the search for the correlation between the value of the fundamental constants since the beginning of time and the value of life today by the universe’s evolution – the search for life in the universe we are in – is random.
+According to this formal identity, the evolution of the universe toward the value of life should occur with quantum speedup and would consequently be teleological in character.
+In the universe we are in, the same should hold for the evolution of the universe toward life.
+It is then easy to show that this teleological character should be inherited by the evolution (and the evolutions) of life on Earth, derivable from that of the universe by tracing out everything else.
+In fact, evolution times should be the same for the evolution of life on Earth and its image in the evolution of the universe and the resources employed by the former evolution cannot exceed those involved by the latter that comprises them.
+Of course, that the universe evolves in a unitary quantum way, although consistent with the theory of decoherence, for which quantum processes decohere by losing information in their environment, is just an assumption.
+It should instead be firm the derivation from it of the teleological character of the evolutions of the living.
+In fact, the Fine-tuned Universe version of the Anthropic Principle is universally accepted and the quantum cosmological interpretation of it is straightforward.
+Summing up, we would conclude with the conditional statement that, if quantum cosmology were right, the evolutions of the living would be teleological in character.
+Let us also observe that this would be a paradoxical possibility.
+For example a way of making teleological a Darwinian evolution could be to assume that those random DNA mutations that, in the future, will turn out to survive because more suited to the environment, because of that, would become more probable.
+By the way, this would be exactly as paradoxical as the causal loop explanation of the quantum speedup.
+The only difference is that the quantum computational speedup belongs to a quantum world that is notoriously the place of queer phenomena.
+A teleological Darwinism would instead concern the macroscopic world we see around us.
+In effect, evolutions attracted by their future would be a radically different way of viewing physical reality.
+In the next Section 2, we try to position the present work within the literature that challenges the still dominant notion of causal evolution.
+In Section 3, we review and further develop the quantum causal loop description of quantum algorithms.
+Section 4 is the (of course, speculative) transposition of the teleological description of quantum algorithms to natural evolutions.
+Section 5 is on the prospects for further research.
+Section 6 gives the conclusions.
 2 Positioning the work
-We try and position the present work.
 Of course, we are not alone in challenging the predominant idea that causality only goes forward in time.
-Besides the retrocausal approach to the quantum computational speedup $\left[1-5\right]$, we should mention time-symmetric quantum mechanics $\left[20-24\right]$, in particular its study of the possibility that a future choice changes a past quantum measurement outcome $\left[23\right]$, and ongoing attempts to study quantum processes without the need to preassume a definite causal structure $\left[25-28\right]$.
-The present approach is naturally inspired by time-symmetric quantum mechanics and has in common with the attempts in question the fact of being open to the violation of the ordinary notion of causality, not in common the way of looking at the quantum description of quantum algorithms.
-For example, in $\left[27\right]$, quantum algorithms are considered examples of processes with a definite causal structure because, in their current quantum description, causality always goes forward in time.
-So far so good but, consequently, they are excluded (by the approach in question) from the phenomena that can accommodate causality violations.
+Other approaches to quantum mechanics that share this feature are time-symmetric quantum mechanics $\left[19-23\right]$, particularly the study of the possibility that a future choice changes a past quantum measurement outcome $\left[22\right]$, and ongoing attempts to study quantum processes without the need to assume a definite causal structure $\left[24-27\right]$.
+The present retrocausal approach to the quantum speedup is naturally inspired by time-symmetric quantum mechanics and has in common with the attempts in question the fact of looking beyond the ordinary notion of causality, not in common the way of looking at quantum algorithms.
+For example, in $\left[26\right]$, quantum algorithms are considered examples of processes with a definite causal structure because, in their current quantum description, causality always goes forward in time.
+So far so good but, consequently, quantum algorithms are excluded from the phenomena that can accommodate causality violations.
 Of course, for us, this means putting too much faith in their current quantum description.
-As we have seen, with their quantum computational speedups, quantum algorithms should be a most clear example of violation of one-way causality – what obviously implies changing their current quantum description.
+To us, with their quantum computational speedups, quantum algorithms are the seminal example of causality violation.
 Some concepts of the work have precedents.
-One is the local explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[29\right]$.
+One is the celebrated EPR paper $\left[28\right]$.
+In it, Einstein, Podolsky, and Rosen already argued that the quantum description is incomplete.
+More specifically, they claimed that the quantum description of quantum nonlocality is incomplete because it mathematically predicts something that it does not physically explain, namely the “spooky action at a distance”.
+As we will see, time-symmetrizing the ordinary quantum description of quantum nonlocality, as we did with that of quantum algorithms, completes it.
+It generates a quantum superposition of causal loops that play the role of the hidden variables those authors envisaged to complete the quantum description.
+Another precedent is the local, retrocausal, explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[29\right]$.
 Consider the measurements in the same basis of two maximally entangled observables in a state of spatial separation of the respective subsystems.
 Costa de Beauregard assumed that the outcome of the first of the two measurements locally propagates backward in time, by the inverse of the unitary transformation that spatially separated the two subsystems, until these were not yet spatially separate.
-Here, still locally, it would tell the other observable the state to assume in view of its future measurement.
-This retrocausal, local, explanation of quantum nonlocality is similar to ours.
-The difference is that sending backward in time an entire measurement outcome can originate paradoxes and, of course, is not imposed by quantum mechanics.
+There, always locally, it would change the state of the other observable in a way that ensures the appropriate correlation between the two future measurement outcomes.
+This retrocausal explanation of quantum nonlocality is similar to ours as it also involves a causal loop.
+The difference is in the character of the causal loops.
+Sending backward in time an entire measurement outcome can originate paradoxes and, of course, is not imposed by quantum mechanics but is instrumental to finding a local explanation of quantum nonlocality.
+As already said, the causal loops originated by time-symmetrizing the ordinary quantum description are instead imposed by quantum mechanics.
+Therefore, first, they must be there independently of the fact of explaining quantum non-locality locally; second, as long as quantum mechanics is consistent, they should be exempt from the paradoxes that plague the usual kind.
 Interestingly, at the time, Costa de Beauregard was an assistant of De Broglie, who forbade him to publish that work for years until the publication, by Wheeler and Feynman, of their emitter-absorber theory that also resorted to retrocausality – see $\left[30\right]$ for an account of this.
@@
 By the way, despite the importance of the author’s name, Costa de Beauregard’s explanation of quantum nonlocality never took off.
-Other precedents are Wheeler’s Participatory Anthropic Principle $\left[31,32\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[13\right]$.
-These two principles share with the present work the notion of mutual causality between the observation today and the value of the fundamental physical laws and constants since the beginning of time mediated by a quantum evolution of the universe in between.
-Dating from before Deutsch’s 1985 discovery, the two principles in question make no reference to the quantum speedup of the universe’s evolution and, a-fortiori, its consequent teleological character.
-4 The time-symmetric quantum causal loops – formalization
-We formalize the notion of the time-symmetric quantum causal loops on the four drawer instance of Grover $\left[6\right]$ quantum search algorithm.
-As anticipated in the Rationale, this immediately generalizes first to any number of drawers, then to any unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement.
+Still other precedents are Wheeler’s Participatory Anthropic Principle $\left[31,32\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[12\right]$.
+These two principles share with the present work the notion of mutual causality between the observation today and the “value” of the fundamental physical laws and constants since the beginning of time mediated by a quantum evolution of the universe in between.
+Dating to before Deutsch’s 1985 discovery, of course the two principles in question make no reference to a quantum speedup of the universe’s evolution and its consequent teleological character.
+3 Completing the quantum description of quantum algorithms
+This section is divided into two parts.
+The first begins with a constructive critique of the ordinary quantum description of Grover algorithm and ends with a generalization of the results obtained.
+We show (i) why the quantum description in question, being causal in character, must be considered incomplete, (ii) how it should be completed, and (iii) how both things extend to a very general class of quantum correlation processes.
+The second part formalizes the completion procedure consisting in a time-symmetrization of the ordinary quantum description.
+3.1 Updating the EPR critique of the quantum description
+Of course, the first assertion that the quantum description is incomplete goes back to the famous EPR paper by Einstein, Podolsky and Rosen $\left[28\right]$.
+As is well known, Einstein in particular, argued that the ordinary quantum description of quantum nonlocality is incomplete because it does not physically explain the “spooky action at a distance” that it mathematically predicts.
+They also stressed the need to devise ”hidden variables” that would complete the quantum description.
+The present critique of the quantum description could be considered an update of that of Einstein and the others because:
+(i) it shows that Einstein et al.’s assertion that the quantum description of quantum nonlocality is incomplete can be extended to quantum algorithms, of course discovered exactly half a century after the publication of the EPR paper;
+(ii) the reason for the incompleteness of the quantum description turns out to be the same in both cases;
+(iii) in the quantum computational context, it turns out to be evident that the quantum description should be completed by time-symmetrizing it: completing in the same way the quantum description of quantum nonlocality, locally (physically) describes the spooky action at a distance.
+First, we develop our critique of the quantum description on that of the four drawers instance of Grover quantum search algorithm $\left[1\right]$.
+For convenience, we summarize the part of this critique anticipated in the introduction.
+In its simple instance, the problem solved by Grover algorithm is as follows.
+Bob, the problem-setter, hides a ball in a chest of four drawers.
+Alice, the problem-solver, is to locate the ball by opening drawers.
+In the classical case, she has to open up to three drawers.
+In the quantum case – by the optimal quantum search algorithm of Grover – she always solves the problem by opening just one drawer.
+This would obviously be impossible in the classical case.
+It is natural to ask ourselves how it is possible that, in the quantum world, a ball hidden in one of four drawers can always be located by opening a single drawer.
+Of course, if one goes into the details of the quantum algorithm, one sees the ingenious interplays between quantum state superpositions and interferences that lead to producing the location of the ball by opening just one drawer.
+However, these interplays are naturally different from quantum algorithm to quantum algorithm and, as far as we know, in mainstream quantum information there is no known way of unifying a quantitative reason for which they yield a quantum speedup.
+A unified, quantitative, reason is provided in the retrocausal approach to the quantum speedup of ours and coauthors but, as things are now, this unconventional approach never took off within the scientific community.
+In the next subsection we address the unification problem.
+We will see that a unifying quantitative explanation of the quantum speedups becomes quite evident once the quantum descriptions of quantum algorithms is complete.
+It is a unifying explanation since it is the same for all quantum algorithm.
+Since deriving it requires an unconventional way of reasoning, open of course to retrocausality, we will give first the rationale then, in subsection 3.2., the formalizations.
+3.1.1 Grover quantum search algorithm as a quantum superposition of causal loops
+We have already shown in the Introduction that the sheer existence of the optimal quantum speedup of Grover algorithm logically implies the mutually exclusive $or$ of causal loops.
+In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer logical operations (computation steps).
+In the optimal quantum case, the number of computations steps is that needed to solve the problem in an optimal logical way benefiting of the advanced knowledge in question.
+Always in the introduction, we wandered what was the impact of the above fact on the current quantum description of the quantum algorithm, noticing that one should take a quantum description not limited to the process of solving the problem but extended to the upstream process of setting the problem.
+Our ordinary quantum description of the quantum algorithm therefore became:
+An initial measurement in a quantum superposition (or mixture) of all the possible problem-settings selects one of them at random.
+This is followed by the unitary computation of the corresponding solution, which of course does not change the problem-setting selected by the initial measurement, and the final measurement of the solution, which does not change the state immediately before it.
+We called this quantum description ordinary because it is still the customary, causal, quantum description that cannot describe causal loops.
+We observed that, involving only the measurements of commuting observables (the problem-setting and the solution), quantum algorithms are submitted to the same (classical) logic of the logical implication in question.
+Therefore, Grover algorithm should host the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of its optimal quantum speedup.
+Of course, this is not represented by its present ordinary quantum description, but we show that it is after completing this quantum description by time-symmetrizing it.
+This time-symmetrization is, of course, submitted to the requirement of generating the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
+As we shall see, the only possible way of obtaining this quantum superposition is by requiring that the initial and final measurements time-symmetrically contribute to the selection of the sorted out pair of one-to-one correlated measurement outcomes.
+Things should go as follows.
+In each time symmetrization instance, the initial measurement selects one of the halves of the information that specifies the sorted out problem-setting (and thus also the corresponding solution) – instead of selecting a single problem-setting as customary, it selects a suitable quantum superposition of problem-settings.
+Its measurement outcome unitarily propagates forward in time becoming, immediately before the final measurement, a quantum superposition of tensor products, each one of the problem-settings selected by the initial measurement multiplying the corresponding solution.
+The final measurement of the solution completes the selection.
+Note that, encoding the tensor product of the sorted out problem-setting and the corresponding solution, the final measurement outcome is the same of the ordinary quantum description.
+Eventually, time-symmetrically, this measurement outcome propagates backward in time, by the inverse of the time-forward unitary evolution, until replacing the previous initial measurement outcome.
+At this point, we should make two important observations:
+(i) The backward in time propagation, which inherits both selections, is an instance of the time-symmetrization of the quantum algorithm.
+Disregarding the time-direction of causality, read from left to right, it is the unitary part of the ordinary quantum description of the quantum algorithm back again.
+Since this is the same in each time-symmetrization instance, also the quantum superposition of all instances is that unitary part.
+Summing up, we can say that the time-symmetrization leaves the unitary part of the ordinary quantum description of the quantum algorithm mathematically unaltered.
+What was incomplete is the physical part of the quantum description, that is, the description of the behavior of causality along the unitary part of the quantum process.
+(ii) In each time-symmetrization instance, the forward and the backward in time propagations, together, constitute a causal loop that changes its own past (the outcome of the initial measurement).
+As we will see in the formalizations of subsection 3.2, this also changes the initial complete ignorance of the solution on the part of the problem-solver into knowledge of one of the possible halves of the information about it.
+There is a corresponding reduction of the computational complexity of the problem to be solved and the dimension of the Hilbert space in which the search for the solution takes place.
+The number of computation steps needed to solve the problem in an optimal quantum way is the number needed to solve in an optimal logical way the reduced computational complexity problem.
+We are of course dealing with one of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
+It is easy to show that the above time-symmetrization is the only possible way of generating the causal loops whose quantum superposition the quantum algorithm must host.
+In fact, to have a causal loop, in the first place the selection of the information that specifies the sorted out pair of correlated measurement outcomes must properly share between the initial and final measurements with the selection performed by the initial measurement propagating forward in time until becoming the state immediately before the final measurement and that successively performed by the final measurement propagating backward in time until replacing the previous initial measurement outcome.
+Of course, this creates the forward and backward in time parts of a causal loop.
+Moreover, to (exactly) generate the causal loops logically implied by the sheer existence of the optimal quantum speedup, the sharing should be even.
+Note that this is nothing but Laplace’s demon brought into the quantum world.
+Here, due to the possibility of having quantum state superpositions, there is more freedom.
+Indeed, it is a freedom exploited by the time-symmetrization process to create the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
+Eventually, let us observe that the quantum causal loops we are dealing with are generated by a mandatory time-symmetrization procedure – mandatory to generate the causal loops that must be hosted in the quantum algorithm.
+Therefore, as long as quantum mechanics is consistent, these causal loops should be exempt from the paradoxes that plague the causal loops of the usual kind.
+When needed to distinguish them from the latter kind, we will call them time-symmetric causal loops because generated by a time-symmetrization procedure.
+3.1.2 Generalization
+All the above has been developed for Grover algorithm. However, let us show how it generalizes. We should make two premises:
+(i) All the results achieved until now have been obtained with no need of knowing the unitary evolution between the two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement – i.e.
+the quantum algorithm itself.
+We only needed to know that such a unitary evolution exists.
+This was the case until now, the unitary evolution being that of Grover algorithm (it will always be the case in the following).
+(ii) Since the eigenvalue selected by the initial measurement remains unchanged throughout the unitary evolution in question, one is free to postpone that measurement at will.
+When performed simultaneously with the final measurement, the time-symmetrization we are dealing with must already have been put in place to represent the perfect (atemporal) symmetry between the simultaneous measurements of two maximally entangled, one-to-one correlated, observables.
+This is a second motivation for time-symmetrizing the ordinary quantum description of the quantum process in question, independent of the need of reproducing in the quantum algorithm the causal loops logically implied by the sheer existence of the quantum speedup.
+There is, of course, a redundancy between the two motivations.
+We produced both because together they say more.
+Putting (i) and (ii) together, brings us well beyond the optimal Grover algorithm considered until now. In fact we obtain that:
+The complete quantum description of any unitary evolution between an initial and a final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is the quantum superposition of the causal loops generated by time-symmetrizing it.
+We call the above the causal loop rule of quantum correlation. Let us examine a few of its consequences.
+In the first place, to be a quantum superposition of causal loop, the quantum algorithm does no more need to be optimal444 If the solution is not an invertible function of the problem setting, we can have one-to-one correlation between the initial and final measurement outcomes by requiring that the final measurement measures both the solution and the problem-setting.
+.
+The time-symmetrization of its ordinary quantum description always yields that it is as if the problem-solver knew in advance one of the possible halves of the information about the solution.
+In equivalent terms, there is a corresponding reduction of the computational complexity of the problem to be solved, and the dimension of the Hilbert space in which the search for the solution (or, identically, the one-to-one correlation) takes place.
+However, any of these facts just creates the room for a quantum speedup.
+To also have a quantum speedup, it takes that the problem-solver, or more generally the unitary evolution, exploits such a room.
+The room is fully exploited in the optimal quantum case, but it can also be not exploited at all or only partly exploited.
+The latter is the case of the random quantum search algorithm $\left[11\right]$ that yields a suboptimal quantum speedup. Later, we will need to make reference to this quantum algorithm.
+Of course, the quantum process may not even be a quantum algorithm. For example, quantum nonlocality is a process of this kind.
+It in fact consists of a unitary evolution (the process of spatial separation between the two subsystems) between two one-to-one correlated measurement outcomes – those of measuring in the same basis the two maximally entangled observables.
+The process of spatial separation, of course, never changes the eigenvalue selected by any measurement.
+Here, each causal loop locally goes from one to the other measurement outcome via the time the two subsystems were not yet spatially separate where, always locally, it changes the state of both observables in a way that ensures the appropriate correlation between the two future measurement outcomes – see $\left[9\right]$.
+This is exactly the role that Einstein and the others ascribed to the hidden variables they envisaged.
+Thus, the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and the others, in their (EPR) paper, ascribed to the quantum description of quantum nonlocality.
+Of course, in both cases what is incomplete is the ordinary quantum description of a quantum correlation.
+By the way, as is well known, the EPR paper inspired works of great fundamental and practical interest like Bell’s inequality $\left[33\right]$ and Ekert’s quantum cryptography based on Bell’s theorem $\left[34\right]$.
+We do not understand why the argument concerning the incompleteness of the quantum description, central in that paper, has been removed by the scientific community.
+As is also well known, the question of the incompleteness of the quantum description has been at the center of the famous Einstein-Bohr dispute.
+Bohr, who was for the completeness, was eventually decreed the winner by the scientific community.
+Apparently that was once and for all.
+According to the present line of thinking, also this position of science should be changed.
+To us, of course, Einstein should have been the winner.
+Another time, he would have put his finger on a very sensitive point in the quantum mechanical theory.
+Eventually, as we shall see in Section 4., under a quantum cosmological interpretation of the Anthropic Principle, also the evolution of the universe toward the value of life turns out to be a unitary evolution between two one-to-one correlated measurement outcomes (the value of the fundamental constants since the beginning of time and the value of life today) that does not change the eigenvalue selected by the initial measurement (the value of the fundamental constants).
+As such, even if the search for the value of life by the universe’s evolution under the fundamental physical laws and constants were random, it would be endowed with quantum speedup and therefore teleological in character.
+3.2 Formalization
+We formalize the notion of the time-symmetric causal loops on the simplest, four drawer, instance of Grover quantum search algorithm.
+At the end of this subsection we will see how this immediately generalizes.
 The formalization consists of the following steps:
 (1) Extend the usual quantum description of the quantum algorithm, limited to the process of solving the problem, to the upstream process of setting the problem.
-We call this extended quantum description the ordinary quantum description of the quantum algorithm, ordinary because it is still a causal description that cannot describe causal loops.
+We call this extended description the ordinary quantum description, of the quantum algorithm: ordinary because it is still a causal description that cannot describe causal loops.
 (2) This point is propaedeutic to point (4).
-It shows how to relativize with respect to the problem-solver the ordinary quantum description of the quantum algorithm, which is with respect to the customary external observer.
-Of course, the problem-solver must be concealed from the problem-setting selected by the initial measurement, which would give her information about the solution of the problem.
+It shows how to relativize with respect to the problem-solver the ordinary quantum description of the quantum algorithm, which is with respect to the customary observer external to the quantum process.
+Of course, the problem-solver must be concealed from the problem-setting selected by the initial measurement, which would give her full information about the solution of the problem555 We must assume that the problem-solver knows everything about the structure of the problem, thus also the relation between problem-setting and solution.
+.
 (3) Complete the ordinary quantum description of the quantum algorithm by time-symmetrizing it.
-Mathematically, the quantum description of the process between the two correlated measurement outcomes remains unaltered.
-However, the description of the behavior of causality along it changes: the single causal process of the ordinary quantum description is replaced by a quantum superposition of time-symmetric causal loops.
-(4) Relativize the quantum description of the causal loops generated at point (3) with respect to the problem-solver.
-This produces the quantum description of the quantum superposition of causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
-4.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
+(4) Relativize with respect to the problem-solver the time-symmetric causal loops generated by the time-symmetrization of point (3).
+This produces the quantum superposition of causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
+3.2.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
 The usual quantum description of quantum algorithms is limited to the process of solving the problem, where the quantum speedup resides, and lacks the initial measurement.
@@
 We extend the usual quantum description to the upstream process of setting the problem.
-An initial measurement of the problem-setting, performed in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings, selects one of them at random.
+An initial measurement of the problem-setting in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random.
 This is followed by the unitary evolution that computes the corresponding solution and the final measurement of the solution.
@@
 Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors.
-By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow444 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball.
-We omit this operantion that changes nothing for us.
+By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow666 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball.
+We omit this operation that would change nothing.
 .
@@
 Note that, of $\hat{\mathbf{U}}_{1,2}$, it is sufficient to know that it does this, whereas we will never need to know how it does it.
-Of course, we must also know that there exists a unitary transformation that does this, Grover algorithm in the present case.
+Of course, we must also know that there exists a unitary evolution that does this, it is the unitary part of Grover algorithm in the present case.
+Correspondingly, throughout this section, the unitary part of Grover algorithm will always remain just the symbol $\hat{\mathbf{U}}_{1,2}$.
 At the end of Alice’s problem-solving action, at time $t_{2}$, register $A$ contains the solution of the problem, i.e. the number of the drawer with the ball – bottom-right corner of the diagram.
 Eventually, by measuring $\hat{A}$ (the content of $A$), which is already in one of its eigenstates, Alice acquires the solution with probability $1$ – the quantum state in the bottom-right corner of the diagram naturally remains unaltered throughout this measurement.
-Correspondingly, there is a unitary transformation between the initial and final measurement outcomes; the process between them is reversible despite comprising the final measurement.
-4.2 Relativizing the ordinary quantum description of the quantum algorithm with respect to the problem-solver
+Correspondingly, there is a unitary transformation that connects the initial and the final measurement outcomes; the process between them is reversible despite comprising the final measurement (which does not change the quantum state).
+3.2.2 Relativizing the ordinary quantum description of the quantum algorithm with respect to the problem-solver
 The ordinary quantum description of the quantum algorithm of Table I is with respect to the customary external observer, namely an observer external to the quantum process.
@@
 In fact, it would tell her, immediately after the initial measurement of $\hat{B}$, the number of the drawer with the ball selected by Bob, which is also the solution of the problem – see the state of register $B$ in the bottom-left corner of Table I diagram.
-Of course, in the quantum description of the quantum algorithm with respect to Alice, she must be concealed from the outcome of the initial measurement555 More in general, one should assume that the problem-solver has a complete knowledge of the problem, thus also of the relation between problem-setting and solution.
-Knowledge of the problem-setting selected by the initial measurement would tell her also the solution.
-.
+Of course, in the quantum description of the quantum algorithm with respect to Alice, she must be concealed from the outcome of the initial measurement.
 Since Alice’s problem-solving action never changes the number of the drawer with the ball selected by Bob and the observables $\hat{B}$ and $\hat{A}$ commute, describing this concealment is simple.
 It suffices to postpone after the end of Alice’s problem-solving action the projection of the quantum state associated with the initial measurement of $\hat{B}$ (or, indifferently, to postpone the very measurement of $\hat{B}$).
-The result is the following description of the quantum algorithm with respect to Alice:
+The result is the following quantum description of the quantum algorithm with respect to Alice:
 time $\begin{array}[]{ccc}\begin{array}[]{c}\text{time}\text{ }t_{1}\text{, meas.
@@
 Since $\hat{B}$ and $\hat{A}$ commute, the final measurement of $\hat{A}$ must project this superposition on the tensor product of the number of the drawer with the ball initially selected by Bob and the corresponding solution – bottom-right corner of Table II diagram.
-Note that this relativization relies on relational quantum mechanics $\left[33,34\right]$, where the quantum state is relative to the observer.
+Note that this relativization relies on relational quantum mechanics $\left[35,36\right]$, where the quantum state is relative to the observer.
 In the following, we will continue to call the quantum description with respect to the external observer the ordinary quantum description of the quantum algorithm.
 Instead we will always specify that the other quantum description is that with respect to the problem-solver.
-4.3 Time-symmetrizing the ordinary quantum description of the quantum algorithm
+3.2.3 Time-symmetrizing the ordinary quantum description of the quantum algorithm
 As we have already said, in order to obtain the quantum superposition of the causal loops whose mutually exclusive $or$ is implied by the sheer existence of an optimal quantum speedup, we should first time-symmetrize the ordinary quantum description of the quantum algorithm, then relativize with respect to the problem-solver the causal loops obtained.
@@
 Note that, since the symbol “$\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$” is mathematically equivalent to “$\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$”, this instance, up to the time-direction of causality, is the unitary process between the two one-to-one correlated measurement outcomes back again.
-Also note that causality performs a causal loop that changes its own past; the outcome of its initial measurement of $\hat{B}_{l}$ is changed into the outcome of the initial measurement of $\hat{B}$ of the ordinary quantum description; compare the state of register $B$ under the left vertical arrow with the state in the bottom-left corner of the diagram.
+Also note that causality performs a causal loop that changes its own past; the outcome of its initial measurement of $\hat{B}_{l}$ is changed into the outcome of the initial measurement of $\hat{B}$ of the ordinary quantum description; compare the state of register $B$ under the left vertical arrow with the state in the bottom-left corner of Table III diagram.
 We should eventually take the quantum superposition of all the time-symmetrization instances.
@@
 Correspondingly, its subsequent backward in time propagation by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ (bottom line of the diagram), mathematically (disregarding the direction of causality), is always the ordinary quantum description of the unitary process between the two correlated measurement outcomes, namely the bottom line of Table I diagram.
-Summing up, mathematically, time-symmetrization leaves the unitary process between the two measurement outcomes unaltered but changes the structure of causality along it.
+Summing up, mathematically, time-symmetrization leaves the quantum description of the unitary process between the two measurement outcomes unaltered but changes the structure of causality along the process.
 The causal process of the ordinary quantum description is replaced by a quantum superposition of causal loops.
-4.4 Relativizing the causal loops with respect to the problem-solver
+3.2.4 Relativizing the causal loops with respect to the problem-solver
 The causal loops generated by time-symmetrizing the ordinary quantum description of the quantum algorithm, which is relative to the external observer, do not yet map on those logically implied by the sheer existence of the quantum speedup.
@@
 time $\begin{array}[]{ccc}\text{time }t_{1}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\\ &&\\ \begin{array}[]{c}(|01\rangle_{B}+|11\rangle_{B})\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table V)
-Note that the entire causal loop is the zigzag diagram of table IV.
+The entire causal loop is the zigzag diagram of table IV.
 It is a causal loop that changes its own past.
-The initial state in the top-left part of the diagram, where the state of register $B$ tells that Alice is completely ignorant of the number of the drawer with the ball, changes into the initial state in the bottom-left part of the diagram, where the state of register $B$ tells that Alice knows in advance, before beginning her problem-solving action, that the ball is in drawer either $01$ or $11$.
+The initial state in the top-left part of the diagram, where the state of register $B$ tells us that Alice is completely ignorant of the number of the drawer with the ball, changes into the initial state in the bottom-left part of the diagram, where the state of register $B$ tells us that Alice knows in advance, before beginning her problem-solving action, that the ball is in drawer either $01$ or $11$.
 Since what happens along the time-symmetric causal loop obeys classical logic, the number of drawer openings needed to find the solution in the optimal quantum case is that needed in the optimal logical way, namely just one.
@@
 In equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place quadratically reduces.
-The generalization of the above is straightforward.
-Note that, in all our considerations, we never needed to know the unitary transformation between the two measurement outcomes $\hat{\mathbf{U}}_{1,2}$.
-It sufficed to know that it unitarily changes its input state, where the two observables ($\hat{B}$ and $\hat{A}$ ) are unrelated, into its output state, where they are one-to-one correlated, without changing the eigenvalue of the observable selected by the initial measurement.
-First, the causal loops explanation of the quantum speedup, and the fact that the latter is quadratic in kind, remain there if we increase the number of quantum bits in the two registers of an even number and in an equal way.
-In fact, all our considerations remain identical.
-This means, in particular, Grover algorithm with number of drawers any even power of $2$.
-The same, of course, holds, for any unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement.
-This will apply to the teleological evolutions addressed in the next section.
-5 The physical basis of the notion of teleological evolution
-We have seen, by the way rather simply, that the complete quantum description of quantum algorithms is a quantum superposition of time-symmetric causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps.
-In the optimal quantum case, the number of computation steps is that needed to solve the problem in the optimal logical way benefiting of the advanced knowledge in question.
-In equivalent terms, the computational complexity of the problem to be solved or the dimension of the Hilbert space in which the search for the solution takes place correspondingly reduces.
-As already observed, in quantum search algorithms, also a random quantum search for the solution is endowed with (a suboptimal) quantum speedup.
-Therefore, we can say that quantum algorithms are evolutions toward a goal (the solution of the problem) with an attractor in the goal they will produce in the future.
-Of course, this is necessarily a causal loop and responds to the notion of teleological evolution.
-As is well known, before the scientific revolution, the evolutions of the living were often considered to be teleological in character.
-The teleological notion was then dismissed with the advent of modern science for the alleged absence of a physical basis.
-That was, of course, under the causal view of evolutions that is dominant in science also today.
-Now, the fact that quantum algorithms turn out to be teleological evolutions should already cast a doubt on this way of seeing.
-It also, naturally, raises the question whether the time-symmetric quantum causal loops that stand at the basis of their teleological character could be the missing physical basis of the teleological character of natural evolutions.
-As already seen in the Rationale, under a quantum cosmological interpretation of the Anthropic Principle, the search for the value of life by the evolution of the universe under the fundamental physical laws and constants has the same form a quantum search algorithm.
-This search can therefore occur with quantum speedup and does even if it is random.
-Consequently, the evolution of a universe toward the value of life turns out to be teleological in character.
-Of course, in the universe we are in, what turns out to be teleological in character is the evolution of the universe toward sentient life.
-Eventually, as we have seen, the evolution of life on earth, in principle obtainable from that of the universe by tracing out all the rest, even if decohered, should inherit the quantum speedup and the consequent teleological character of the evolution of the universe toward life.
-In the end, the notion that the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithm are also the physical basis of the teleological character of natural evolutions would seem to be physically plausible, this, of course, within the quantum cosmological view.
-This section is a discussion of the many impacts that the present teleological view of natural evolutions might have on different areas of science and knowledge; of course, not forgetting the natural philosophy that gave rise to the notion of teleological evolution.
-5.1 Decohered evolutions of life endowed with quantum speedup: how would they look?
-All natural evolutions are necessarily described in a probabilistic way.
-We can think of the stochastic description of a Darwinian evolution: an alternation of random changes in the species DNA pool each followed by natural selection of the DNA changes more suited to the environment.
-Of course, in a classical stochastic process, transition probabilities can depend on the past of the process but not on the future of it.
-Instead, a decohered process endowed with quantum speedup must have an attractor in its future.
-This means that those DNA mutations that in the future will survive natural selection must make, in the present, the transition toward them more probable.
-By the way, since the consequent teleological character of the evolution would be hidden in transition probabilities, the process would be difficult to distinguish from a classical stochastic process.
-Since also in this case the time-symmetric causal loops we are dealing with should be involved, let us call time-symmetric also the just outlined kind of stochastic evolution.
-We mention what could be an emblematic example of time-symmetric stochastic evolution.
-It is known that, for many millions of years, the ancestors of the flying birds were feathered dinosaurs who did not fly.
-Very light, strong but flexible, cave, and air-proof, feathers are of course ideal for flight.
-So, why anticipating something designed for flight millions of years in advance?
-The usual explanation is of course a causal one: at first, feathers appeared through random mutations and survived natural selection as good thermal insulators, later they proved useful for speeding escape from predators, perhaps for gliding from trees to the ground, and only eventually they turned out to be what was needed to conquer the air.
-According to the present teleological view of the evolutions of the living, even during the flightless period of the feathered dinosaurs, those random DNA mutations that would have led in the future to the conquest of the air – obviously an extremely successful outcome from the point of view of survival – would have more likely occurred.
-The above is puzzling, of course.
-We should keep in our mind that we are under the quantum cosmological assumption that might make macroscopic evolutions, the partial trace over all the rest of a quantum evolution of the universe, as strange as the microscopic quantum world.
-5.2 Teleological evolution of the universe toward life – a toy model
-A toy model of a teleological evolution of the universe toward life can be trivially mapped on the simplest instance of Grover algorithm.
+The generalization of all the above is straightforward.
+Of course, everything immediately generalizes to any number of drawers.
+Note that, in all our considerations, we never needed to know the unitary transformation ($\hat{\mathbf{U}}_{1,2}$) between the two measurement outcomes.
+It sufficed to know that there exists a unitary transformation that changes its input state, where the two observables $\hat{B}$ and $\hat{A}$ are unrelated, into its output state, where they are one-to-one correlated, without changing the eigenvalue of the observable selected by the initial measurement.
+4 The possible physical basis of natural teleological evolutions
+We have seen in the Introduction how, under a quantum cosmological interpretation of the Anthropic Principle, the evolution of the universe toward the value of life (in particular, either sentient life or absence of life) under the fundamental physical laws and constants has the form of a quantum search algorithm.
+As such, it should be endowed with quantum speedup and would therefore be teleological in character even in the case that the search for the value of life is assumed to be random.
+In the universe we are in, the search for the value of life becomes search for life, of course endowed with the same quantum speedup.
+We have also seen that, as a consequence, the evolution of life on Earth, obtainable from that of the universe by tracing out all the rest, even if decohered, would inherit the quantum speed up of the evolution of the universe toward life and would therefore be teleological too.
+The same would necessarily hold for the evolutions of species, etc..
+In this section, after encoding a toy model of the evolution of the universe toward the value of life on Grover quantum search algorithm, we will discuss the many impacts that the present physical notion of teleological natural evolution should have on different areas of science and knowledge.
+4.1 Teleological evolution of the universe toward the value of life – a toy model
+A toy model of a teleological evolution of the universe toward the value of life can be trivially mapped on the simplest instance of Grover algorithm.
 We should imagine a value of the fundamental physical constants observable $\hat{B}$ and a value of life observable $\hat{A}$.
@@
 Initially, $\hat{B}$ should be in a quantum superposition (or, indifferently, mixture) of all its eigenstates (i.e of all the possible universes each with its fundamental physical constants).
-We can assume that also $\hat{A}$, the value of life observable, is initially in a quantum superposition of all its eigenstates.
+We can assume that also $\hat{A}$, the value of life observable, is initially in an independent quantum superposition of all its eigenstates.
 The universe’s unitary evolution under the fundamental physical laws and constants would have finally, at the present time, built up a quantum superposition of tensor products, each an eigenstate of $\hat{B}$ tensor product the corresponding eigenstate of $\hat{A}$ – see Table II.
@@
 As shown by its time-symmetrization (Table IV), it can undergo with a quantum speedup, and does even if the quantum search for the value of sentient life is random.
-5.3 How to model a teleological natural evolution
+We should note that the present toy model is less crude than it may seem.
+In fact, it immediately generalizes to any number $N$ of drawers.
+This means that any amount of information can be encoded in the process we are dealing with.
+Moreover, the unitary transformation $\hat{\mathbf{U}}_{1,2}$ between the initial and final one-to-one correlated measurement outcomes is any unitary transformation that does not change the eigenvalue selected by the initial measurement.
+Therefore, in principle, this toy model could even become a faithful discrete model of the actual evolution of the universe toward the value of life.
+4.2 Decohered teleological evolutions: what would they look like?
+The supposedly teleological evolutions of life would be, so to speak, around us in the macroscopic world.
+Decohered and endowed with quantum speedup it is natural to wonder what they would look like.
+Now, all natural evolutions are necessarily described in a probabilistic way.
+We can think of the stochastic description of a Darwinian evolution: an alternation of random changes in the species DNA pool each followed by natural selection of the DNA changes more suited to the environment.
+Of course, in a classical stochastic process, transition probabilities can depend on the past of the process but not on the future of it.
+Instead, a decohered process teleological in character must have an attractor in its future.
+This means that those DNA mutations that in the future will survive natural selection because more fit to the environment must make, when they occur in the present, the transition toward them more probable.
+By the way, since the consequent teleological character of the evolution would be hidden in transition probabilities, the evolution would be difficult to distinguish from a classical stochastic evolution.
+Let us call time-symmetric the present teleological kind of stochastic evolution, given that its description would be generated by a time-symmetrization procedure.
+Let us mention what could be an emblematic example of time-symmetric stochastic evolution.
+It is known that, for many millions of years, the ancestors of the flying birds were feathered dinosaurs who did not fly.
+Very light, strong but flexible, cave, and air-proof, feathers are of course ideal for flight.
+So, why anticipating something designed for flight millions of years in advance?
+The usual explanation is of course a causal one: at first, feathers appeared by random mutation and survived natural selection as good thermal insulators, later they proved useful for speeding escape from predators, perhaps for gliding from trees to the ground, and only eventually they turned out to be what was needed to conquer the air.
+According to the present teleological view of the evolutions of the living, even during the flightless period of the feathered dinosaurs, those random DNA mutations that would have led in the future to the conquest of the air – obviously an extremely successful outcome from the point of view of survival – would have more likely occurred.
+The above explanation is puzzling, of course.
+We should keep in our mind that we are under the quantum cosmological assumption that should make macroscopic evolutions, the partial trace over all the rest of a quantum evolution of the universe, as puzzling as the microscopic quantum world.
+4.3 How to model teleological natural evolutions
 We address the problem of how to model a decohered evolution endowed with quantum speedup – say that of the living on Earth after the beginning of the liquid water stage of it.
 The canonical way would be to obtain this evolution from that of the universe toward life by tracing out all the rest.
-Of course, in the present case, this holds only in principle.
-However, it would be interesting to apply the canonical way to a toy model of the situation we are dealing with.
-One should identify the purely quantum evolution of a system (the universe) endowed with quantum speedup, where the achievement of the goal is mostly localized in a subsystem thereof (the Earth).
+Although, in the present case, this obviously holds only in principle, it would be interesting to apply it to a toy model of the situation.
+One should identify the purely quantum evolution of a system (standing for the universe) endowed with quantum speedup, where the achievement of the goal is mostly localized in a subsystem thereof (standing for the Earth).
 Then, one should obtain a decohered evolution of the subsystem from that of the system by tracing out all the rest.
@@
 If it were completely independent, it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup.
-Therefore, an obvious (approximate) way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact it is decohered and model it in a quantum way.
+Therefore, an obvious, approximate, way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact it is decohered and model it in a quantum way.
 There are encouraging precedents to this.
-For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[35\right]$.
+For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[37\right]$.
 As already said, a further possibility should be to model a decohered evolution endowed with quantum speedup as a stochastic process with transition probabilities time-symmetrically depending on both the past and the future of the process, something just outlined for the time being.
@@
 On the other, as this vision is considered without physical basis and therefore misleading, there is outright hostility to it.
-This tension is captured by the following sentence by the evolutionary biologist Haldane $\left[36\right]$:
+This tension is captured by the following sentence by the evolutionary biologist Haldane $\left[38\right]$:
 Teleology is like a mistress to a biologist: he cannot live without her, but he is unwilling to be seen with her in public.
@@
 No biologist, of course, does not mean no scientist.
-A paradigmatic example is the American philosopher Nagel $\left[37\right]$, who argues that the principles that explain the emergence of life should be teleological, rather than materialistic or mechanistic in character.
+A paradigmatic example is the American philosopher Nagel $\left[39\right]$, who argues that the principles that explain the emergence of life should be teleological, rather than materialistic or mechanistic in character.
 We would put together Nagel’s view in support of the teleological character of the evolutions of the living and Haldane’s assertion that a biologist cannot live without teleology to affirm that the notion of teleological natural evolution is more “scientific” than the mainstream one of causal natural evolution.
 This in the sense that it better answers empirical evidence – of course the first canon of science.
-The only reason biologists would be against teleology in spite of empirical evidence would come from another science.
+The only reason biologists would be against teleology despite empirical evidence would come from another science.
 We mean the notion – current in today’s physics – that a teleological evolution is physically impossible.
 Of course, the sheer notion that quantum computation is teleological in character may already put a strain on the position of today’s biology: actually, although in the quantum world, teleological evolutions are physically possible.
-The strain should increase with the notion that, under a quantum cosmological interpretation of the Anthropic Principle, the evolution of the universe toward life should be teleological in character.
+The strain might increase with the notion that, under a quantum cosmological interpretation of the Anthropic Principle, the evolution of the universe toward the value of life, toward life in the universe we are in, should be teleological in character.
 It should further increase with the consequent notion that also the evolution of life on Earth, obtainable in principle from that of the universe by tracing out all the rest, even if decohered, should inherit the teleological character of the universe’s evolution.
 Summing up, given the above, the impact of the present teleological view of natural evolutions on today’s evolutionary biology would seem to be almost a head on collision.
-Almost because, on the one side, in recent years a group of evolutionary biologists $\left[38\right]$ expressed the need to rethink Darwin’s theory of evolution.
+Almost because, on the one side, in recent years a group of evolutionary biologists $\left[40\right]$ expressed the need to rethink Darwin’s theory of evolution.
 That was after the discovery that some species’ evolution steps occurred far too quickly to be solely ascribed to random changes in the species’ DNA pool.
 Of course, the present teleological view of evolutions, with the associated quantum speedup, would offer a way out.
-On the other side, there are scientists who have began to apply essentially quantum mechanical features like quantum entanglement and nonlocality to the macroscopic game of species, showing that this can explain some empirically ascertained features of it otherwise difficult to explain $\left[35\right]$.
+On the other side, there are scientists who have began to apply essentially quantum mechanical features like quantum entanglement and nonlocality to the macroscopic game of species, showing that this can explain some empirically ascertained features of it otherwise difficult to explain $\left[37\right]$.
 4.5 Impact on philosophy
-Of course, the current theory of teleological evolutions would rehabilitate an ancient notion of natural philosophy $\left[39\right]$.
-We should therefore be allowed to look for the possibility that there are other relations between this theory and philosophy.
+Of course, the current theory of teleological natural evolutions has a precise conceptual precedent in a very ancient notion of natural philosophy $\left[41\right]$.
+We should therefore be allowed to look for other possible relations between this theory and philosophy.
 As we are going to see, their abundance is a surprising interdisciplinary aspect of the present work.
 In effect, the teleological theory in question seems to revamp a full body of ancient philosophical views sort of opposite (and complementary) to the body of philosophical views emerged with the scientific revolution.
-Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view666 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops.
+Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view777 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops.
 and the causal view.
@@
 Consider the opposition between materialism and idealism.
-The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[31\right]$, the anthropic cosmological principles of Barrow and Tipler $\left[13\right]$, and the present theory is a mutual causality view.
-We find a similar situation in the opposition between the mechanistic and teleological views of evolutions, which obviously respond to the causal and mutual causality views of evolutions, respectively.
-For what concerns reductionism vs monism, we note that deriving the teleological character of the decohered evolutions of the living systems on Earth from that of the evolution of a quantum universe is an essentially monistic view.
-It presupposes that what we see around and within us is essentially inseparable from the whole.
+The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[31\right]$, and the anthropic cosmological principles of Barrow and Tipler $\left[12\right]$ is a mutual causality view.
+We find a similar situation in the opposition between the mechanistic and teleological views of evolutions, which obviously respond, respectively, to the causal and mutual causality views of evolutions.
+Regarding reductionism vs.
+monism, we note that deriving the teleological character of the decohered evolutions of living systems on Earth from that of the evolution of a quantum universe is an essentially monistic view.
+It presupposes that the description of what we see around and within us is essentially inseparable from the description of the whole.
 Let us mention other theories in the history of philosophy reminiscent of the mutual causality between sentient life today and the physical laws with their physical constants since the beginning of time:
-(i) For Plato, “Ideas [of our mind] are objective perfections that exist in themselves and for themselves and, at the same time, are the causes of natural phenomena and constitute their unity” $\left[40\right]$.
+(i) For Plato, “Ideas [of our mind] are objective perfections that exist in themselves and for themselves and, at the same time, they are the causes of natural phenomena and constitute their unity” $\left[42\right]$.
 In modern language, the cause of natural phenomena that constitute their unity are the fundamental physical laws.
@@
 It is as if the number of the drawer with the ball was determined by both the problem-setter and the problem-solver.
-Also Plato’s notion that ideas – physical laws – are objective perfections could have a correspondence with the fine tuned universe version of the Anthropic Principle, when it asserts that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
-The above should also go along with Fritjov Capra’s $\left[41\right]$ notion that there is identity between the fundamental states of consciousness described by eastern theosophies and the fundamental laws of quantum physics.
+Also Plato’s notion that ideas – physical laws – are objective perfections could have a correspondence with the fine tuned universe version of the Anthropic Principle, the fact it asserts that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
+The above should also go along with Fritjov Capra’s $\left[43\right]$ notion that there is identity between the fundamental states of consciousness described by eastern theosophies and the fundamental laws of quantum physics.
 (ii) We do not need to be anthropocentric.
@@
 The assumption that spatial intelligence is achieved with quantum speedup excludes that it is only the result of “ordinary” natural evolution (obeying the principle of causality).
-Indeed a quantum speedup in the development of spatial intelligence would imply the mutual causality in question.
+Indeed a quantum speedup in the development of spatial intelligence – and the consequent teleological character of this development – would imply the mutual causality in question.
 (iii) The present notion of teleological evolution, namely of an evolution toward a goal with an attractor in the goal it will reach in the future, could also be seen in terms of a “systematically lucky” evolution, lucky from the standpoint of the objective of reaching its goal, indeed because attracted by it.
 Interestingly, also the notion of a systematically lucky evolution seems to have a precedent in pre-scientific thinking.
-It would be the Stoic notion of $\pi\rho o\nu o\iota\alpha$, foreknowledge $\left[42\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge.
+It would be the Stoic notion of $\pi\rho o\nu o\iota\alpha$, foreknowledge $\left[44\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge.
 Foreknowledge, or advanced knowledge, would have pushed the evolution toward its goal in a luckier and thus faster way than what is possible in the case of a causal evolution, exactly as the advanced knowledge about the solution does in quantum algorithms.
 This notion of $\pi\rho o\nu o\iota\alpha$ would also fit well with the notion of teleological evolution understood as a stochastic processes with transition probabilities time-symmetrically affected by the past and future of the process, with the luckier transitions more probable than in the classical case.
-(iv) It is natural to wander whether the present time-symmetric quantum causal loops view of the universe’s evolution toward sentient life, where the sentient life stage is an attractor of the evolution of the universe toward it, could be related to Nietzsche’s intuition of an eternal return consequent to our will for it.
+(iv) It is natural to wander whether the present time-symmetric causal loops view of the universe’s evolution toward sentient life, where the sentient life stage is an attractor of the evolution of the universe toward it, could be related to Nietzsche’s intuition of an eternal return consequent to our will for it.
 (v) Michelangelo’s “Creation of Adam” on the ceiling of the Sistine Chapel could be the manifesto of the present theory.
-There is a mirror symmetry between creator and created $\left[43\right]$.
+There is a mirror symmetry between creator and created $\left[45\right]$.
 Could this be interpreted as the time-symmetry – and the related mutual causation – between the fundamental physical laws with their physical constants since the beginning of time and the existence of sentient life today?
-Furthermore, the fact that the figures and shapes portrayed behind the creator figure would seem to be an anatomically accurate portrait of the human brain $\left[44\right]$, could it allude to the teleological character of the evolution of our intelligent species: the human brain being an attractor of the search for it?
+Furthermore, the fact that the figures and shapes portrayed behind the creator figure would seem to be an anatomically accurate portrait of the human brain $\left[46\right]$, could it allude to the teleological character of the evolution of our intelligent species: the human brain being an attractor of the search for it?
 We would not feel like dismissing these analogies as sheer coincidences.
 As a matter of fact, in idealistic philosophy – in Schelling – aesthetic insight would reproduce the fundamental mechanisms of the production of life.
+Overall, it would seem that the current ”scientific” theory of teleological evolutions reevaluates the idealistic side of philosophy against the materialistic side.
+However, the latter would also have its revenge, since a scientific justification of idealism could be ascribed to a materialistic view.
+To sum up, it seems clear that the contrast between the causal and mutual causality views touches a deep mechanism in the history of thought.
+In an inevitably biased way, we would ascribe it to the unifying power of the current theory of teleological evolutions.
 5 Research prospects
@@
 For what concerns quantum physics:
-(i) Look for a synthesis between the present notion of the time-symmetric quantum causal loops and time-symmetric quantum mechanics $\left[20-24\right]$ to which it is obviously inspired.
-(ii) Look for toy models that capture the mechanism for which the evolution of life on Earth (the mostly autonomous evolution of a subsystem), although decohered, inherits the quantum speedup and consequently the teleological character of the evolution of the universe (the system) toward life.
-(iii) The mechanism described in (ii) might apply to other life processes for which the presence of a quantum behavior is suspected, for example, to the possible quantum behavior of the brain.
-Although mostly autonomous from the rest of the universe, this behavior could derive from that of a quantum universe by tracing out everything external to the brain.
-Note the radical difference between this view and the usual one where people looks for an autonomous quantum behavior in the macroscopic, warm and wet brain.
-By the way, we could introduce a distinction between two kinds of quantum behaviors: autonomous, like those of microscopic systems sufficiently isolated from all the rest, and derived, namely quantum behaviors of even macroscopic subsystems obtained from that of a quantum universe – the system – by tracing out all the rest.
-(iv) Look for the possible impact of the teleological view on quantum cosmology.
-According to current assumptions, even the inanimate part of the latter, always an evolution toward life, should be endowed with quantum speedup and thus have teleological character.
-(v) The Anthropic Principle is limited to stating the correlation between the “value” of life and the value of the fundamental physical constants since the beginning of time.
-One could also look for a correlation between the value of life and the character of the fundamental physical laws since the beginning of time.
-Of course, also these would have an attractor in the goal of creating life.
+(i) Look for a synthesis between the present notion of the time-symmetric causal loops and time-symmetric quantum mechanics $\left[19-23\right]$ to which it is obviously inspired.
+(ii) Look for toy models that capture the mechanism for which the mostly autonomous evolution of a subsystem, although decohered, inherits the quantum speedup and consequent teleological character of the evolution of a purely quantum system.
+Of course, here, the quantum system stands for the universe and the subsystem for the living Earth subsystem.
+Note that the present notion for which the decohered behavior of a subsystem inherits quantum features from the behavior of the quantum system might apply to other cases.
+Possible examples are the life processes suspected to have higher than classical efficiency and consciousness suspected to need quantum nonlocality, the quantum system being here, of course, the universe.
+(iii) Look for a possible impact of the present teleological view on quantum cosmology.
+According to current assumptions, even the “inanimate” part of the universe’s evolution, an evolution toward life too, should be endowed with quantum speedup and thus have teleological character.
+(iv) The Anthropic Principle is limited to asserting the correlation between the “value” of life today and the value of the fundamental physical constants since the beginning of time.
+One could also look for a relation between the value of life today and the character of the fundamental physical laws since the beginning of time.
+Of course, in the universe we are in, also the latter character should have an attractor in the goal of creating life.
 About the natural sciences, looking for:
@@
 (II) Toy models of teleological Darwinian processes.
-These evolutions could be either approximated as pure quantum processes endowed with quantum speedup or described as decohered stochastic processes with transition probabilities time-symmetrically affected by the past and the future of the process.
-(III) An application of the time-symmetric quantum causal loops to quantum self-organization.
-An example is studying the teleological formation of a self-replicating system.
-There seem to be an enormous jump of complexity in the transition from a soup of life-precursor molecules into a self-replicating system.
+These evolutions could be either approximated by pure quantum processes endowed with quantum speedup or described as decohered stochastic processes with transition probabilities time-symmetrically affected by the past and the future of the process.
+(III) An application of the time-symmetric causal loops to quantum self-organization.
+An example could be studying the teleological formation of a self-replicating system.
+There might be an enormous jump of complexity in the transition from a soup of life-precursor molecules to a self-replicating system.
 The idea that this transition has an attractor in the full fledged self-replicating system can therefore be attractive.
@@
 6 Conclusions
-The work is, of course, highly unconventional. It might be useful to summarize in a consequential way the line of reasoning followed.
-We have (trivially) shown that the sheer existence of the optimal quantum speedup of Grover quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops.
+It might be useful to summarize the key points of our unconventional argument.
+We have trivially shown that the sheer existence of the quantum speedup of the optimal Grover quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops.
 In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and used this knowledge to produce the solution with correspondingly fewer computation steps.
 The number of computation steps needed to solve the problem in an optimal quantum way is that needed to solve it in an optimal logical way using the advanced knowledge in question.
-Since quantum algorithms only involve the measurements of commuting observables, they are submitted to the same (classical) logic.
-Therefore, Grover algorithm must host the quantum superposition of the causal loops in question.
-However, being causal in character as that of any other quantum process, its customary quantum description cannot describe causal loops; we ascribe this to its incompleteness.
-Before going to the way of remedying this incompleteness, we must eliminate another, trivial, one.
-The customary quantum description of quantum algorithms is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution.
-It is incomplete since it lacks the initial measurement.
-Thus, as a first step, we extend the customary quantum description of the process of solving the problem to the upstream process of setting the problem, by adding the initial measurement of the problem-setting in a quantum superposition (or mixture) of all the possible problem-settings.
-We have thus a unitary evolution between two one-to-one correlated measurement outcomes that does not change the problem-setting selected by the initial measurement.
-This is a key feature of quantum algorithms, not visible in their customary quantum description777 If the solution is not an invertible function of the problem-setting, to have one-to-one correlation between the two measurement outcomes, it suffices that the problem-solver, at the end, measures both the solution and the problem-setting.
-.
-We call this kind of process the ordinary quantum description of the quantum algorithm, ordinary because it is still a causal description that cannot describe causal loops.
-The second step is to change the causal structure of the ordinary quantum description of the quantum algorithm by time-symmetrizing the latter.
-We time-symmetrically share between the initial and final measurements the selection of the information that specifies the sorted out pair of correlated measurement outcomes, in a quantum superposition of all the possible ways.
-Each time-symmetrization instance, mathematically, yields the unitary evolution between the two measurement outcomes back again, but the structure of causality along it changes.
-It becomes a causal loop that changes its own past.
-Relativized with respect to the problem-solver (who should be concealed from any outcome of the initial measurement), the causal loop in question becomes one of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
-Eventually, we should take the quantum superposition of all the causal loops.
-Of course, the time-symmetrization leaves the ordinary quantum description mathematically unaltered but changes its single causal process into a quantum superposition of causal loops.
-Until now, the time-symmetrization in question has been imposed by the necessity of reproducing, in the quantum description of Grover algorithm, the causal loops logically implied by the sheer existence of its (optimal) quantum speedup.
-However, as we have seen, the same time-symmetrization is also redundantly imposed by a symmetry implicit in any unitary evolution between an initial and a final, one-to-one correlated, measurement outcomes that does not change the eigenvalue selected by the initial measurement.
-As a consequence, any such evolution is a quantum superposition of time-symmetric causal loops.
-If the unitary evolution builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement, like in the case of Grover algorithm, in each time-symmetrization instance the dimension of the Hilbert space in which the search for one-to-one correlation takes place reduces.
-This creates the room for a quantum speedup.
-There is a quantum speedup even in the case that the search for one-to-one correlation is random.
-The process is a quantum superposition of time-symmetric causal loops also in the case that the two observables are already maximally entangled at the beginning of it.
-It is the case of quantum nonlocality, where the quantum process starts with the spatial separation between two subsystems whose respective observables are already maximally entangled.
-Here the causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space-separate, where they locally tell either observable the state to assume in view of its future measurement.
-They play the role of the hidden variables envisaged by Einstein and the others to locally explain quantum nonlocality.
-The quantum causal loop description of quantum algorithms highlights their teleological character, namely their being evolutions toward a goal (the solution) with an attractor in the goal they will produce in the future (the solution again).
-This naturally raises the question whether the time-symmetric quantum causal loops we are dealing with could also be the physical basis of the teleological character of the evolutions of the living for which the notion of teleological evolution was conceived.
-This immediately raises the problem of explaining a quantum speedup in evolutions of the living that can be the size of the Earth and have durations of billions of years.
-The fact it is obviously impossible to scale up the quantum features of the microscopic world to such levels suggests the only possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe.
-Of course, at the same time we should adopt quantum cosmology, namely the assumption that the universe evolves in a unitary quantum way.
-We should now call into play the “Fine-tuned Universe” version of the Anthropic Principle.
+On the one hand, being trivial, this claim is perfectly verifiable.
+On the other hand, a trivial demonstration of the existence of causal loops, a notion strongly questioned, should therefore be a given.
+This can be considered the fulcrum of the present work, and the source of its unconventionality.
+If, so simple, it has never been noticed outside the present retrocausal approach to the quantum speedup, it must be because a tacit no-retrocausality precept has prevented looking that way.
+The following is the consequent line of thinking.
+Since quantum algorithms only involve the measurements of commuting observables, they are submitted to the same (classical) logic of the logical implication in question.
+Therefore, Grover algorithm must host the quantum superposition of the causal loops we are dealing with.
+However, of course, this cannot be in its quantum description that, being causal in character as that of any quantum process, cannot describe causal loops.
+It should be noted that the usual quantum description of quantum algorithms presents another problem.
+It is limited to the process of solving the problem where the quantum speedup mathematically resides and is clearly incomplete.
+In fact, it consists of the unitary computation of the solution and the measurement of the latter.
+It lacks the initial measurement.
+Remedying this incompleteness is another key point of the work.
+Extending the usual quantum description to the upstream process of setting the problem by measuring the problem-setting in the quantum superposition (or mixture) of all the possible problem-settings makes the quantum description of the quantum algorithm time-symmetric.
+The quantum description of Grover algorithm becomes that of a unitary evolution between an initial and a final one-to-one correlated measurement outcomes.
+Of course, being still causal, this extended quantum description cannot describe causal loops.
+However, time-symmetrizing it generates the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup.
+We should require that the initial and final measurements time-symmetrically contribute to the selection of their pair of one-to-one correlated measurement outcomes.
+The selection performed by the initial measurement should propagate forward in time until becoming the state immediately before the final measurement.
+The selection performed by the final measurement should propagate backward in time until replacing the previous outcome of the initial measurement.
+This is a causal loops that changes its own past – the initial measurement outcome.
+It is indeed one of the causal loops logically implied by the sheer existence of the quantum speedup.
+Of course, we should take the quantum superposition of all the possible time-symmetrization instances, each for each possible way of time-symmetrically sharing the selection of the pair of measurement outcomes between the initial and final measurements.
+This leaves the quantum description of the unitary evolution between the initial and final measurement outcomes mathematically unaltered but changes the description of the structure of causality along that evolution.
+The single causal process of the ordinary quantum description is replaced by the quantum superposition of causal loops logically implied by the sheer existence of the quantum speedup.
+Being generated by a time-symmetrization, we call these causal loops time-symmetric.
+Note that the time-symmetrization in question turns out to be the only possible way of obtaining a quantum description of the quantum algorithm that describes the causal loops the quantum algorithm must host.
+It is therefore mandatory.
+We have shown that this immediately generalizes to any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement (the problem-setting in quantum algorithms).
+The ordinary quantum description of any such process is incomplete and is completed by time-symmetrizing it.
+This leaves the ordinary quantum description of the unitary part of the quantum process mathematically unaltered but replaces its single causal process by a quantum superposition of time-symmetric causal loops.
+In particular, the generalization covers the case of quantum nonlocality, where the unitary evolution is the process of spatial separation of the two subsystems in a state of maximal entanglement between the respective observables.
+There is one-to-one correlation between the outcomes of measuring these observables in the same basis.
+Here, the time-symmetric quantum causal loops play the role of the hidden variables envisaged by Einstein and the others to complete the ordinary quantum description of quantum nonlocality.
+Eventually, the quantum causal loop description of quantum algorithms highlights their teleological character, that is, their being evolutions toward a goal (the solution) with an attractor in the goal they will produce in the future (the solution again).
+This naturally raises the question whether the time-symmetric causal loops we are dealing with could also be the physical basis of the teleological character of the evolutions of the living for which the notion of teleological evolution was conceived.
+We believe that investigating this possibility is incumbent on the part of physics toward the natural sciences, after its having denied for centuries the possibility that teleological evolutions exist.
+This immediately raises the problem of explaining the presence of quantum features (the quantum speedup) in evolutions of the living that can be the size of the Earth and have durations of billions of years.
+The obvious impossibility that the quantum features of the microscopic world scale up to such levels suggests the only possible solution: that they scale down from the most macroscopic and enduring evolution there is, that of the universe.
+Of course, at the same time we should assume, with quantum cosmology, that the evolution of the universe is unitary quantum.
+The fact that, once given the amount of resources involved, the quantum speedup is related to the time taken by the quantum process, under suitable conditions makes it inheritable by a subsystem (a natural evolution) of the quantum system (the universe).
+Surprisingly, the “Fine-tuned Universe” version of the Anthropic Principle puts all the above things together.
 It asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value of the fundamental physical constants since the beginning of time and the “value” of life today, in particular, either “existence of sentient life” or “absence of life”.
 We assume for simplicity that this correlation is one-to-one.
-With the evolution of the universe quantum, also the correlation asserted by the Anthropic Principle becomes quantum.
-We are thus dealing with a unitary evolution that one-to-one correlates two initially unrelated observables (the value of the fundamental constants and the value of life observables) without changing the eigenvalue of the observable selected by the initial measurement.
-We have seen that this kind of unitary evolution can occur with quantum speedup and does so even in the case that the search for one-to-one correlation is random.
-The evolution of the universe toward the value of life should therefore take place with quantum speedup and thus be teleological in character.
-In the universe we are in, this would apply to the evolution of the universe toward sentient life.
-We have also seen that the evolution of life on Earth, in principle obtainable from that of the universe toward sentient life by tracing out all the rest, should inherit the teleological character of the latter evolution.
+The assumption that the evolution of the universe is unitary quantum, of course forces us to adopt a quantum interpretation of the Anthropic Principle.
+It is now the unitary evolution of the universe that builds up a quantum correlation between the value of the physical constants since the beginning of time and the value of life today.
+This is formally identical to a quantum search algorithm that unitarily builds up a quantum correlation between the value of the problem-setting since its beginning and the value of the solution and its end.
+We have seen that any unitary evolution that builds up one-to-one correlated between an initial and a final measurement outcome without changing the eigenvalue selected by the initial measurement can occur with quantum speedup and does even if the search for that correlation is random.
+The latter is of course a conservative assumption; that there has been that building up by the evolution of the universe is what the Anthropic Principle tells us.
+In conclusion, the unitary evolution of the universe toward the value of life occurs with quantum speedup and is thus teleological.
+In the universe we are in, the same, of course, holds for the evolution of the universe toward life.
+The evolution of life on Earth, in principle obtainable from that of the universe toward life by tracing out all the rest, should inherit the teleological character of the latter evolution.
 In fact: (i) evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe and (ii) the resources employed by the former evolution cannot exceed those employed by an image that comprises them.
-In conclusion, the notion of the time-symmetric quantum causal loops turns out to be the common physical basis of such diverse physical phenomena as the quantum computational speedup, the consequent teleological character of quantum algorithms, quantum nonlocality, and – under a quantum cosmological interpretation of the Anthropic Principle – the teleological character of the evolutions of the living.
-Of course, that of teleological evolution is a puzzling notion that radically changes a long-standing way of viewing the role of causality in the physical world.
-Regarding the quantum phenomena in the microscopic world, we think to have provided clear evidence that the notion of the time-symmetric quantum causal loops is a necessary consequence of the notion of quantum computational speedup.
-Regarding the teleological character of the macroscopic evolutions of life, the situation becomes even more puzzling.
-We limit ourselves to noticing that this is a necessary consequence of the quantum cosmological assumption.
-Seeing the macroscopic evolutions of life as the partial trace over everything else of a quantum evolution of the universe toward life brings into the macroscopic world of everyday experience some of the strange characters of the microscopic quantum world.
+In conclusion, we have seen that the ordinary, causal, quantum description of a class of processes involving quantum correlation is incomplete, exactly in the sense argued by Einstein and the others in their EPR paper, and is completed by time-symmetrizing it.
+This explains such diverse quantum phenomena as the quantum speedup and the consequent teleological character of quantum algorithms, quantum nonlocality, and, under a quantum cosmological interpretation of the Anthropic Principle, the teleological character of the evolutions of life.
+It also recovers a pre-scientific view of physical reality that had waned with the advent of modern science and its classical-causal view of physical phenomena.
 Acknowledgments

### v7

@@
 We show that, under a quantum cosmological interpretation of the Anthropic Principle, the same quantum superposition of causal loops underlying the teleological character of quantum algorithms becomes the missing physical basis of the teleological character of natural evolutions.
-1 Introduction
+1 Extended summary
 In this work we propose a physical basis for the notion of teleological evolution, that is, an evolution toward a goal with an attractor in the same goal it will produce in the future.
@@
 What we think should change that (also centuries old) position of science is the relatively recent discovery of the quantum computational speedup.
-In this work, by quantum (computational) speedup we mean its strict form, that is, the fact that quantum algorithms solve problems with fewer logical operations (computation steps) than classically possible.
-As we try to be interdisciplinary, let us provide an example of quantum speedup right away.
+In this work, by quantum (computational) speedup we mean its strict form, that is, the fact that quantum algorithms solve problems with fewer logical operations (computation steps) than classically possible – sometimes, the quantum speedup is with respect to the best known classical algorithm but one does not whether it is over what is classically possible.
+As we try to be interdisciplinary, let us provide an example of (strict) quantum speedup right away.
 Consider the following problem.
@@
 We will show that this is the immediate consequence of a retrocausal approach to the quantum speedup of ours and co-authors $\left[5-9\right]$.
-The present work is divided into two parts.
+We chose the extended summary form because of the unconventionality of the work: there are many concepts that need to be explained from the beginning.
+The work is divided into two parts.
 The first is a review and further clarification of that retrocausal approach to the quantum speedup: it will justify the teleological character of quantum algorithms.
-The second is whether the same justification could constitute the physical basis of the teleological character of the evolutions of the living for which the teleological notion was conceived.
+The second is whether the physical basis of the teleological character of quantum algorithms could also be the physical basis of a teleological character of the evolutions of the living for which the teleological notion was conceived.
 The present introduction is divided into two corresponding parts.
@@
 We can conclude that the optimal quantum speedup of Grover algorithm logically implies the mutually exclusive $or$ of well defined causal loops, each for each possible way of taking half of the information about the solution.
-In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution by opening fewer drawers.
+In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and used this knowledge to produce the solution by opening fewer drawers.
 We can formulate the following advanced knowledge rule that, as checked in $\left[8\right]$, also holds for all the major quantum algorithms for the solution of oracle problems:
 The number of computation steps (oracle queries) needed to solve the (oracle) problem in an optimal quantum way is the number needed to solve it in an optimal logical way benefiting of the advanced knowledge of one (any one) of the possible halves of the information about the solution.
-Later we will show why this rule applies to any problem for which a problem setting and a corresponding solution can be identified.
+Later we will show why this rule applies to any problem for which a problem-setting and a corresponding solution can be identified.
 By the way, it is a rule that does not exist in mainstream quantum computing science.
@@
 As is well known, this is a matter of high ingenuity.
-Of course, this goes along with the fact that mainstream quantum computing science, as mainstream quantum mechanics, relies on the notion of causal evolution that is incompatible with that of causal loop.
+Of course, this goes along with the fact that mainstream quantum computing science, as mainstream quantum mechanics, relies on the notion of causal evolution (i.e., with causality always going in the same time-direction) that is incompatible with that of causal loop.
 Let us make a digression.
@@
 So, we would like to emphasize that the forest is right now in front of our eyes.
-The notion of causal loops is obviously controversial; probably most people believe that causal loops cannot exist except perhaps in very exotic physical situations such as the closed timelike curves allowed by general relativity.
+The notion of causal loops is obviously controversial; probably most people believe that causal loops cannot exist except perhaps in rather exotic physical situations.
 Now we would have, moreover trivially, demonstrated that the quantum speedup of a quantum algorithm logically implies their existence.
 This is the forest!
-As we will venture into very exotic situations ourselves, but in ordinary non-relativistic quantum mechanics, we would like to suggest that we keep this basic fact in mind: it is the source of all exoticism.
-Back to our line of thinking, of course we should ask ourselves what is the impact of the fact that a quantum speed up logically implies the mutually exclusive $or$ of causal loops on our way of seeing the quantum algorithm, that is, on the quantum description that we make of it.
+Since this will have puzzling consequences, we suggest keeping in mind this fact that the very existence of the quantum speedup implies causal loops.
+Of course, it is the causal loops that are the source of all the weirdness in question.
+Back to our line of thinking, we should ask ourselves what is the impact of the fact that a quantum speed up logically implies the mutually exclusive $or$ of those causal loops on our way of seeing the quantum algorithm, that is, on the quantum description that we make of it.
 To this end, we should make reference to a quantum description of the quantum algorithm less incomplete than the usual one, which is limited to the process of solving the problem (where the quantum speedup resides) and lacks the initial measurement.
 Of course, the canonical quantum description of a quantum process must include the initial and final measurements and the unitary evolution in between.
-Of course, this problem is easily adjusted.
+Naturally, this problem is easily adjusted.
 It suffices to extend the usual quantum description of the process of solving the problem to the upstream process of setting the problem.
 The extended quantum description is thus as follows.
-An initial measurement, in a quantum superposition (or mixture) of all the possible problem settings, selects one of them at random222 The problem-setter would be free to unitarily change the randomly selected problem-setting into a wanted one.
-We disregard this possibility that, in the present context, changes nothing.
-.
-This is followed by the unitary computation of the corresponding solution and by the final measurement that reads it without changing the quantum state333 With four drawers, Grover algorithm yields the solution of the problem with certainty; when the drawers are more than four, by Grover algorithm we will always mean Long’s version of it $\left[10\right]$, which always provides the solution with certainty.
+An initial measurement, in a quantum superposition (or mixture) of all the possible problem settings, selects one of them at random.
+This is followed by the unitary computation of the corresponding solution and by the final measurement that reads the solution without changing the quantum state222 With four drawers, Grover algorithm yields the solution of the problem with certainty; when the drawers are more than four, by Grover algorithm we will always mean Long’s version of it $\left[10\right]$, which always provides the solution with certainty.
 .
@@
 This extension of the usual quantum description of quantum algorithms to the process of setting the problem will play a key role in the present work.
-We will call it the ordinary quantum description of the quantum algorithm, ordinary because it is still the ordinary causal description of mainstream quantum mechanics.
+Let us call it the ordinary quantum description of the quantum algorithm, ordinary because it is still the ordinary, causal, quantum description of mainstream quantum mechanics.
 We can go now to the impact of the logical implication in question on our way of describing the quantum algorithm.
@@
 Of course, we should then take the quantum superposition of all the possible ways of evenly sharing.
-We will see that all this, mathematically, leaves the unitary part of the ordinary quantum description of the quantum algorithm unaltered but changes the description of the behavior of causality along that part.
+We will see that all this, mathematically, leaves the description of the unitary part of the quantum algorithm unaltered but changes that of the behavior of causality along that part.
 The single, unitary, causal process of the ordinary quantum description of the quantum algorithm changes into the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup.
 We will (easily) show that this time-symmetrization is the only possible way to generate the quantum superposition of the causal loops in question, which in turn must be hosted within the quantum algorithm.
-Therefore, according to the current line of thinking, it is imposed by logic and quantum mechanics.
+Therefore, according to the present line of thinking, it is imposed by logic and quantum mechanics.
 Most importantly to present ends, the fact that, in each causal loop, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and used this knowledge to produce the solution with fewer computation steps highlights the teleological character of quantum algorithms.
@@
 The two things together make the results obtained in the case of Grover algorithm applicable to a very general class of quantum processes involving quantum correlation. In fact, they imply that:
-Any unitary evolution between an initial and final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is the quantum superposition of the causal loops generated by time-symmetrizing its ordinary quantum description.
+The complete quantum description of any unitary evolution between an initial and final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is the quantum superposition of the causal loops generated by time-symmetrizing its ordinary quantum description.
 We call the above the causal loop rule of quantum correlation.
 We can see that it is a generalization also when the process is still a quantum algorithm.
-In fact, until now we have remained committed to Grover algorithm that is an optimal one.
+In fact, until now, all the results obtained are for an optimal quantum algorithm, Grover algorithm.
 However, also the time-symmetrization of the ordinary quantum description of non-optimal quantum algorithms yields a quantum superposition of causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution.
-The point is that this time she may not exploit this information at all, or not fully exploit it.
-For example, the latter is the case of the random quantum search algorithms $\left[11\right]$, which provides a suboptimal quantum speedup.
+The point is that this time she is free to exploit any amount of this information.
+For example, this covers the case of the random quantum search algorithm $\left[11\right]$, which provides a suboptimal quantum speedup.
 We will be interested in this kind of quantum algorithms further on.
-In view of applying the causal loop rule of quantum correlation to processes that have the form of quantum algorithms but are not that, let us note that we can do without the anthropomorphic notion of a problem-solver that produces the solution with fewer computation steps because she knows in advance half of the information about it.
-We can also avoid to speak of a “quantum algorithm”.
+In view of applying the causal loop rule of quantum correlation to processes that have the form of quantum algorithms but are not that, let us note that we can do without the anthropomorphic notion of a problem-solver that produces the solution with fewer computation steps because she knows in advance half of the information about it; we can also avoid to speak of “quantum algorithm”.
 We can say that, in each time-symmetrization instance, there is a reduction of the dimension of the Hilbert space in which the search for the one-to-one correlation between the initial and final measurement outcomes takes place and consequently the correlation in question can be achieved with a quantum speedup.
@@
 The two observables can be maximally entangled before the process begins.
-As we shall see, quantum nonlocality is exactly a process of this kind and the quantum superposition of causal loops generated by its time-symmetrization explains the EPR (Einstein, Podolsky, and Rosen) paradox.
+As we shall see, quantum nonlocality is exactly a process of this kind and the quantum superposition of causal loops generated by its time-symmetrization explains the EPR (Einstein, Podolsky, and Rosen) paradox – it provides a local explanation of quantum nonlocality.
 1.2 The possible teleological character of the evolutions of the living
 The teleological character of quantum algorithms naturally raises the question of whether the quantum superposition of causal loops that underlie it can also be the missing physical basis of the teleological character of the evolutions of the living for which the teleological notion was conceived.
-Having established the teleological character of quantum algorithms, probing this possibility would even seem to be a duty of the physical science to the natural sciences.
+Having established that, although in the quantum world, there is a physical basis for the notion of teleological evolution, probing this possibility could even be considered a duty of the physical science toward the natural sciences.
 One difficulty, of course, would be to explain the presence of quantum features – the quantum speedup – in the evolutions of life that can be Earth-sized and last billions of years.
-However, the obvious impossibility that the quantum features of the microscopic world scale up to such levels opens up another possibility: that they scale down from the largest and longest existing evolution, that of the universe.
+However, the obvious impossibility that the quantum features of the microscopic world scale up to such levels points to another possibility: that they scale down from the largest and longest existing evolution, that of the universe.
 Of course, at the same time, we should adopt quantum cosmology $\left[12,13\right]$, i.e., the assumption that the universe – a perfectly isolated system since by definition there is nothing outside it – evolves in a unitary quantum way.
-Interestingly, the corresponding quantum cosmological interpretation of the Anthropic Principle $[14-18]$ brings all things together.
+Surprisingly, the corresponding quantum cosmological interpretation of the Anthropic Principle $[14-18]$ brings all things together.
 The Fine-tuned Universe version of this principle asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental constants since the beginning of time and the “value” of life today (in particular: “presence of sentient life” or “absence of life”).
-The reason for this correlation is that the slightest variation in the value that the fundamental constants have in the universe we are in would produce a radically different universe, most likely incapable of hosting life.
+As is well known, this principle stems from the fact that the slightest variation in the value of the fundamental constants would have produced a radically different universe, most likely incapable of hosting life.
 Let us assume, for simplicity, that this correlation is one-to-one.
@@
 These quantum algorithms yield a quantum speedup even if the search for the correlation is random $\left[11\right]$.
-It is indeed reasonable (and conservative) to assume that the search for the correlation between the value of the fundamental constants since the beginning of time and the value of life today by the universe’s evolution – the search for life in the universe we are in – is random.
+It is indeed reasonable (and conservative) to assume that the search for this correlation by the universe’s evolution under the fundamental physical laws and constants – its search for life in the universe we are in – is random.
 According to this formal identity, the evolution of the universe toward the value of life should occur with quantum speedup and would consequently be teleological in character.
@@
 In fact, evolution times should be the same for the evolution of life on Earth and its image in the evolution of the universe and the resources employed by the former evolution cannot exceed those involved by the latter that comprises them.
-Of course, that the universe evolves in a unitary quantum way, although consistent with the theory of decoherence, for which quantum processes decohere by losing information in their environment, is just an assumption.
-It should instead be firm the derivation from it of the teleological character of the evolutions of the living.
-In fact, the Fine-tuned Universe version of the Anthropic Principle is universally accepted and the quantum cosmological interpretation of it is straightforward.
-Summing up, we would conclude with the conditional statement that, if quantum cosmology were right, the evolutions of the living would be teleological in character.
-Let us also observe that this would be a paradoxical possibility.
-For example a way of making teleological a Darwinian evolution could be to assume that those random DNA mutations that, in the future, will turn out to survive because more suited to the environment, because of that, would become more probable.
-By the way, this would be exactly as paradoxical as the causal loop explanation of the quantum speedup.
-The only difference is that the quantum computational speedup belongs to a quantum world that is notoriously the place of queer phenomena.
-A teleological Darwinism would instead concern the macroscopic world we see around us.
-In effect, evolutions attracted by their future would be a radically different way of viewing physical reality.
+Summing up, coherently with the present line of thinking, we should conclude with the conditional statement that, if the universe evolved in a unitary quantum way, the evolutions of the living would be teleological in character.
 In the next Section 2, we try to position the present work within the literature that challenges the still dominant notion of causal evolution.
@@
 Of course, we are not alone in challenging the predominant idea that causality only goes forward in time.
-Other approaches to quantum mechanics that share this feature are time-symmetric quantum mechanics $\left[19-23\right]$, particularly the study of the possibility that a future choice changes a past quantum measurement outcome $\left[22\right]$, and ongoing attempts to study quantum processes without the need to assume a definite causal structure $\left[24-27\right]$.
-The present retrocausal approach to the quantum speedup is naturally inspired by time-symmetric quantum mechanics and has in common with the attempts in question the fact of looking beyond the ordinary notion of causality, not in common the way of looking at quantum algorithms.
-For example, in $\left[26\right]$, quantum algorithms are considered examples of processes with a definite causal structure because, in their current quantum description, causality always goes forward in time.
-So far so good but, consequently, quantum algorithms are excluded from the phenomena that can accommodate causality violations.
-Of course, for us, this means putting too much faith in their current quantum description.
-To us, with their quantum computational speedups, quantum algorithms are the seminal example of causality violation.
+Other approaches to quantum mechanics that share this feature are time-symmetric quantum mechanics $\left[19-23\right]$, particularly the study of the possibility that a future choice changes a past quantum measurement outcome $\left[22\right]$, and ongoing attempts to study quantum processes without the need to assume a definite causal structure $\left[24-28\right]$.
+The present retrocausal approach to the quantum speedup is naturally inspired by time-symmetric quantum mechanics and, of course, has in common with the attempts in question the fact of looking beyond the ordinary notion of a causality that always goes in the same time-direction, not in common the approach.
+In fact, these attempts explicitly rule out the possibility that quantum algorithms, with their causally ordered quantum circuits, can accommodate causality violations $\left[26,27\right]$.
+For us instead, with their quantum speedup over what is classically possible, they are the paradigmatic example of causality violation.
+Evidently, the two approaches address different kinds of causality violations.
 Some concepts of the work have precedents.
-One is the celebrated EPR paper $\left[28\right]$.
+One is the celebrated EPR paper $\left[29\right]$.
 In it, Einstein, Podolsky, and Rosen already argued that the quantum description is incomplete.
@@
 It generates a quantum superposition of causal loops that play the role of the hidden variables those authors envisaged to complete the quantum description.
-Another precedent is the local, retrocausal, explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[29\right]$.
+Another precedent is the local, retrocausal, explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[30\right]$.
 Consider the measurements in the same basis of two maximally entangled observables in a state of spatial separation of the respective subsystems.
@@
 Therefore, first, they must be there independently of the fact of explaining quantum non-locality locally; second, as long as quantum mechanics is consistent, they should be exempt from the paradoxes that plague the usual kind.
-Interestingly, at the time, Costa de Beauregard was an assistant of De Broglie, who forbade him to publish that work for years until the publication, by Wheeler and Feynman, of their emitter-absorber theory that also resorted to retrocausality – see $\left[30\right]$ for an account of this.
+Interestingly, at the time, Costa de Beauregard was an assistant of De Broglie, who forbade him to publish that work for years until the publication, by Wheeler and Feynman, of their emitter-absorber theory that also resorted to retrocausality – see $\left[31\right]$ for an account of this.
 This tells us a lot on the physicist’s position toward retrocausality, even at the top levels.
 By the way, despite the importance of the author’s name, Costa de Beauregard’s explanation of quantum nonlocality never took off.
-Still other precedents are Wheeler’s Participatory Anthropic Principle $\left[31,32\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[12\right]$.
+Still other precedents are Wheeler’s Participatory Anthropic Principle $\left[32,33\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[12\right]$.
 These two principles share with the present work the notion of mutual causality between the observation today and the “value” of the fundamental physical laws and constants since the beginning of time mediated by a quantum evolution of the universe in between.
@@
 3.1 Updating the EPR critique of the quantum description
-Of course, the first assertion that the quantum description is incomplete goes back to the famous EPR paper by Einstein, Podolsky and Rosen $\left[28\right]$.
+Of course, the first assertion that the quantum description is incomplete goes back to the famous EPR paper by Einstein, Podolsky and Rosen $\left[29\right]$.
 As is well known, Einstein in particular, argued that the ordinary quantum description of quantum nonlocality is incomplete because it does not physically explain the “spooky action at a distance” that it mathematically predicts.
@@
 We call the above the causal loop rule of quantum correlation. Let us examine a few of its consequences.
-In the first place, to be a quantum superposition of causal loop, the quantum algorithm does no more need to be optimal444 If the solution is not an invertible function of the problem setting, we can have one-to-one correlation between the initial and final measurement outcomes by requiring that the final measurement measures both the solution and the problem-setting.
+In the first place, to be a quantum superposition of causal loop, the quantum algorithm does no more need to be optimal333 If the solution is not an invertible function of the problem setting, we can have one-to-one correlation between the initial and final measurement outcomes by requiring that the final measurement measures both the solution and the problem-setting.
 .
@@
 Of course, in both cases what is incomplete is the ordinary quantum description of a quantum correlation.
-By the way, as is well known, the EPR paper inspired works of great fundamental and practical interest like Bell’s inequality $\left[33\right]$ and Ekert’s quantum cryptography based on Bell’s theorem $\left[34\right]$.
+By the way, as is well known, the EPR paper inspired works of great fundamental and practical interest like Bell’s inequality $\left[34\right]$ and Ekert’s quantum cryptography based on Bell’s theorem $\left[35\right]$.
 We do not understand why the argument concerning the incompleteness of the quantum description, central in that paper, has been removed by the scientific community.
@@
 It shows how to relativize with respect to the problem-solver the ordinary quantum description of the quantum algorithm, which is with respect to the customary observer external to the quantum process.
-Of course, the problem-solver must be concealed from the problem-setting selected by the initial measurement, which would give her full information about the solution of the problem555 We must assume that the problem-solver knows everything about the structure of the problem, thus also the relation between problem-setting and solution.
+Of course, the problem-solver must be concealed from the problem-setting selected by the initial measurement, which would give her full information about the solution of the problem444 We must assume that the problem-solver knows everything about the structure of the problem, thus also the relation between problem-setting and solution.
 .
@@
 Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors.
-By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow666 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball.
+By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow555 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball.
 We omit this operation that would change nothing.
@@
 Since $\hat{B}$ and $\hat{A}$ commute, the final measurement of $\hat{A}$ must project this superposition on the tensor product of the number of the drawer with the ball initially selected by Bob and the corresponding solution – bottom-right corner of Table II diagram.
-Note that this relativization relies on relational quantum mechanics $\left[35,36\right]$, where the quantum state is relative to the observer.
+Note that this relativization relies on relational quantum mechanics $\left[36,37\right]$, where the quantum state is relative to the observer.
 In the following, we will continue to call the quantum description with respect to the external observer the ordinary quantum description of the quantum algorithm.
@@
 In the universe we are in, the search for the value of life becomes search for life, of course endowed with the same quantum speedup.
-We have also seen that, as a consequence, the evolution of life on Earth, obtainable from that of the universe by tracing out all the rest, even if decohered, would inherit the quantum speed up of the evolution of the universe toward life and would therefore be teleological too.
+We have also seen that, as a consequence, the evolution of life on Earth, obtainable from that of the universe by tracing out all the rest, even if decohered, should inherit the quantum speed up of the evolution of the universe toward life and would therefore be teleological too.
 The same would necessarily hold for the evolutions of species, etc..
@@
 The above explanation is puzzling, of course.
+However, let us note that it is not more puzzling than the complete quantum description of the quantum computational speedup.
 We should keep in our mind that we are under the quantum cosmological assumption that should make macroscopic evolutions, the partial trace over all the rest of a quantum evolution of the universe, as puzzling as the microscopic quantum world.
@@
 There are encouraging precedents to this.
-For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[37\right]$.
+For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[38\right]$.
 As already said, a further possibility should be to model a decohered evolution endowed with quantum speedup as a stochastic process with transition probabilities time-symmetrically depending on both the past and the future of the process, something just outlined for the time being.
@@
 On the other, as this vision is considered without physical basis and therefore misleading, there is outright hostility to it.
-This tension is captured by the following sentence by the evolutionary biologist Haldane $\left[38\right]$:
+This tension is captured by the following sentence by the evolutionary biologist Haldane $\left[39\right]$:
 Teleology is like a mistress to a biologist: he cannot live without her, but he is unwilling to be seen with her in public.
@@
 No biologist, of course, does not mean no scientist.
-A paradigmatic example is the American philosopher Nagel $\left[39\right]$, who argues that the principles that explain the emergence of life should be teleological, rather than materialistic or mechanistic in character.
+A paradigmatic example is the American philosopher Nagel $\left[40\right]$, who argues that the principles that explain the emergence of life should be teleological, rather than materialistic or mechanistic in character.
 We would put together Nagel’s view in support of the teleological character of the evolutions of the living and Haldane’s assertion that a biologist cannot live without teleology to affirm that the notion of teleological natural evolution is more “scientific” than the mainstream one of causal natural evolution.
@@
 It should further increase with the consequent notion that also the evolution of life on Earth, obtainable in principle from that of the universe by tracing out all the rest, even if decohered, should inherit the teleological character of the universe’s evolution.
-Summing up, given the above, the impact of the present teleological view of natural evolutions on today’s evolutionary biology would seem to be almost a head on collision.
-Almost because, on the one side, in recent years a group of evolutionary biologists $\left[40\right]$ expressed the need to rethink Darwin’s theory of evolution.
+Obviously, all this would go against the ban currently placed by the natural sciences on the teleological view.
+Instead, it might go along with the fact that, in recent years, a group of evolutionary biologists $\left[41\right]$ expressed the need to rethink Darwin’s theory of evolution.
 That was after the discovery that some species’ evolution steps occurred far too quickly to be solely ascribed to random changes in the species’ DNA pool.
-Of course, the present teleological view of evolutions, with the associated quantum speedup, would offer a way out.
-On the other side, there are scientists who have began to apply essentially quantum mechanical features like quantum entanglement and nonlocality to the macroscopic game of species, showing that this can explain some empirically ascertained features of it otherwise difficult to explain $\left[37\right]$.
+Of course, the present teleological view of evolutions, with the associated quantum speedup, would offer a possible explanation.
+It would also go along with the fact that some scientists have began to apply essentially quantum mechanical features like quantum entanglement and nonlocality to the macroscopic game of species, showing that this can explain some empirically ascertained features of it otherwise difficult to explain $\left[38\right]$.
 4.5 Impact on philosophy
-Of course, the current theory of teleological natural evolutions has a precise conceptual precedent in a very ancient notion of natural philosophy $\left[41\right]$.
-We should therefore be allowed to look for other possible relations between this theory and philosophy.
-As we are going to see, their abundance is a surprising interdisciplinary aspect of the present work.
+Of course, the current theory of teleological natural evolutions has a precise conceptual precedent in a very ancient notion of natural philosophy $\left[42\right]$.
+We should therefore be allowed to look for other possible cases where philosophical thinking has anticipated aspects of it.
 In effect, the teleological theory in question seems to revamp a full body of ancient philosophical views sort of opposite (and complementary) to the body of philosophical views emerged with the scientific revolution.
-Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view777 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops.
+Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view666 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops.
 and the causal view.
-The causal view and the mutual causality view seem to have respectively subtended two opposing views throughout the history of philosophy.
+The causal view and the mutual causality view seem to have respectively subtended two opposing school of thought throughout the history of philosophy.
 The causal view would underlie the materialist, mechanistic and reductionist views, while the mutual causality view would underlie the idealist, teleological and monist views respectively.
 Consider the opposition between materialism and idealism.
-The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[31\right]$, and the anthropic cosmological principles of Barrow and Tipler $\left[12\right]$ is a mutual causality view.
+The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[32\right]$, and the anthropic cosmological principles of Barrow and Tipler $\left[12\right]$ is a mutual causality view.
 We find a similar situation in the opposition between the mechanistic and teleological views of evolutions, which obviously respond, respectively, to the causal and mutual causality views of evolutions.
@@
 Let us mention other theories in the history of philosophy reminiscent of the mutual causality between sentient life today and the physical laws with their physical constants since the beginning of time:
-(i) For Plato, “Ideas [of our mind] are objective perfections that exist in themselves and for themselves and, at the same time, they are the causes of natural phenomena and constitute their unity” $\left[42\right]$.
+(i) For Plato, “Ideas [of our mind] are objective perfections that exist in themselves and for themselves and, at the same time, they are the causes of natural phenomena and constitute their unity” $\left[43\right]$.
 In modern language, the cause of natural phenomena that constitute their unity are the fundamental physical laws.
@@
 Also Plato’s notion that ideas – physical laws – are objective perfections could have a correspondence with the fine tuned universe version of the Anthropic Principle, the fact it asserts that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
-The above should also go along with Fritjov Capra’s $\left[43\right]$ notion that there is identity between the fundamental states of consciousness described by eastern theosophies and the fundamental laws of quantum physics.
+The above should also go along with Fritjov Capra’s $\left[44\right]$ notion that there is identity between the fundamental states of consciousness described by eastern theosophies and the fundamental laws of quantum physics.
 (ii) We do not need to be anthropocentric.
-This mutual causality would also be between physical laws and spatial intelligence, often superior to our in non-human animals.
+This mutual causality would also be between fundamental physical laws and constants on the one side and spatial intelligence, often superior to our in non-human animals, on the other.
 The assumption that spatial intelligence is achieved with quantum speedup excludes that it is only the result of “ordinary” natural evolution (obeying the principle of causality).
@@
 Interestingly, also the notion of a systematically lucky evolution seems to have a precedent in pre-scientific thinking.
-It would be the Stoic notion of $\pi\rho o\nu o\iota\alpha$, foreknowledge $\left[44\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge.
+It would be the Stoic notion of $\pi\rho o\nu o\iota\alpha$, foreknowledge $\left[45\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge.
 Foreknowledge, or advanced knowledge, would have pushed the evolution toward its goal in a luckier and thus faster way than what is possible in the case of a causal evolution, exactly as the advanced knowledge about the solution does in quantum algorithms.
@@
 (iv) It is natural to wander whether the present time-symmetric causal loops view of the universe’s evolution toward sentient life, where the sentient life stage is an attractor of the evolution of the universe toward it, could be related to Nietzsche’s intuition of an eternal return consequent to our will for it.
-(v) Michelangelo’s “Creation of Adam” on the ceiling of the Sistine Chapel could be the manifesto of the present theory.
-There is a mirror symmetry between creator and created $\left[45\right]$.
+(v) Michelangelo’s “Creation of Adam” on the ceiling of the Cappella Sistina could be the manifesto of the present theory.
+There is a mirror symmetry between creator and created $\left[46\right]$.
 Could this be interpreted as the time-symmetry – and the related mutual causation – between the fundamental physical laws with their physical constants since the beginning of time and the existence of sentient life today?
-Furthermore, the fact that the figures and shapes portrayed behind the creator figure would seem to be an anatomically accurate portrait of the human brain $\left[46\right]$, could it allude to the teleological character of the evolution of our intelligent species: the human brain being an attractor of the search for it?
-We would not feel like dismissing these analogies as sheer coincidences.
-As a matter of fact, in idealistic philosophy – in Schelling – aesthetic insight would reproduce the fundamental mechanisms of the production of life.
-Overall, it would seem that the current ”scientific” theory of teleological evolutions reevaluates the idealistic side of philosophy against the materialistic side.
-However, the latter would also have its revenge, since a scientific justification of idealism could be ascribed to a materialistic view.
-To sum up, it seems clear that the contrast between the causal and mutual causality views touches a deep mechanism in the history of thought.
-In an inevitably biased way, we would ascribe it to the unifying power of the current theory of teleological evolutions.
+To sum up, it seems clear that the contrast between the causal and mutual causality views corresponds to an opposition that characterizes the entire history of thought.
+We would ascribe this unified view of the history of philosophy to the unifying power of the present “scientific” notion of teleological evolution that, under quantum cosmology, also extends to the macroscopic world.
 5 Research prospects
@@
 (i) Look for a synthesis between the present notion of the time-symmetric causal loops and time-symmetric quantum mechanics $\left[19-23\right]$ to which it is obviously inspired.
-(ii) Look for toy models that capture the mechanism for which the mostly autonomous evolution of a subsystem, although decohered, inherits the quantum speedup and consequent teleological character of the evolution of a purely quantum system.
-Of course, here, the quantum system stands for the universe and the subsystem for the living Earth subsystem.
+(ii) Look for toy models that capture the mechanism for which the mostly autonomous evolution of a subsystem, although decohered, inherits the quantum speedup and the consequent teleological character of the evolution of a purely quantum system.
+Of course, here, the quantum system stands for the universe and the subsystem for the Earth subsystem.
 Note that the present notion for which the decohered behavior of a subsystem inherits quantum features from the behavior of the quantum system might apply to other cases.
-Possible examples are the life processes suspected to have higher than classical efficiency and consciousness suspected to need quantum nonlocality, the quantum system being here, of course, the universe.
+Possible examples are the life processes suspected to have higher than classical efficiency and consciousness suspected to be endowed with quantum nonlocality, the quantum system being here, of course, a quantum universe.
 (iii) Look for a possible impact of the present teleological view on quantum cosmology.
@@
 Restoring the dismissed notion of teleological evolution would obviously be a paradigm shift for today’s natural sciences.
-(II) Toy models of teleological Darwinian processes.
+(II) Toy models of teleological Darwinian evolutions.
 These evolutions could be either approximated by pure quantum processes endowed with quantum speedup or described as decohered stochastic processes with transition probabilities time-symmetrically affected by the past and the future of the process.
+Note that there might be a divide between the teleological view of evolution and the, of course dominant, causal/classical view.
+An evolution endowed with quantum speedup might not only mean a faster evolution; the mutual causality implicit in it could produce forms of evolution that are precluded to the causal/classical type.
 (III) An application of the time-symmetric causal loops to quantum self-organization.
@@
 It might be useful to summarize the key points of our unconventional argument.
-We have trivially shown that the sheer existence of the quantum speedup of the optimal Grover quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops.
+We have (trivially) shown that the sheer existence of the quantum speedup of the optimal Grover quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops.
 In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and used this knowledge to produce the solution with correspondingly fewer computation steps.
 The number of computation steps needed to solve the problem in an optimal quantum way is that needed to solve it in an optimal logical way using the advanced knowledge in question.
-On the one hand, being trivial, this claim is perfectly verifiable.
-On the other hand, a trivial demonstration of the existence of causal loops, a notion strongly questioned, should therefore be a given.
-This can be considered the fulcrum of the present work, and the source of its unconventionality.
-If, so simple, it has never been noticed outside the present retrocausal approach to the quantum speedup, it must be because a tacit no-retrocausality precept has prevented looking that way.
 The following is the consequent line of thinking.
 Since quantum algorithms only involve the measurements of commuting observables, they are submitted to the same (classical) logic of the logical implication in question.
-Therefore, Grover algorithm must host the quantum superposition of the causal loops we are dealing with.
-However, of course, this cannot be in its quantum description that, being causal in character as that of any quantum process, cannot describe causal loops.
-It should be noted that the usual quantum description of quantum algorithms presents another problem.
-It is limited to the process of solving the problem where the quantum speedup mathematically resides and is clearly incomplete.
-In fact, it consists of the unitary computation of the solution and the measurement of the latter.
-It lacks the initial measurement.
-Remedying this incompleteness is another key point of the work.
-Extending the usual quantum description to the upstream process of setting the problem by measuring the problem-setting in the quantum superposition (or mixture) of all the possible problem-settings makes the quantum description of the quantum algorithm time-symmetric.
-The quantum description of Grover algorithm becomes that of a unitary evolution between an initial and a final one-to-one correlated measurement outcomes.
-Of course, being still causal, this extended quantum description cannot describe causal loops.
-However, time-symmetrizing it generates the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup.
-We should require that the initial and final measurements time-symmetrically contribute to the selection of their pair of one-to-one correlated measurement outcomes.
-The selection performed by the initial measurement should propagate forward in time until becoming the state immediately before the final measurement.
-The selection performed by the final measurement should propagate backward in time until replacing the previous outcome of the initial measurement.
+Therefore, Grover algorithm must host the quantum superposition of the causal loops we are dealing with – logically mutual exclusive possibilities go in quantum superpositions in the quantum world.
+However, of course, this cannot be in the quantum description of the quantum algorithm that, being causal, cannot describe causal loops.
+It should be noted that we adopt a canonical quantum description of quantum processes, consisting of initial measurement, unitary evolution, and final measurement.
+In quantum algorithms, the initial measurement selects the problem-setting in a quantum superposition (or mixture) of all the possible problem-settings, the unitary evolution computes the corresponding solution, and the final measurement reads it.
+Time-symmetrizing this, still causal, ordinary quantum description of the quantum algorithm generates the quantum superposition of the causal loops logically implied by the sheer existence of its quantum speedup.
+We require that the initial and final measurements time-symmetrically contribute to the selection of their pair of one-to-one correlated measurement outcomes; that the selection performed by the initial measurement propagates forward in time until becoming the state immediately before the final measurement and the selection performed by the final measurement propagates backward in time until replacing the previous outcome of the initial measurement.
 This is a causal loops that changes its own past – the initial measurement outcome.
@@
 Of course, we should take the quantum superposition of all the possible time-symmetrization instances, each for each possible way of time-symmetrically sharing the selection of the pair of measurement outcomes between the initial and final measurements.
-This leaves the quantum description of the unitary evolution between the initial and final measurement outcomes mathematically unaltered but changes the description of the structure of causality along that evolution.
+This leaves the quantum description of the unitary evolution between the initial and final measurement outcomes mathematically unaltered but changes the behavior of causality along the evolution.
 The single causal process of the ordinary quantum description is replaced by the quantum superposition of causal loops logically implied by the sheer existence of the quantum speedup.
 Being generated by a time-symmetrization, we call these causal loops time-symmetric.
-Note that the time-symmetrization in question turns out to be the only possible way of obtaining a quantum description of the quantum algorithm that describes the causal loops the quantum algorithm must host.
+Note that the time-symmetrization in question turns out to be the only possible way of obtaining a quantum description of the quantum algorithm with the quantum superposition of causal loops it must host.
 It is therefore mandatory.
@@
 The ordinary quantum description of any such process is incomplete and is completed by time-symmetrizing it.
-This leaves the ordinary quantum description of the unitary part of the quantum process mathematically unaltered but replaces its single causal process by a quantum superposition of time-symmetric causal loops.
+Let us repeat ourselves: this leaves the ordinary quantum description of the unitary part of the quantum process mathematically unaltered but changes the description of the behavior of causality along that part.
+The single causal process of the ordinary quantum description is replaced by a quantum superposition of time-symmetric causal loops.
 In particular, the generalization covers the case of quantum nonlocality, where the unitary evolution is the process of spatial separation of the two subsystems in a state of maximal entanglement between the respective observables.
@@
 This naturally raises the question whether the time-symmetric causal loops we are dealing with could also be the physical basis of the teleological character of the evolutions of the living for which the notion of teleological evolution was conceived.
-We believe that investigating this possibility is incumbent on the part of physics toward the natural sciences, after its having denied for centuries the possibility that teleological evolutions exist.
+It can be argued that investigating this possibility is incumbent on the part of the physical science toward the natural sciences, after its having asserted for centuries that teleological evolutions are unphysical.
 This immediately raises the problem of explaining the presence of quantum features (the quantum speedup) in evolutions of the living that can be the size of the Earth and have durations of billions of years.
 The obvious impossibility that the quantum features of the microscopic world scale up to such levels suggests the only possible solution: that they scale down from the most macroscopic and enduring evolution there is, that of the universe.
-Of course, at the same time we should assume, with quantum cosmology, that the evolution of the universe is unitary quantum.
-The fact that, once given the amount of resources involved, the quantum speedup is related to the time taken by the quantum process, under suitable conditions makes it inheritable by a subsystem (a natural evolution) of the quantum system (the universe).
+Of course, at the same time, we should assume, with quantum cosmology, that the evolution of the universe is unitary quantum.
 Surprisingly, the “Fine-tuned Universe” version of the Anthropic Principle puts all the above things together.
@@
 It is now the unitary evolution of the universe that builds up a quantum correlation between the value of the physical constants since the beginning of time and the value of life today.
-This is formally identical to a quantum search algorithm that unitarily builds up a quantum correlation between the value of the problem-setting since its beginning and the value of the solution and its end.
-We have seen that any unitary evolution that builds up one-to-one correlated between an initial and a final measurement outcome without changing the eigenvalue selected by the initial measurement can occur with quantum speedup and does even if the search for that correlation is random.
-The latter is of course a conservative assumption; that there has been that building up by the evolution of the universe is what the Anthropic Principle tells us.
+We have seen that any unitary evolution that builds up one-to-one correlation between an initial and a final measurement outcome without changing the eigenvalue selected by the initial measurement can occur with quantum speedup and does even if the search for that correlation is random.
+That the search for the correlation is random is of course a conservative assumption; that there has been this search by the evolution of the universe is what the Anthropic Principle says.
 In conclusion, the unitary evolution of the universe toward the value of life occurs with quantum speedup and is thus teleological.
@@
 In fact: (i) evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe and (ii) the resources employed by the former evolution cannot exceed those employed by an image that comprises them.
-In conclusion, we have seen that the ordinary, causal, quantum description of a class of processes involving quantum correlation is incomplete, exactly in the sense argued by Einstein and the others in their EPR paper, and is completed by time-symmetrizing it.
-This explains such diverse quantum phenomena as the quantum speedup and the consequent teleological character of quantum algorithms, quantum nonlocality, and, under a quantum cosmological interpretation of the Anthropic Principle, the teleological character of the evolutions of life.
-It also recovers a pre-scientific view of physical reality that had waned with the advent of modern science and its classical-causal view of physical phenomena.
+In conclusion, we have seen that the ordinary, causal, quantum description of any unitary evolution between an initial and a final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is incomplete and is completed by time-symmetrizing it.
+This generates a quantum superposition of time-symmetric causal loops that explain such diverse quantum phenomena as the quantum speedup and the consequent teleological character of quantum algorithms, quantum nonlocality, and, under a quantum cosmological interpretation of the Anthropic Principle, the teleological character of the evolutions of life.
+Although the time-symmetric causal loops we have been dealing with should be exempt from the paradoxes that plague the usual notion of causal loop insofar as they are imposed by logic and quantum mechanics, they remain puzzling for the consequences they entail.
+Being related to quantum correlation, they can also be thought to be pervasive.
+Once verified, their existence could introduce a paradigm shift in quantum mechanics.
 Acknowledgments
-Thanks are due to Eliahu Cohen, Artur Ekert, Avshalom Elitzur and David Finkelstein for useful comments throughout the whole of the present approach to the quantum speedup, Daniel Sheehan for organizing the San Diego AAAS Pacific Division conferences on retrocausality, a far-looking forum for the discussion of frontier, also unorthodox, physics, and Mario Rasetti for organizing the first series of international workshops on quantum communication and computation, the Elsag Bailey-ISI Turin Villa Gualino workshops of the years 1991-96.
+Thanks are due to Eliahu Cohen, Artur Ekert, Avshalom Elitzur and David Finkelstein for useful comments throughout the whole of the present retrocausal approach to the quantum speedup, Daniel Sheehan for organizing the San Diego AAAS Pacific Division conferences on retrocausality, a far-looking forum for the discussion of frontier, also unorthodox, physics, and Mario Rasetti for organizing the first series of international workshops on quantum communication and computation, the Elsag Bailey-ISI Turin Villa Gualino workshops of the years 1991-96.